A Bibliography of Publications about the Java Programming Language, 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

05 June 2023 Version 1.231

Abstract

This bibliography records books about the Java Programming Language and related software.

Title word cross-reference

3 [DiP18b, FLZ^{+}18, GBC12, JEC^{+}12, ZXL16]. $39.95 [Ano18]. 4 + 1 [SRB18]. _{T P} [LTK17]. C_{p} [AÖ11]. K [PLL^{+}18, SS19, SD16b, SGG^{+}17]. N [ADJG19, WZK^{+}19]. Z_{p} [AÖ11].

-core [PLL^{+}18]. -Means [SS19]. -overlap [ADJG19]. -safety [SD16b]. -Tier [WZK^{+}19].

/multi [Taf13]. /multi-threaded [Taf13].

’12 [Hol12]. 12th [Fox17a].

2 [HD17]. 2002 [FLL^{+}13]. 2003 [BCR13]. 2008 [HGCA11]. 2012 [HTW14, Hol12]. 2015 [LSBV17]. 27th [KP15].

5 [Dan18, KHR11].

6 [Jen12].

7 [Ano15, EV13, J^{+}12]. 75 [HWM11].

8 [BKP16, CWGA17, LYBB14, SAdB^{+}16, UFM15].

9 [Bla18, LSBV17]. 938 [Gun14]. 978 [Ano15]. 978-1-4493-1103-2 [Bro12]. 978-1-4919-4946-7 [Ano15]. 978-1-68050-288-6 [Ano18]. 9th [Gve13].

Abbreviated [SRTR17]. ABS [SAdB^{+}16]. absence [AGH^{+}17]. Abstract [AGR12, BDT10, DLR16, KPP12, XMA^{+}14, DLM10, DLR14, FSC^{+}13, KMMV14, NSDD17, SSK13]. Abstraction [BW12, Bro12, GY16, SKKR11, PL12, ZMG^{+}14, ZFK^{+}16]. Abstractions [NYCS12, RFBJ14, UR15, SPP^{+}10]. accelerated [PQTGS17]. Accelerating [KMZN16, ZLBF14, Cha18]. Accelerator [MAK19, OIA^{+}13]. accelerators [PWA13]. Access [CSGT17, HBT12, TT11, TNTN12, BP19, BB17, KT14, MHM10, RHN^{+}13, XHH12]. Accessibility [STST12, VBMDP16]. Acculock [XXZ13]. accuracy [MDHS10]. Accurate [Jaf13, RRB17, XXCL19, ZBB15, XXZ13]. ACDC [AHK^{+}15]. ACDC-JS [AHK^{+}15]. across [DD13, DFR13, HLSK13]. action [KB17, UFM15]. Actions [RK19, MTSH16]. Active [BSAL18, YMHB19]. Actor [RCB17]. actors [PGA18, Sub11]. Ada [Car11, Sch10a, WCB16]. adaptable [ADI13]. adaptation [VBAM10a]. Adapter [SK12]. Adaptive [AFG^{+}11, IHWN12, NFV15, RXK^{+}17, YPMM12, CL17, PKO^{+}15, PDPM^{+}16, SEPV19, VBAM10b]. add [DLM10]. adding [MZC10a]. Addition [RK19]. addressing [GD10, VBMDP16]. Adequacy [PSJ18]. Adequate [GGZ^{+}15]. ADiJaC [SD16a]. Admitted [YXS^{+}19]. Adoption [PBMH13, PGA18]. Adriaan [Ngo12]. Advanced [Hor11, VBAM10a, dJM18, Jen12]. Advances [FHP^{+}12]. Adversarial [FF10]. Aegis [Nil12a]. Æminium [SNS^{+}14]. affects [LO15]. affordable [BM14]. Agent [AFGG11, PE11, RVP11, Den18].

Agent-Based [PE11]. agent-oriented [RVP11]. aggregates [BCR11]. Agility [Bro12]. Ahead [BLH12, JMB12, PKPM19]. Ahead-of-Time [JMB12, PKPM19]. Aided [KP15]. air [OD18, PPS16]. Ajax [MvDL12]. Ajax-Based [MvDL12]. algebraic [Lei17]. algebras [IvdS16, ZCdSOvdS15]. Algorithm [JJCO19, YCYC12, ZW13, MT13, MLM17, Por18, Gun14]. Algorithmic [FHP^{+}12]. Algorithms [BF18, GT10b, Gra15]. Aliasing [NS12]. Alignment [NBB18]. alike [DAA13]. All-in-one [SV18]. Allocation [CPST14, WZK^{+}19, YPMM12, CPST15, OOK^{+}10]. allocation-site-based [CPST15]. Almost [NWB^{+}15, SC16]. alternatives [SHU16]. Alting [WBM^{+}10]. always [AJL16]. Analyses [Kri12, TN19, CDBD18, HB13, KMZN16, PMP^{+}16, ZMG^{+}14]. Analysis [ADJG19, AGM^{+}17, Bul18, CPV15, Hol12, JJCO19, KCD12, LHR19, MvDL12, NS12, RDCP12, RRB19, RPP19, SGD15, SW12, SDC^{+}12, SLES15, SLE^{+}17, SR17, VP16, ZKB^{+}16, AM14, Bra14, CFH^{+}13, CDMR19, DHS15, GYB^{+}11, HCN14, HWLM11, KSW^{+}14, KT14, KvGS^{+}14, KPP^{+}18, KRR19, LSBV16, LSBV17, LT14, LTMS18, LX19, MTL15, MKZ^{+}14, MCC17, MB12, NSDD17, NS13, PIR17, PLR18, Puf13, RLBV10, RRB17, SPPH10, SMSB11, SBK13, SP10b, SNCM19, TLX17, TWX^{+}10, TLMM13, TL17, TPG15, WA19, ZMNY14, ZWSS15, CH17]. Analytics [BBB^{+}17, KB17, STCG13]. Analyzer [Spo16, Fer13, GN16, SMP10]. Analyzing [PLL^{+}18, ZDK^{+}19, BTR^{+}13, PSNS14]. Android [CNS13, MMP^{+}12, STY^{+}14, THC^{+}14, ZHL^{+}12, ZKB^{+}16, vdMvdMV12]. AngularJS [RVT18]. Ann [CSdL16]. annotated [TJLL18]. annotation [CV14, KATS12]. annotation-based [KATS12]. annotations [CSdL16, GBS14, MGS19]. announcement [SPAK10]. anomalies [FRM^{+}15]. answering [KM10]. any [FIF^{+}15]. anytime [STCG13]. anywhere [STCG13]. AOP [WAB^{+}11]. AOT [Ser18, WKJ17]. Apache [CJ17, FRM^{+}15]. apart [LBF12]. API [FH16, MPM^{+}15, PTRV18, TWNH12, WA19, YKSL17]. APIs [HBS16, RDP16, Sam12, SRB18, VM10]. app [Ngo12, Sta10]. Apple [Ano13]. Application [BH12, CCA^{+}12, KF11, KB11, LZ12, RDCP12, RLMM15, SWF12, WZK^{+}19, AYZI10, AAB^{+}10, AÖ11, CNRG19, Del13, FRGPLF^{+}12, HWLM11, LBF12, OUY^{+}13, SE12, WA19, WSH^{+}19, WAB^{+}11, XHH12, HD17]. Application-Aware [LZ12]. Application-Replay [BH12]. Applications [GMPS12, GD12, MAHK16, MGI14, MvDL12, MMP15, NKH16, NWB^{+}15, OwKPM15, RPP19, SLES15, VP16, WBA^{+}11, AMT17, AST^{+}16, AC16, AMWW15, ADI13, ABFM12, DSEE13, BOF17, BFS^{+}18, BBXC13, EABVGV14, GMC^{+}13, HLO15, JH11, MTL15, MZC10a, MZC10b, PLR14, PKC^{+}13, RHSD15, R^{+}13, RVP11, RW17, Ryu16, Sch10b, SAdB^{+}16, SGV12, SPP^{+}10, TWX^{+}10, VB18, WHIN11, XGD^{+}19, vdMvdMV12]. applying [CMM17]. Approach [BDT10, CSF^{+}16, DLPT14, KKW14, LYM^{+}18, STST12, ADI13, FGB^{+}19, CHM13, CSKB12, DHM^{+}12, HLO15, HdM17, J^{+}12, MZC10a, MvH15, PSW11, RVP11, RO12, SEPV19, SNS^{+}14, YSCX17]. approachable [WHV^{+}13]. approaches [CDMR19, GD10, MD15, SS14]. approximate [CNS13]. Approximation [RvB14]. Approximations [SS12]. apps [BM18, CNS13, MMP^{+}12, Ngo12, Sta10]. Architectural [CSGT17, KKK^{+}17, Cha18]. Architecture [GMPS12, Wan11, XXCL19, AMWW15, Del13, Gon11]. Architectures [KKK^{+}17, RKN^{+}18, ABCR10, Hos12, MS10, ZP14]. arena [TRE^{+}13]. arithmetic [TGZ17]. Arm [DiP18b]. Arquillian [Ame13]. array [SV15b]. arrays [FBH17, SBF^{+}10]. arrows [FZ17]. art [Lew13]. ASM [AGR17]. Aspect [ABMV12, BH10, VBAM10b, VBMA11, WBA^{+}11]. Aspect-Oriented [ABMV12, BH10, VBAM10b, WBA^{+}11]. Aspectizing [TNTN12]. AspectJ [AC10]. aspects [LVG10]. Assertion [MM12]. Assertion-Based [MM12]. Assertional [LL15]. assertions [VYY10]. Assessing [GTSS11, PSJ18, VBZ^{+}18, JACS10]. assessment [IS18]. assignment [KT15]. AST [DRN14, HWW^{+}15, ZLBF14]. asymmetric [CBGM12]. asymptotic [ODL15]. Asynchronous [KW11, SK12, WK12, AZMT18, FZ17, KW10, LML17]. atomic [WAB^{+}11]. Atomicity [GGRSY17, JLP^{+}14, BHSB14, BNS12, GGRSY15, UMP10]. atomics [PPS16]. Attack [BH12]. Attacks [MSSK16, SBE^{+}19, VS11]. attribute [SHU16]. attributes [GD10]. augmentation [DAA13]. Augmenting [ZYY^{+}19]. authentication [XHH12]. authorship [FMS^{+}11]. auto [SKBL11]. auto-tuning [SKBL11]. AutoFix [YSCX17]. automata [LKP19, TLX17, ZWZ^{+}14]. Automated [BH17, BSOG12, BMOG12, MS14, RGEV11, SDM12, TJLL18, UPR^{+}18, ASdMGM14, MRMV12, YSCX17, ZFK^{+}16]. Automatic [GGRSY14, GGRSY15, GGRSY17, IS18, KKW11, LXP18, MDS^{+}17, MM16, PQD12, PBM^{+}19, RK19, SZ11, SD16a, SJPS10, SS16, WM10, XMD^{+}17, ZLNP18, ABK^{+}16, FM13, PG12]. automatically [TB14, VB18]. Automating [YXS^{+}19]. Autonomic [DLPT14, SEK^{+}19]. Autonomous [GMPS12]. average [LDL14]. avoid [XR10]. Avoiding [FRC^{+}17, ZBB17, OD18]. avoids [PPS16]. Aware [JYKS12, LZ12, BBXC13, CL17, EQT10, SSB^{+}14a, SZZ^{+}19, SGV12]. awareness [VGS14]. axiomatic [TVD10].

B [DLZ^{+}13]. back [Car11]. Background [PWSG17, PWSG19]. Backstage [PS11]. Bad [dGRdB^{+}15]. baggage [KFB^{+}12]. balances [FMBH15]. balancing [PDPM^{+}16]. Ball [DD13]. Bar [WCG^{+}18]. Barrier [CHMY19, CHMY15, VB14a]. barriers [HJH10, WBM^{+}10]. Based [AFGG11, DLR16, GM12, GGZ^{+}15, GGC18, LTD^{+}12, MvDL12, MM12, PTML11, PiLCH11, PE11, RBL12, RT14, SGD15, SLS^{+}12, ST15, SWF12, YPMM12, AYZI10, AZLY18, AST^{+}16, ADI13, BBF^{+}10, BBP13, BB17, BL15, CDTM10, CNRG19, CSKB12, CJ17, CJ19, CPST14, CPST15, EKUR10, GT10a, GMC^{+}13, GGC19, HWM14, HWI^{+}12, HOKO14, HWLM11, IHWN12, IRJ^{+}12, JEC^{+}12, JMO14, KATS12, KS13, KRCH14, KvRHA14, KS14, Lon10a, Lon10b, MCC17, MB12, MCY^{+}10, MTSH16, Ott18, PDPM^{+}16, PSW11, SZ11, SBK13, SMP10, SPY^{+}16, SV17, SNS^{+}14, UIY10, UPR^{+}18, VSG17, XHH12, YP10, YKA^{+}19, ZYZ^{+}12, ZYY^{+}19]. Basic [NBB18, CZ14]. basic-block [CZ14]. basics [Zak12]. basierte [Ric14]. Battery [ST15]. battlefield [WT10]. Bayesian [BSA14, RKHN18]. Bazel [McN19]. BeagleBone [Ric14]. before [TD15]. Beginning [McN19]. begone [MRMV12]. Behavior [Sun18, LWB^{+}15, RLBV10, TABS12, WXR16]. Behavioral [LN15, AMWW15]. behaviors [PCL14]. behaviour [SMS^{+}12]. Beliefs [BA17]. Ben [Teo12]. Benchmark [GBC12, SMSB11]. Benchmarking [CKS18, AHK^{+}15, HCLH18, MDM17]. benchmarks [KHM^{+}11, RGEV11]. benefit [HH13]. best [Sch13]. Better [Bro12, TD15]. Between [ADJG19, PVB17, ZLHD15, BKP16, CMM17, CSKB12, CSF^{+}16, LSBV16, LSBV17, RDP16, SH12]. beyond [Mor18]. Big [BF18, GTS^{+}15, NWB^{+}15, NFN^{+}18, RVK15, BOF17, BBXC13, RVK19, SSG^{+}14, WR10, XGD^{+}19]. billions [DRN14]. Binary [WWG^{+}18, XXCL19]. bindings [VGRS16]. bird [Guy14]. Birthmark [PiLCH11]. Bitcoin [TD17]. BIXSAN [VS11]. Blame [KT15]. BLeak [VB18]. Bloat [MSS10, NWB^{+}18, XMA^{+}14, BRGG12, BBXC13, XR10]. bloat-aware [BBXC13]. block [CZ14, KBL14]. block-level [KBL14]. blocking [DW10]. Blockly [AMWW15]. Blueshell [PWA13]. Board [GLGA19]. boilerplate [ZCdSOvdS15]. Book [Ano15, Ano18, Bro12, Del13, Gve13, Kie13, Ngo12, Teo12, Teo13]. Bookshelf [Ano18]. Boosting [ASV^{+}16, AC16]. Bootstrapping [CBLFD12]. Bottle [DSEE13]. bottlenecks [DSEE13]. bottom [ZMNY14]. bottom-up [ZMNY14]. boundary [RDP16]. Bounded [NWB^{+}15, GMT14]. Bounds [SW12, GvRN^{+}11]. boxes [BDGS13]. Brain [VBZ^{+}18]. breaking [VB14a]. Breakpoint [ZW13]. breakpoints [PS12]. Brewing [WZL^{+}18]. Bridging [PVB17]. Bringing [CV14, HRS^{+}17, STS^{+}13]. Broken [dGRdB^{+}15, AZMT18]. Browser [MSSK16, PVB17, FIF^{+}15, VS11, VB14a, WGW^{+}11, YK14]. Browsers [HLSK13]. Browsing [LYM^{+}18]. Browsix [PVB17]. BUbiNG [BMSV18]. Budget [GM12]. buffered [DLZ^{+}13]. buffers [Gun14]. Bug [RPP19, LWH^{+}10]. Bugs [OBPM17, XMD^{+}17, ECS15, MDS^{+}17, ODL15, Ryu16]. Build [BMDK15, BNE16, ELW15, MAH12, WSH^{+}19]. Building [McN19, Sta10, SS19, HWW^{+}15, Ngo12]. built [DTM^{+}18]. built-in [DTM^{+}18]. Business [CCA^{+}12]. Bytecode [BDT10, BSOG12, FHSR12, NS12, RDCP12, Rey13, SEK^{+}19, AdCGGH16, CZ14, CNRG19, DLM10, SP10b, SMP10, VB14b].

C [BB12, CDG^{+}17, GBC12, KB11, LSBV16, LSBV17, NED^{+}13, SRTR17, Sta10, YSCX17, Zak18, ZWSS15]. C# [SSK13]. C/C [BB12, NED^{+}13]. CA [KP15]. cache [IN12, ZP14]. caches [NGB16]. calculations [VSG17]. Calculi [FFF17]. calculus [AH10, PS10a]. Call [FGR12, PULO16, ZWZ^{+}14, Xue12, SSB^{+}14a]. Call-site [SSB^{+}14a]. calling [HB13, SSB^{+}14a, ZWZ^{+}14]. Calls [SW12, SS16]. came [Car11]. can [TPG15]. can’t [WA19]. capabilities [Ame13]. capability [RDF15]. capo [SMSB11]. capturing [BKC^{+}13]. Card [GMPS12, BL15, ABFM12, MLM17, MLM19, dCMMN12]. Cards [BH12, GMNC13, GMPS12]. care [EKUR10]. Caring [DAA13]. carry [Ame13]. Cartesian [SD16b]. Case [LMZP19, ZMM^{+}16, dGRdB^{+}15, AMWW15, HNTL12, JK11, MT13, SPPH10, Vit14]. Cassandra [FRM^{+}15]. cast [MHN19]. Casting [MHN19]. casts [SH12]. categorising [CMM17]. Catena [TD17]. Causal [MRF18]. Causes [OBPM17, FRM^{+}15]. CAV [KP15]. Cay [Gve13]. CC [LSBV16, LSBV17]. CCA [FLZ^{+}18, ZXL16]. Center [Hol12]. centric [DHM^{+}12, FOPZ14]. CERT [LMS^{+}12]. chain [KSR14]. Challenges [GM12, SWMV17, Sie17, SR17, AACR18]. Change [YXS^{+}19, YQTR15, MPR12]. Change-Level [YXS^{+}19]. Changes [MvDL12, CJ19, PTRV18]. Changing [SSG^{+}14]. Channel [Bul18]. channels [AGH^{+}17, LS11]. characteristics [ABC18]. Characterizing [CJ17]. check [CS12, GvRN^{+}11]. Checking [BNE16, CSF^{+}16, Cho14, FSK12, JC10, JYKS12, ABFM12, BHSB14, BNS12, CVG^{+}17, DLM10, FLL^{+}13, HMDE12, KATS12, KvRHA14, LT11, RR14, RAS16, RDF15, TVD10, VYY10]. checkpointing [SGV12]. checkpointing-enabled [SGV12]. Checks [FMBH15]. CHERI [CDG^{+}17]. Children [GLGA19]. chip [PS10b, Puf13, RS12, SPS17]. chip-multiprocessor [PS10b]. chip-multiprocessors [RS12]. Choice [JCMM19, WBM^{+}10]. CICS [R^{+}13]. CIL [BBF^{+}10]. circular [Gun14, SZ10]. Circus [ZLCW14, MCW19]. City [Hol12]. Class [BS13, CSF^{+}16, NCS10, CSKB12, HC10, MHM10, SC16, SM12, TSD^{+}12]. Classes [And14, SVB^{+}17, WT11, CZ14, CS12, SZ10, TSD^{+}12, VBDPM16]. Classfiles [SD16a]. Classification [PBM^{+}19, SS14]. Classifiers [BSA14]. Classifying [MHM10, PBB19]. Classless [WZdSOS17]. clicker [HA13]. Client [MS14, OBPM17, CH17, GGC19, KRH16]. client-server [GGC19]. Client-Side [OBPM17, KRH16]. Client-State [MS14]. clients [SRB18]. Clojure [ECG12, FH11, VS10]. Cloned [SSL18]. Closing [ZLHD15]. Closures [BO11, BO12, BO13]. Cloud [VDV17, WZK^{+}19, BFS^{+}18, GGC18, LZYP16, TLMM13]. cloud-based [GGC18]. clustered [PDPM^{+}16]. clustering [MKK^{+}12, MKK^{+}13]. clusters [TRTD11]. Cocoa [Sta10]. Code [ADJG19, BH17, BNE16, CJ19, HC11, MSS19, MM16, PKPM19, RVK15, RLMM15, SRTR17, SVB^{+}17, SV15a, SED14, WWG^{+}18, XXCL19, AGR17, AK13, CCFB15, DRN14, FLZ^{+}18, FH16, FMS^{+}11, GMNC13, GGC19, IS18, LVG10, MKK^{+}12, MKK^{+}13, NG13, OJ12, PTRV18, PBB19, PMP^{+}16, PSW11, RFRS14, RBV16, RVK19, RO12, SSK13, Tai13, UTO13, VSG17, WKJ17, WGF11, WBA^{+}11, WAB^{+}11, WWS13, ZHL^{+}12, ZXL16, ZWSS15]. Code-Issue-Introducing [CJ19]. coding [LMS^{+}12, LMS^{+}13]. Coefficient [ADJG19]. Coffin [Teo12]. coherent [ZP14]. Cohesion [RC17]. Cold [BZD17, WGF11]. Collect [JCMM19]. collected [AGGZ10]. collecting [AHK^{+}11]. Collection [ASV^{+}16, BF18, GM12, MAK19, QSaS^{+}16, ST15, URJ18, ASME18, AFGM15, BP10, BOF17, KPHV11, KBL14, NGB16, ODL15, PZM^{+}10, PDPM^{+}16, SP10a, SBM14, Sie10, SJBL10, SKBL11, UIY10, UJR14]. Collections [GS12, DTM^{+}18, Lon10a, Lon10b, PL12, SV15b, SV17]. collectives [RTET15, TRTD11]. Collector [BH12, GTS^{+}15, BCR13, BVGV14b, Puf13]. Collectoren [Sch13]. collectors [GTSS11, Sch13, XGD^{+}19]. coloring [SS10]. Colt [BKP16, WN10]. CoMA [AGR12]. Combating [NWB^{+}18]. Combination [BSA14]. Combinatorial [YHY13]. combinators [MHBO13]. Combining [BDGS13, MSS19, MGI17]. commensal [BRWA14]. comments [PBB19, ZYY^{+}19]. Commercial [ZMM^{+}16]. commodity [BK14]. Common [PiLCH11]. Communication [JQJ^{+}16, RTE^{+}13, SK12, BJBK12, ETR^{+}15, TTD^{+}11]. communications [ETTD12, RTET15, TTD12]. Communities [ZMM^{+}16]. COMP [CKS18]. Compact [HWM10, HWM11, JJL17]. Comparative [KB11, CDMR19, KFBK^{+}15, SSL18]. comparing [MD15]. Comparison [BKP16, ADI13, BJBK12, HH13, KvRHA14, SMS^{+}12]. Comparisons [GGZ^{+}15]. Compartmental [WGW^{+}11]. compatibility [DJB16, OIA^{+}13]. compatible [ABCR10, Hor12]. Competition [CKS18]. Compilation [DLR16, PKPM19, CGJ^{+}16, CMS^{+}12, DLR14, FSC^{+}13, IHWN12, JLP^{+}14, JK13, JMO14, KS13, KHL^{+}13, Lei17, MD15, MGI17, Ser18, ZBB15]. compiled [NED^{+}13, RO12, TMVB13]. Compiler [JMB12, Loc18, NKH16, NWB^{+}15, BBF^{+}10, BRWA14, CIAD13, Cle16, HWM14, IHWN12, KMLS15, KS14, KC12, LSWM16, MDM17, Ott18, Rub14, TTS^{+}10, TWSC10, VB14b, ZYZ^{+}12]. compiler-compiler [KS14]. compiler-runtime [TWSC10]. compilers [Hos12, LMK16, RSB^{+}14].

Compiling [Fee16, Hos12].

complementation [BS13]. Complete [BO13, BR15, JC10, NNT^{+}19, Sch14, Gri17, PSR15, RGM13, RRB17]. completeness [KBPS17]. completing [BS13]. completion [FH16]. Complexity [SSH17]. Compliance [GD12]. compliant [MZC10a]. component [AST^{+}16, CSKB12, GT10a]. component-based [AST^{+}16, GT10a]. components [BMSZ17, FOPZ14, KS14]. Composable [SS10]. Composing [EABVGV14]. Composition [SK12, AGH^{+}17, AH10, SZ10, VM15]. compositional [BGOS18]. Comprehension [BGK17]. Comprehensive [STST12, VBMA11, ZKB^{+}16, MKZ^{+}14]. Compressing [Gun14]. Computation [BW12, LYM^{+}18, ZHL^{+}12]. Computation-Intensive [LYM^{+}18]. computational [Bra14, SSG^{+}14, VF10]. computations [KFBK^{+}15, TLMM13]. Computer [HWM11, OAC18, DNB^{+}12, KP15]. Computing [Hol12, MPR12, NBB18, PWSG17, PWSG19, SHU16, TWNH12, WN10, AdSCdR^{+}19, Dan18, HCLH18, LZYP16, Rub14, TTD^{+}11, VF10, TRE^{+}13]. con [SMSB11]. conceptual [Tai13]. Concurrency [BG17, Bro12, SWF12, BVGVEA11a, CHM13, DMS11, HAW13, KHL^{+}17, PPS16, Sub11, TD15, UR15]. Concurrent [MSM^{+}16, PS12, Sie10, BP19, BMSZ17, EP14, Gra15, HJH10, KBL14, MSM^{+}10, OW16, PTF^{+}15, RVP11, STR16, SNS^{+}14, WLL19, YS10, YKA^{+}19]. concurrent-by-default [SNS^{+}14]. Conditional [XMD^{+}17, SS16]. Conference [DDDF17, Hol12, KP15, LMK16, PDPM^{+}16]. Configurations [PSJ18]. conflict [ABC18]. Conformance [AGR12, SKR17]. Confused [BH12]. Connecting [NFN^{+}18]. conquer [SBF^{+}10]. Consequences [OBPM17]. conservative [SBM14]. Consistency [CSF^{+}16, CS12, DNB^{+}12, FRM^{+}15, ZBB17]. consistent [BCR13]. constrained [KSR14]. constraint [FMBH15, SHU16]. Constraints [SGD15, LSSD14]. construction [CIAD13, RGEV11]. constructors [MME14]. constructs [PCL14, PTF^{+}15]. consumers [DAA13]. Consumption [MV16]. container [XR13]. containers [XR10]. Context [HWM13, MM16, TL17, HB13, IvdS16, LTMS18, LX19, SSB^{+}14a, ZYY^{+}19]. Context-sensitive [HWM13]. Contextual [MSSK16]. Continuous [Teo12]. Continuously [DTLM14]. Contracts [YQTR15, HBT12, KT15, KKW11]. Control [FGR12, FHSR12, TT11, TNTN12, AdCGGH16, BNP^{+}18, BL15, FWDL15, LSWM16, RHN^{+}13, STS^{+}13, TABS12, WLL19, XHH12]. controlling [BKC^{+}13, YDFF15]. Convention [Hol12]. conversions [CMM17]. Converter [YWW^{+}18]. Cooperative [YDFF15, HdM17]. Coordinating [MAHK16]. coordination [BMSZ17]. copy [FBH17]. copyrightable [Sam12]. Core [Hor11, HC13, RDCP12, RTE^{+}13, MS10, PLL^{+}18, TRTD11, Gve13]. cores [GTSS11, SKBL11]. Cornell [Gve13]. corpus [HCN14, LSBV16, LSBV17, TMVB13]. correct [AdCGGH16, AJL16, DJLP10, PS10a]. Correctness [LL15, BENS12, Cho14]. Correlation [SDC^{+}12, XHH12]. Corrigendum [LSBV17]. Cost [MSS19]. costs [OD18]. counter [LSSD14]. counters [IN12]. Counting [Bul18]. Course [Wan11, Zak12]. Coverage [CSS^{+}16, GGZ^{+}15, MSS19, RGB18]. Coverage-Based [GGZ^{+}15]. Coverage-directed [CSS^{+}16]. CPS [PDDD17]. CPU [PKO^{+}15]. Crawling [BMSV18, MvDL12]. Creating [YMHB19, HC10, VBAM10b]. Creation [SK12]. crisis [AT16]. Critical [HL13, MCW19, WK12, WCB16, ZLCW14, AGR17, DTLM14, GMC^{+}13, NM10a, Nil12b, RS12, SDH^{+}17, CWW13, LWC17]. Cross [GSS^{+}18, MDM17, OTR^{+}18, WBHN18, XXCL19, AMWW15, BKC^{+}13, GSS^{+}16, KMZN16]. Cross-Architecture [XXCL19]. cross-cutting [AMWW15]. Cross-Language [GSS^{+}18, MDM17, GSS^{+}16]. Cross-Layer [OTR^{+}18]. Cross-OS [XXCL19]. Cross-Platform [WBHN18]. cross-program [KMZN16]. cross-thread [BKC^{+}13]. Crowdsourcing [BH17]. CrowdSummarizer [BH17]. crypto [PTRV18]. Cryptography [GPT12]. CSM [dARPH^{+}19]. CSS [Ano15, HLO15, Sta10]. Curve [GPT12]. customizations [LVG10]. customized [HB13]. cutting [AMWW15]. Cyclic [BMOG12, RS12].

D

[DiP18b, FLZ^{+}18, GBC12, JEC^{+}12, ZXL16]. DAA [DR10]. Dalvik [AFGM15]. dark [MHN19]. Data [Bra14, BMOG12, BA17, BF18, GM12, GTS^{+}15, GT10b, JJCO19, NKH16, NWB^{+}15, NFN^{+}18, NWB^{+}18, TAF^{+}18, YWW^{+}18, ZLNP18, dMRH12, AFGM15, BK14, BB17, BOF17, BBXC13, BJBK12, CDTM10, CRP^{+}10, DFR13, DHM^{+}12, EKUR10, FOPZ14, KB17, LDL14, MRA^{+}17, MTSH16, NL14, SAdB^{+}16, SSG^{+}14, SGG^{+}17, UMP10, WKJ17, WCG14, XXZ13, XMA^{+}10, XGD^{+}19, ZIvdS17]. data-centric [DHM^{+}12, FOPZ14]. Data-Driven [JJCO19]. Data-Intensive [NWB^{+}18]. Data-Parallel [NKH16, CRP^{+}10]. database [Dei10, EKUR10, TABS12]. databases [EKUR10, MLGA11]. Dataflow [BR12]. Datalog [ZMG^{+}14]. dataset [MDS^{+}17]. David [Kie13]. Days [Sev12b]. DBT [KS13]. dead [SK13]. Deadlock [CHMY19, CHMY15, SR14a, SR14b]. deallocated [AFGM15]. Dean [Bro12]. Debt [YXS^{+}19]. debugging [ASdMGM14, BM14, KS14, TB14, VB18, ZFK^{+}16]. December [LSBV17]. Deciding [SGD15]. decision [RBV16]. Declarative [DRN14, RSI12, FOPZ14, WCST19, MME^{+}10]. Decomposition [AGH^{+}17, PLL^{+}18]. deconstructing [ACS^{+}14]. decoupled [LPA13]. deduplication [HOKO14]. Default [BG17, SNS^{+}14]. defects4j [MDS^{+}17]. defined [FMS^{+}11]. Definite [NS12]. Definition [SSB14b, AK13, SSB01]. Definitive [Oak14]. delegation [GBS13]. delimited [PDDD17]. DelphJ [GBS13]. demand [FWDL15, SNCM19, ZHL^{+}12]. demand-driven [FWDL15, SNCM19]. DemoMatch [YKSL17]. demonstrations [YKSL17]. Deoptimization [KRCH14]. depend [LCW18]. dependability [GD10]. Dependence [PDDD17, JWMC15]. Dependence-driven [PDDD17]. dependences [BKC^{+}13, WLL19]. dependencies [ELW15]. Dependency [ZBvdB^{+}19]. Dependent [CHJ12, LE16]. deploying [R^{+}13]. deprecation [SRB18]. depth [HTBB16, Rau14]. Design [AC16, CNRG19, ETTD12, MLGA11, Puf13, RTE^{+}13, SW12, TRTD11, TKL^{+}15, VGRS16, YCYC12, BBXC13, CSdL16, GSD^{+}15, IRJ^{+}12, Lon10a, Lon10b, OA17, SAdB^{+}16, SMSB11, VM10, Xue12]. Designing [Sev12b, KHR11]. Desktop [GS11]. destructive [FF10]. detect [GGC19]. Detecting [BK12, HLO15, PiLCH11, XR10, FF10]. Detection [BH10, BSOG12, KCD12, MS14, RD15, XMA^{+}14, AMT17, BGOS18, CSK17, LMK16, LS11, ODL15, PG12, RDF15, RGB18, RW17, SR14a, SR14b, SS14, WCG14, XXZ13, XR13]. detector [WFF18]. detectors [LWH^{+}10]. Determinacy [AM14]. Determination [YXS^{+}19]. deterministic [DNB^{+}12, MvH15]. develop [WA19]. developer [EV13, Top11, ZZK13]. Developers [Bro12, BMR14, DJB16, HH13, Wam11]. Developing [FGB^{+}19, R^{+}13]. Development [ABK^{+}16, AYZI10, MT13, PBM^{+}19, AGR17, BM18, FRGPLF^{+}12, GT10a, PSW11, SKR17, SH12, WBA^{+}11, ZDS14]. Device [TTD^{+}11, XHH12]. Devices [GPT12, JQJ^{+}16, MV16, ETR^{+}15, Xue12]. DFC [BR12]. diagnosis [RW17]. DiAl [STCG13]. dialects [BIvdS17]. difference [PS11]. differential [CSS^{+}16]. Differentiation [FHP^{+}12, PQD12, SD16a]. digital [JMO14]. dimensional [TGZ17]. Directed [STR16, CSS^{+}16, EP14, Lei17, NG13, NED^{+}13, WM10]. directives [VGS14]. Discovering [Sev12a]. discovery [YKSL17]. discrete [DDDF17]. Disease [PE11]. Disjunctive [JJCO19]. Dissimilar [Has12]. Distance [ZW13]. distributable [CRAJ10]. Distributed [BVEAGVA10, CWGA17, LTD^{+}12, LM15, MAHK16, MRF18, NFN^{+}18, PE11, YMHB19, AdSCdR^{+}19, BVGVEA10, BVGVEA11b, BVGV14b, CDBD18, CRAJ10, EABVGV14, STCG13, SS19]. distributing [TGZ17]. divide [SBF^{+}10]. Do [HH13, LMZP19, Han15]. Does [BRGG12, Rub14]. DOJ [hEYJD12]. DOM [GGC18]. DOM-Based [GGC18]. Domain [KSPK12, CSdL16, EEK^{+}13, HWW^{+}15, PIR17]. domain-specific [CSdL16, EEK^{+}13, HWW^{+}15]. dominance [CPST14]. Doppio [VB14a]. DoubleChecker [BHSB14]. down [Ker15, ZMNY14]. DRAM [OTR^{+}18]. drf [MSM^{+}16]. DRFX [MSM^{+}10, SMN^{+}12]. Driven [CCA^{+}12, JJCO19, YPMM12, BM18, FGB^{+}19, CHM13, FWDL15, HZZK19, LKP19, MTL15, PDDD17, SR14b, SNCM19]. drug [EKUR10]. DSL [KARO12]. DSLs [KHR11, RO12, SC16]. DSSAT [dARPH^{+}19]. DSSAT-CSM [dARPH^{+}19]. DSU [PVH14]. Dual [AD16]. Dual-Pivot [AD16]. Dynamic [AGM^{+}17, ABMV12, ASF17, BFS^{+}18, CHMY15, CHMY19, LMZP19, MRF18, MvDL12, PTHH14, RDF15, SMP19, WWG^{+}18, XMA^{+}14, ZKB^{+}16, AF12, BDB11, BK14, BCD13, BOF17, CSV15, CDBD18, CPST15, DTM^{+}18, ELW15, GYB^{+}11, HB13, KRCH14, KRR^{+}14, KT14, LWH^{+}10, LVG10, MKZ^{+}14, Nil12b, NG12, NED^{+}13, RLBV10, RCR^{+}14, RRB17, SR14b, SPKT18, SJPS10, SH12, TPG15, VBAM10b, WXR16, WFF18, WBA^{+}11, WAB^{+}11, WWS13, WWH^{+}17, ZBB15]. dynamic-memory [GYB^{+}11]. Dynamically [WWG^{+}18, CZ14, CMS^{+}12, hEYJD12]. Dynamically-Generated [WWG^{+}18]. Dynamo [BDB11].

e-Science [SGV12]. ease [DRN14]. Easy [Jaf13, CRP^{+}10]. economic [CSV15]. economics [SJBL10]. Ecosystem [YMHB19]. Edition [Ano15, Gve13, LYBB14]. editor [EKR^{+}12]. Editorial [Fox17a]. Editorials [Fox17b, HTW14, RHT13]. EdSketch [HZZK19]. EDSLs [RDP16]. Educator [BA17]. EE [Jen12, MCC17]. Effect [BSO18, JK11, CCFB15]. Effective [BMR14, PTML11, RD15, CSdL16, KPP^{+}18, Kie13]. Effectively [UR15]. effects [FH16, HAW13, Lei17]. Efficiency [OTR^{+}18, SEPV19]. Efficient [DVL13, GPT12, HWM11, HB13, KT14, KW10, OOK^{+}10, RSF^{+}15, RFBJ14, SYZZ^{+}14, SMN^{+}12, TLX17, TN19, TD17, AK13, BHSB14, CRP^{+}10, ETR12, HWM10, KKW11, MRA^{+}17, MSM^{+}10, Pos19, Sie17, SGV12, SWB^{+}15, SV15a, TRTD11, UMP10, VWJB10, XXZ13, ZDK^{+}19, SV18]. Efficiently [FBH17, BKC^{+}13, FOPZ14]. Einsatzszenarien [Sch13]. Einsteiger [Ric14]. Elektronik [Ric14]. Elektronik-Projekte [Ric14]. Elephant [RGM13]. Elimination [RKN^{+}18, GvRN^{+}11]. elision [NM10a]. Elliptic [GPT12]. Eloquent [Hav11]. emass [Por18]. Embedded [Fox17b, HTW14, JMB12, KARO12, Pau14, SLES15, SLE^{+}17, TKL^{+}15, VK12, Dei10, Fox17a, GMC^{+}13, HTLC10, KHR11, LMK16, LTK17, OIA^{+}13, RHT13, SC16, SDH^{+}17, SFR^{+}14, UIY10, Xue12, ZYZ^{+}12]. embedding [KMLS15, SC16]. emerging [CDMR19]. Empirical [LSBV16, LSBV17, SS13, WXR16, BJBK12, FH16, HH13, KPP^{+}18, MHN19, MHR^{+}12, NCS10, SH12, Tai13, VBDPM16, VBMDP16]. Employing [CC15]. Emscripten [Zak18]. emulated [THC^{+}14]. Emulation [XXCL19]. emulator [KS13]. Enabled [GPT12, DR10, ETR^{+}15, RBL12, SGV12]. encapsulation [DDM11]. End [GM12, DAA13]. End-to-End [GM12]. end-user [DAA13]. Energy [OTR^{+}18, CL17, PCL14]. energy-aware [CL17]. enforcement [IF16]. enforcing [JWMC15]. Engine [SMP19, MGI17, Ngo12, OUY^{+}13, Tar11, Ngo12]. Engineering [CCA^{+}12, GT10a, MLM17, MLM19, VF10]. engineers [Bra14]. engines [KRH16, SSG^{+}14]. enhanced [LMK16, WBA^{+}11]. enhancement [WCST19]. Enhancing [BDT10, BVGVEA13, DcSG12, HC10]. Ensuring [HDK^{+}11]. Enterprise [Ano14, AAB^{+}10]. entities [ETR12]. Entry [BK12]. enumeration [SSH17]. Environment [Köl10, PTML11, RK19, EKR^{+}12]. Environments [BF18, EABVGV14, GTL^{+}10, HOKO14, KF11, RDP16, RCB17, SGV12]. equality [GRF11]. Equilibrium [YMHB19]. Equivalence [BO12]. equivalent [TLX17]. equivocation [TD17]. ERAM [Sch10a].

Erratum [HWM11]. error [eBH11]. ES5 [DFHF15, Mor18]. ES6 [Mor18]. Escape [SLES15, SLE^{+}17]. Espresso [WZL^{+}18]. Essential [Ngo12]. estimation [LMK16]. etched [VSG17]. Ethereum [Dan17]. eval [Mil13, MRMV12]. Evaluating [BGK17, BLH12, MDHS10]. Evaluation [CSZ17, GBC12, JMB12, OCFLI14, TTS^{+}10, Wan11, CSK17, MRA^{+}17, MD15, WWH^{+}17, XGD^{+}19]. Evaluator [JB12]. Event [KW11, MV16, BBP13, KW10, MTL15, WK12, YP10]. event-based [BBP13, YP10]. event-driven [MTL15]. EventBreak [PSNS14]. ever [Gra15]. everyone [Hor12]. Evolution [CC15, GMPS12, Mei14, JK11, MAH12, NCS10, WBA^{+}11, WAB^{+}11, WWS13]. evolving [ZZK13]. Exact [ZW13]. Examples [BNP11, Del13]. Exception [LT14, ECS15, HWM14, LT11]. Exceptionization [YKM17]. Exceptions [ASF17, AdCGGH16, HdM17, SMN^{+}12, ZBB17]. Execution [Bul18, MSS19, NNTK17, NNT^{+}19, OwKPM15, SWMV17, BNP^{+}18, Cha18, HZZK19, JJL17, JhED11, LLL13, MMP^{+}12, RCB17, SPPH10]. execution-driven [HZZK19]. Executions [WCG^{+}18, ASdMGM14, PPS16, STR16]. executives [RS12]. Exemplar [ZW13]. exhaustive [DHS15]. exhibitionism [VBMDP16]. existential [AT16]. Exogenous [BMSZ17]. Experience [ABMV12, OW16, Sch10a, FGB^{+}19, CBLFD12, TRE^{+}13, WT10]. Experiment [BKP16, MDS^{+}17, HWLM11]. experimental [XGD^{+}19]. explicit [NGB16]. exploit [Ano13]. Exploitation [SSMGD10, HTBB16, MLM19]. Exploiting [NKH16, QSaS^{+}16]. exploration [FWDL15]. explorative [AHK^{+}15]. Exploratory [BKP16, ECS15]. explore [Dan18]. EXPLORER [FWDL15]. Exploring [JK13, JWMC15, SE12]. exposed [VBDPM16]. Express [JQJ^{+}16].

Expression [NS12, PIR17]. Expressions [NM10b, GK15, MKTD17]. expressive [VYY10]. Extended [DDDF17, FGR12, FLL^{+}13, JC10, LMK16, PDPM^{+}16]. Extending [AC10, BVGVEA11a, LPA13, PTHH14]. Extensible [ZIvdS17, ER14, KMLS15, MHBO13]. Extension [RSI12, WA19, LE16, MLGA11, PdMG12]. extensions [MPR12, Zha12]. Extensive [LMZP19, Wan11]. Extracting [CJ19, CCA^{+}12, KM10]. extraction [LKP19]. Extremal [LTD^{+}12]. Eye [OAC18, RLMM15, Guy14]. Eye-Tracking [OAC18, RLMM15].

F [GMT14, TTD12]. F-bounded [GMT14]. F-MPJ [TTD12]. FAA [Sch10a]. FACADE [NWB^{+}15]. face [XHH12]. Facebook [Ano13, HOSC16]. Facets [ASF17, AF12]. facilities [BVGVEAFG11]. Factors [PGA18]. FAD.js [BB17]. failboxes [ZBvdB^{+}19]. failing [STR16]. failures [CRAJ10]. false [HWI^{+}12]. familiarized [Ame13]. family [KHM^{+}11, KvRHA14]. family-based [KvRHA14]. Fast [CVG^{+}17, CSGT17, HyG12, SBM14, SLF14, YMHB19, Zak18, BB17, KMMV14, KCP^{+}17, LX19, MDM17, MHBO13, SV15b, WSH^{+}19]. Faster [BMDK15, JC10, AJL16]. Fault [RK19, RBL12]. Fault-Tolerance [RK19]. Faults [SRTR17, KPP^{+}18, ZZK13]. FC [YWW^{+}18]. Featherweight [RvB14]. feature [AH10, KvRHA14, OJ12, SS19]. feature-based [KvRHA14]. features [MKK^{+}12, MKK^{+}13]. Feedback [NED^{+}13, NG13, WM10]. Feedback-directed [NED^{+}13, NG13, WM10]. fields [PQTGS17]. FIFO [QSaS^{+}16]. filtering [HWI^{+}12]. find [Ryu16]. Finding [AZMT18, RPP19, XMA^{+}10]. Fine [BVGVEAFG11, DRN14]. fine-grained [DRN14]. Fingerprints [MSSK16]. Finite [BLH12, MB12]. Finite-State [BLH12]. first [SC16, TSD^{+}12]. first-class [SC16, TSD^{+}12]. fix [TPG15]. Fixing [SRTR17, LTZ14, YSCX17]. flexibility [SBF^{+}10]. Flexible [ES14, MSM^{+}16, PKC^{+}13, RHN^{+}13, BCD13, KHR11, Por18, ZW10]. Flint [LTZ14]. Floating [Jaf13, AJL16]. Floating-Point [Jaf13, AJL16]. Flow [ASF17, FHSR12, LMK16, SS12, AdCGGH16, AF12, ABFM12, BK14, BL15, FWDL15, HBS16, KHL^{+}13, LSWM16, PMTP12, STA18, YSCX17]. Flow-sensitive [LMK16, STA18]. FlumeJava [CRP^{+}10]. fly [UJR14, URJ18]. folding [CPST14]. Footprint [GS12, WHIN11]. Forecasting [CC15]. foreign [LWH^{+}10]. forge [Ler10]. fork [MZC10a]. fork/join [MZC10a]. form [GK15]. Formal [DLPT14, KR12, SW12, HdM17, PSR15, SZ11]. formalised [CWW13]. formalization [BP19]. Format [YWW^{+}18]. Forsaking [GBS13]. FORSETI [CSV15]. Forward [FOPZ14]. Foundation [CJ17]. Four [MSS10]. FPGA [OUY^{+}13]. fragmentation [PZM^{+}10]. fragmentation-tolerant [PZM^{+}10]. Fragments [PBM^{+}19, OA17]. frames [SJPS10]. Framework [CCA^{+}12, Den18, FFF17, LM15, PWSG17, PWSG19, RBL12, SEK^{+}19, TN19, Ame13, AC16, DDDF17, ER14, FRGPLF^{+}12, GGC19, JEC^{+}12, KMLS15, Lon10a, Lon10b, MT13, PGA18, PKO^{+}15, RR14, STY^{+}14, ZW10, ZDS14]. frameworks [PPMH15]. Francisco [KP15]. free [DTLM14, FC11, GK15, HHB^{+}14, NFV15]. free-form [GK15]. free-lunch [DTLM14]. frequency [ZWSS15]. Frequent [RC17]. Friendly [RBL12]. fringe [MB12, MB12]. Full [SRTR17, CDBD18, DRN14]. full-stack [CDBD18]. Full-Word [SRTR17]. Fully [FSC^{+}13, PG12, ZFK^{+}16].

Functional

[HOSC16, Wam11, Ame13, BVGVEA11b, NFV15, SV18, UFM15, Bro12]. functional-style [UFM15]. functions [LSBV16, LSBV17, SPKT18]. Fundamentals [HC13, Teo13, Gve13]. Fusing [MS13, ETR12, WM10]. fusion [KBPS17]. future [HCLH18, SS16]. fuzzer [Guo17]. Fuzzy [YPMM12]. Fuzzy-Rule-based [YPMM12].

Game [MT14, Wan11]. Games [GLGA19]. Gap [PVB17, ZLHD15]. Garbage [ASV^{+}16, BH12, BF18, GTS^{+}15, JCMM19, MAK19, QSaS^{+}16, Sch13, SKBL11, URJ18, ASME18, AGGZ10, AFGM15, BCR13, BP10, BVGV14b, BOF17, GTSS11, KPHV11, KBL14, NGB16, PZM^{+}10, PDPM^{+}16, Puf13, SP10a, SBM14, Sie10, SJBL10, UIY10, UJR14, XGD^{+}19]. garbage-collection [Sie10]. Gary [Gve13]. GC [NGB16, RGM13, SEPV19]. GC-Wise [SEPV19]. GEMs [BSMB16]. General [CHMY19, AdSCdR^{+}19, AZLY18, CHMY15, EKUR10]. general-purpose [AdSCdR^{+}19]. generalized [WT10]. generate [CS12]. Generated [WWG^{+}18, BM18]. Generating [HJS^{+}10, RDP16, GRF11, KS14, MHBO13, SSK13]. Generation [AGM^{+}17, BH17, YWW^{+}18, CRJ^{+}10, CMM^{+}10, PPMH15, Pha18, PSNS14, Rim12, RO12, SPKT18, UMP10, ZYY^{+}19]. generations [BOF17]. generators [SLF14]. generic [DDM11, Fer13, HH13, ZPL^{+}10, eBH11]. generics [AS14, Gri17, PBMH13]. Genetic [YCYC12, MT13]. Genotyping [YCYC12]. GeoGebra [ABK^{+}16]. geosciences [MCY^{+}10]. Geospatial [CH17]. German [Sch13]. get [Ame13]. Getaway [SLES15, SLE^{+}17]. Gets [BH12]. getters [Mil13]. Getting [GMT14]. Giga [DHS15]. Giga-scale [DHS15]. GitHub [LMZP19]. glimpse [SP16]. Global [PE11].

Global-Scale [PE11]. Globally [YMHB19]. Glotaran [SLS^{+}12]. go [LWB^{+}15, McN19]. Goldilocks [EQT10]. Good [dGRdB^{+}15]. Google [Ngo12, MGI17, Sam12]. GPGPU [PQTGS17]. GPGPU-accelerated [PQTGS17]. GPU [CNRG19, PKO^{+}15]. GPU-based [CNRG19]. GPUs [Hos12]. grade [CRJ^{+}10]. Gradual [RSF^{+}15, SFR^{+}14, TSD^{+}12, Sie17]. grained [DRN14]. grammars [GN16, SHU16]. Granularity [RRB19, CZ14, YKA^{+}19]. Graph [dMRH12, BS13]. Graphical [SLS^{+}12]. Graphics [Cec11, LLL13]. graphs [AdCGGH16, DSEE13, JWMC15, PULO16]. green [BRGG12]. Greenfoot [Köl10]. grid [SGV12, VWJB10, MZC10b]. Gridifying [MZC10b]. grounded [EV13]. Growing [EKR^{+}12]. growth [LDL14]. guarantees [JWMC15, ZHCB15]. GUI [CNS13, VGS14, WBA^{+}11]. GUI-awareness [VGS14]. Guide [Ame13, Oak14, Rau14, Teo13, Top11]. Guided [CNS13, DiP18b, MMP15, GY16, LTMS18, Ott18, PSNS14, RKHN18, SSH17]. Guidelines [GGZ^{+}15, HLSK13, LMS^{+}13].

Hack [Ott18]. handlers [BSO18]. Handling [KW11, ECS15, HWM14, KW10, WK12]. Hands [CSZ17, Teo13]. Hands-on [CSZ17, Teo13]. happened [Han15]. happens [TD15]. happens-before [TD15]. hard [LTK17, Puf13]. Hardware [MAK19, SKKR11, SPS17, CBGM12, IN12, SE12, ZDK^{+}19]. hardwired [OUY^{+}13]. harness [Kie13]. harvesting [AFGM15]. hash [SV15a, SV15b, SV18]. hash-array [SV15b]. hash-tries [SV18]. hashing [GRF11]. HDFS [IRJ^{+}12]. HDL [OUY^{+}13]. health [EKUR10]. heap [CSV15, LDL14, TLX17, Tar11, VYY10, YS10, BVGVEA10]. heap-manipulating [YS10]. Heaps [NFN^{+}18]. Helping [RT14]. Hera [MS10].

Hera-JVM [MS10]. Herman [Kie13].

Heterogeneous [ASV^{+}16, HHB^{+}14, Rub14, AYZI10, ABCR10, DFR13, MS10, SV18]. Heterogeneous-race-free [HHB^{+}14]. Heuristics [MGI14, LMK16]. HHVM [Ott18]. Hidding [RBL12]. hierarchy [BS13]. High [GSS^{+}16, Hol12, IRJ^{+}12, MSM^{+}16, RGB18, SWU^{+}15, URJ18, WN10, Zak10, BRWA14, Hos12, Ngo12, RFBJ14, TTD^{+}11, TGZ17, VWJB10, WFF18, WWH^{+}17, TRE^{+}13]. High-coverage [RGB18]. high-dimensional [TGZ17]. high-level [Hos12, RFBJ14, VWJB10]. High-Performance [URJ18, WN10, GSS^{+}16, BRWA14, Ngo12, TTD^{+}11, WFF18, WWH^{+}17]. higher [KT15, SPKT18]. higher-order [KT15, SPKT18]. highly [BP10, SPP^{+}10]. history [DRN14]. hit [Ano13]. Hoare [SD16b]. hole [Ano13]. Holistic [MAHK16]. Hop [WBHN18, D’H12]. Hopjs [SP16]. Horstmann [Gve13]. hosted [CBLFD12, SYZZ^{+}14]. hot [LMK16]. HotSpot [Sch13, BOF17]. HotWave [ABMV12, VBAM10b]. HPC [JQJ^{+}16]. HTM [CHM16]. HTML [Sta10]. HTML5 [HLO15, NKH16, Ano15]. Hunting [GGC18]. HVM [LTK17]. Hybrid [CHM16, JQJ^{+}16, JMO14, KCD12, VDV17, ZMNY14, ZMM^{+}16, ASME18, ADI13, HyG12, PdMG12, STA18, SWB^{+}15]. Hybris [VDV17]. hygienic [DFHF15]. hypervisor [GMC^{+}13].

i-Jacob [LYM^{+}18]. IaaS [ZLHD15]. Identification [MTSH16, PBM^{+}19, SBE^{+}19, BZD17, FMS^{+}11]. Identifier [SRTR17]. identifiers [FMS^{+}11]. Identifying [IN12, SVB^{+}17]. if [Han15, STA18]. If-transpiler [STA18]. illuminating [BK14]. Image [WN10]. immutability [HMDE12, ZPL^{+}10]. immutable [SV15b]. impact [CMS^{+}12, Gra15, HWLM11, MPR12, WKJ17]. imperative [RFRS14]. implement [HdM17]. Implementation [CSF^{+}16, GPT12, HM12, NBB18, CNRG19, OA17, Por18, VGRS16, YP10]. implementations [CSS^{+}16, OJ12, PS10a]. Implementing [FFF17, GM12, WCB16, EEK^{+}13, FBH17, PMP^{+}16, ZBvdB^{+}19]. implications [BRGG12]. implicit [IvdS16, SPAK10]. imply [BRGG12]. Improve [OTR^{+}18, QSaS^{+}16]. Improved [KRR^{+}14, UIY10, OJ12, XHH12]. Improvement [RC17]. Improving [ACS^{+}14, HWI^{+}12, TWSC10, WWG^{+}18, eBH11, UTO13]. in-depth [HTBB16, Rau14]. in-place [DVL13]. including [Den18]. Incremental [LHR19, DS16, ELW15, UIY10]. independent [IF16, VS11]. industrial [CRJ^{+}10]. inefficiently [XR10]. inefficiently-used [XR10]. Inference [BO13, YHY13, AGGZ10, CGJ^{+}16, HyG12, HMDE12, RKHN18, Zha12]. Inferring [PTRV18, AS14, BENS12]. InfiniBand [ETTD12, IRJ^{+}12]. infinite [ASdMGM14]. Inflow [ZMM^{+}16]. influence [MHR^{+}12]. Informa [HA13]. Information [ASF17, HBS16, KHL^{+}13, RKN^{+}18, SS12, AF12, ABFM12, BVGVEA11b, CMS^{+}12, PMTP12, RRB17, ZYY^{+}19]. Information-flow [HBS16]. Infrastructure [Den18, NG12, WCST19]. Inheritance [LN15, WT11, AST^{+}16, GBS13, NCS10]. Initial [LTD^{+}12]. initialization [AMT17, MME14, WSH^{+}19]. Initialize [WSH^{+}19]. Initiation [FGR12]. Injecting [ZZK13]. Injection [SBE^{+}19]. inline [DJLP10]. Inlining [BA12, STA18, HWM13]. input [Pha18]. insecure [YW13]. Insight [VF10]. instanceof [SMS^{+}12]. Instant [MHBO13]. instantiation [AST^{+}16]. instead [AGH^{+}17, BTR^{+}13]. instrumenting [CZ14]. Integrated [Tar11, YP10]. integrating [SPP^{+}10]. integration [Ame13, HKVG14, dARPH^{+}19, Sch10a]. integrity [HDK^{+}11]. Intel [CDMR19]. intelligence [JACS10]. Intelligent [Pau14]. IntelliMerge [SZZ^{+}19]. Intensive [LYM^{+}18, NWB^{+}18, SAdB^{+}16]. inter [CMM17]. inter-language [CMM17]. Interacting [SK13]. Interaction [WT11]. Interactive [FBG17, AMWW15, Dan18, JH11, MCY^{+}10]. intercession [VM10]. interdependencies [LBF12]. Interface [Liu14, MvDL12, SLS^{+}12, AYZI10, MT14, LT11, LT14]. Interfaces [WT11, Cho14, DLM10, LWH^{+}10, PSNS14, WT10]. interference [YDFF15]. International [Hol12, KP15, Fox17a]. Internetware [LYM^{+}18]. Internetware-Oriented [LYM^{+}18]. Interoperability [GSS^{+}18, GSS^{+}16]. Interpretation [BDT10, DLR16, DLM10, DLR14, NSDD17]. Interpretation-Based [DLR16]. interpreter [D’H12, KMMV14]. interpreters [CNRG19, HWW^{+}15, IvdS16, MD15, SYZZ^{+}14, ZLBF14]. Interprocedural [CPV15, FWDL15, ZMNY14]. Interrupting [AST12]. intersection [KT15]. intra [BJBK12]. intra-node [BJBK12]. Introducing [CJ19, Dan17, DMS11]. Introduction [CIAD13, CSZ17, HTLC10, HTW14, Lew13, RHT13, VK12, Hav11, VF10]. Introductory [BNP11]. intrusively [MZC10a]. Invasive [ADJG19]. Investigation [SS13, FH16, Tai13]. invited [Piz17, Sie17]. invocation [SPAK10, SS19, BVGVEAFG11]. invocations [BVGV14a]. invokedynamic [OCFLI14]. Involvement [ZMM^{+}16]. IP [TKL^{+}15]. iPhone [Sta10]. IR [LSWM16]. irregular [AC16]. ISAs [HNTL12]. ISBN [Ano15, Ano18, Bro12]. ISBN-13 [Bro12].

Isolation [ZLB^{+}13]. Issue [CJ19, DVL13, HL13, HTW14, Puf13, VK12, Fox17a, HTLC10, HGCA11, RHT13]. iterations [DD13]. iterators [ZLBF14]. IVE [CRJ^{+}10]. IVPs [KS15].

J [KMLS15]. J2M [LZYP16]. J2ME [GPT12]. J2ME-Enabled [GPT12]. Jaccie [KS14]. Jacob [LYM^{+}18]. Jalapeno [AFG^{+}11]. JAMES [DDDF17]. JaSTA [HD17]. JaSTA-2 [HD17]. Java [Bro12, Den18, Fox17a, Gve13, HWM11, HTW14, MvH15, Ngo12, Sch13, VK12, AÖ11, KvGS^{+}14, PQTGS17, SAdB^{+}16, ABC18, ASdMGM14, AST12, AFGG11, AYZI10, AdSCdR^{+}19, AS14, AAB^{+}10, Alt12, Ame13, AdCGGH16, AT16, And14, AFGM15, Ano12, Ano13, ABMV12, AGR12, AGR17, ABCR10, ADI13, ABFM12, AK13, BK12, BH17, BMR14, BH12, BDT10, BVGVEA10, BVEAGVA10, BVGVEA11a, BVGVEAFG11, BVGVEA11b, BVGVEA13, BVGV14a, BVGV14b, BS12, BMDK15, BO11, BO12, BO13, BP19, BCR11, BDGS13, BCD13, BD17, BRGG12, BIvdS17, Bla18, BR12, BH10, BR15, BB12, BNP11, BL15, BW12, BA12, BZD17, BSOG12, BMOG12, BKP16, BA17, BJBK12, CIAD13, FGB^{+}19, CSZ17, CZ14, CMM17, CWW13, CV14, CS12, CDTM10, CCFB15, CNRG19, CC15, CRJ^{+}10, CWGA17, CSF^{+}16, CSK17, CCH11, CJ17]. Java [CJ19, CYWD19, CDG^{+}17, Cle16, CDMR19, CKS18, CSdL16, CCA^{+}12, CMM^{+}10, CRAJ10, DJLP10, Dan18, DDDF17, DLM10, DLZ^{+}13, DVL13, DR10, DHS15, DJB16, DMS11, ECS15, EEK^{+}13, ES14, EQT10, Esq11, EABVGV14, Eug13, EV13, ETTD12, ETR^{+}15, FLZ^{+}18, FRGPLF^{+}12, FGR12, Fer13, FFF17, FLL^{+}13, FHSR12, Fox17b, FMS^{+}11, GLGA19, GMPS12, GMNC13, GvRN^{+}11, GYB^{+}11, GM12, GBS14, GD12, GBC12, GS11, GS12, Gon11, GMC^{+}13, GT10b, GJS^{+}13, GJS^{+}14, Gri17, GPT12, GK15, HL13, HD17, HdM17, Has12, HWM10, HWM13, HWM14, HA13, HM12, HTLC10, HKVG14, HTBB16, HH13, HOKO14, HGCA11, Hor11, Hor12, HC13, HC10, HZZK19, HWLM11, HJ12, IHWN12, IN12, IS18, IF16, JC10, JEC^{+}12, JQJ^{+}16, JJL17, Jen12, JB12, JYKS12, JTO12, JH11, J^{+}12, JMB12, JMO14]. Java [KHR11, KHM^{+}11, KMLS15, KS13, KW10, KW11, KPP^{+}18, KM10, KSR14, KSPK12, KDPG18, KS14, KF11, KB11, LSBV16, LSBV17, LTD^{+}12, LMK16, LSWM16, LLL13, LT11, LT14, LZYP16, LXP18, LYBB13a, LYBB13b, LYBB14, LZ12, LKP19, Loc13, Loc18, Lon10a, Lon10b, LMS^{+}12, LMS^{+}13, LO15, LPA13, LWC17, LTK17, LS11, Lyo12, MKZ^{+}14, MS13, MME^{+}10, MLGA11, MDS^{+}17, MCC17, MPM^{+}15, MHN19, MZC10b, MKTD17, MM16, MHM10, MAH12, MB12, MCY^{+}10, McN19, MGS19, MPR12, MLM17, MLM19, MKK^{+}12, MKK^{+}13, MSS10, MCW19, MvH15, MT14, MDHS10, NM10a, NCS10, NS12, Nil12a, Nil12b, NG13, NNTK17, NNT^{+}19, NM10b, NBB18, Oak14, OOK^{+}10, OMK^{+}10, OIA^{+}13, OUY^{+}13, OW16, OJ12, OCFLI14, PS11, PLL^{+}18, PdMG12, PTML11, PMTL14, PTHH14, PL12, PiLCH11, PBMH13, PBB19, PPMH15]. Java [PMP^{+}16, PQD12, PVH14, PTF^{+}15, PS10a, PS10b, PDPM^{+}16, Pos19, PSW11, Puf13, PKC^{+}13, QLBS17, RD15, RDCP12, RTE^{+}13, RTET15, RR14, RS12, RHT13, R^{+}13, RBL12, RAS16, RSI12, Rey13, Rez12, RVP11, RLMM15, RRB19, RB15, RvB14, SSL18, SSB^{+}14a, SE12, SRB18, SRTR17, STST12, SS12, Sch14, Sch13, Sch10a, SPPH10, SKKR11, SDH^{+}17, Sch10b, SSMGD10, SZ10, Set13, SMSB11, SMS^{+}12, SM12, SDM12, SWMV17, SW12, SGV12, SEPV19, SKBL11, SD16a, SJPS10, SLS^{+}12, SKR17, SS14, SABB19, SP10b, SMP10, Spo16, SBE^{+}19, SPP^{+}10, SWB^{+}15, SSB01, SSB14b, ST15, SMP19, SPS17, SSG^{+}14, SS19, STS^{+}13, Sve14, SWF12, TRTD11, TTD^{+}11, TTD12, TRE^{+}13, TLL11, TWX^{+}10, TFPB14, TN19, TWNH12, TNTN12, TGZ17, TJLL18, TKL^{+}15, UR15, UFM15, UPR^{+}18, VSG17, VGRS16, VBDPM16, VBMDP16]. Java [VGS14, VBAM10a, VBAM10b, VBMA11, WGF11, Wam11, WZdSOS17, WCST19, WLL19, WBM^{+}10, WK12, WCB16, WN10, WRI^{+}10, WA19, WHV^{+}13, WHIN11, WZL^{+}18, WBA^{+}11, WAB^{+}11, WWS13, XHH12, XR13, XMD^{+}17, Xue12, YP10, YKM17, YKA^{+}19, YDFF15, ZIvdS17, Zak12, ZP14, ZLCW14, ZHL^{+}12, ZXL16, ZBvdB^{+}19, ZKB^{+}16, ZYY^{+}19, ZWSS15, ZPL^{+}10, ZDS14, dCMMN12, dMRH12, eBH11, hED12, vdMvdMV12, Del13]. Java-Based [AFGG11, SLS^{+}12, ST15, SWF12, CJ17, CJ19, HOKO14, JMO14, KS13, KS14, MB12, MCY^{+}10]. Java-compatible [ABCR10]. Java-like [BDGS13, BCD13, DJLP10, SZ10]. Java-to-HDL [OUY^{+}13]. Java-to-JavaScript [LSWM16]. Java.utils.Collection.sort [dGRdB^{+}15]. Java/JSP [Sch10b]. Java/Scala [Pos19]. JavaBean [MZC10a]. JavaBIP [BMSZ17]. JavaCC [GN16]. JavaCOP [MME^{+}10]. JavAdaptor [PKC^{+}13]. JavaFX [FBG17, Top11]. JavaGI [WT10, WT11]. JavaScript [Ano15, Kie13, Ric14, Teo13, CH17, AMT17, ACS^{+}14, AHK^{+}15, AZMT18, AGM^{+}17, AMWW15, BNP^{+}18, BCF^{+}14, BBP13, Cec11, Cha18, CGJ^{+}16, CVG^{+}17, CBLFD12, Cho14, CDBD18, CHJ12, Dan18, Dei10, Dei11, DcSG12, DiP18a, DiP18b, DFHF15, FMM^{+}11, FM13, FH16, FBH17, FSC^{+}13, FZ17, FOPZ14, GMS12, Guo17, GGC19, HyG12, Hav11, HBS16, HLSK13, HHSS13, HC11, HOSC16, KR12, KSW^{+}14, KRH16, KT14, Ker15, KFBK^{+}15, Kie10, KBL14, KRR19, KARO12, Kri12, LSWM16, Ler10, LVG10, LPGK14, Liu14, LML17, MTL15, MLT17, MPS12, MGI17, MHL15, MRMV12, MTSH16, Mil13, MM12, MMP15, Mor18, NKH16, NSDD17, OBPM17, PWSG17, PWSG19, PGA18, PLR14, PSR15, PLR18, PKPM19, PDDD17, PKO^{+}15, Por18, Rau14, dARPH^{+}19, RLBV10, RGEV11, RHN^{+}13, RW17, Ryu16, RPP19, SMN^{+}18, STA18]. JavaScript [Ser18, Sev12a, Sev12b, SVB^{+}17, SDC^{+}12, Sta10, Ste10, SR17, SFR^{+}14, TAF^{+}18, TT11, VM15, VP16, VB14b, Wal12, WCST19, WXR16, YW13, Zak18, Zak10, dJM18, BM18, KCD12, Mei14, Ano18, Kie13, Teo12, Teo13]. JavaScriptCore [Piz17]. JaVerT [SMN^{+}18]. JAWS [PKO^{+}15]. JBInsTrace [CZ14]. JCloudScale [ZLHD15]. JCML [dCMMN12]. JCSI [ABFM12]. JCSP [WBM^{+}10]. JDiffraction [PQTGS17]. JDK [SRB18]. JDMM [ZP14]. jDSSAT [dARPH^{+}19]. JEqualityGen [GRF11]. JET [LT11]. JGRIM [MZC10b]. Jinn [LWH^{+}10]. JIT [BBF^{+}10, BB17, CMS^{+}12, HWM14, IHWN12, JK13, NED^{+}13, Ott18, RSB^{+}14, WKJ17, ZYZ^{+}12]. JIT-based [BB17]. JITs [KRCH14]. jMarkov [CRAT^{+}12]. JML [CRJ^{+}10, TJLL18]. JML-annotated [TJLL18]. JNI [CDG^{+}17]. Joe [Ano18]. Johnny [WA19]. join [MZC10a]. Jonge [Ngo12]. Journey [Ryu16]. joy [FH11]. JP2 [SSB^{+}14a]. JPC [CMM17]. JPF [BA19, WKG17, WCG^{+}18]. JPR [WKG17]. jQuery [AM14, PIR17]. JR [OW16]. JR-like [OW16]. JRE [CZ14]. JS [AHK^{+}15, Por18]. js-emass [Por18]. Js of ocaml [VB14b]. JSART [MM12]. JSCore [Cha18]. JSetL [RB15]. JSON [BB17]. JSormdb [Dei10]. JSP [Sch10b]. JTabWb [FFF17]. JTRES [HTW14]. JTRES2011 [RHT13]. JTRES2013 [Fox17b]. JTRES2014 [Fox17a]. judgment [CSV15]. Julia [Dan18, Spo16]. Juliet [BB12]. July [Bro12, KP15]. Jump [WBHN18]. jungle [Sew12]. Jupyter [Dan18]. JupyterLab [Dan18]. Just [DLR16, TN19, KHL^{+}13, LMK16, MGI17, TTS^{+}10]. Just-In-Time [TN19, DLR16, KHL^{+}13, LMK16, MGI17, TTS^{+}10]. JVM [AC16, AFG^{+}11, CSS^{+}16, Guy14, MS10, PVH14, R^{+}13, RRB17, SYZZ^{+}14, SV15b, Sub11, WKG17]. JVMs [BK14, ZYZ^{+}12].

K-Java [BR15]. kernel [HDK^{+}11]. Key [BBB^{+}17, DFR13, JB12]. key-value [DFR13]. keynote [McK16]. Kirk [Del13]. KiWi [BBB^{+}17]. KJS [PSR15]. Knoernschild [Del13]. knot [LBF12]. know [DJB16, Gra15, Han15]. Knowledge [KSPK12, UMP10]. known [Han15]. Kraken [Ano14].

Lake [Hol12]. lambda [MKTD17]. lambdas [UFM15]. landscape [Sve14]. Language [DLPT14, GJS^{+}13, GJS^{+}14, GSS^{+}18, JC10, KSPK12, MAHK16, NM10b, Sev12b, SS13, WBHN18, ABCR10, CMM17, CSdL16, DAA13, EKR^{+}12, Fee16, GSS^{+}16, Hos12, HWW^{+}15, KRCH14, LWH^{+}10, LE16, MDM17, SC16, SZ10, SKR17, SNS^{+}14, VB14a, WCG14, WWH^{+}17, ZWSS15, dCMMN12]. language-level [WCG14]. Language-Neutral [WBHN18]. Languages [CSGT17, MSM^{+}16, PTHH14, YKM17, AGGZ10, BCD13, CMS^{+}12, DTM^{+}18, EEK^{+}13, ER14, FMBH15, Han15, HBT12, HJS^{+}10, KRR^{+}14, MSM^{+}10, NED^{+}13, PULO16, SPKT18, SPY^{+}16, Zha12]. LARD [WCG14]. Large [BA17, AST^{+}16, CCFB15, CJ19, LSBV16, LSBV17, MDS^{+}17, MCY^{+}10, PTF^{+}15, WHIN11]. Large-Scale [BA17, CJ19, MDS^{+}17, MCY^{+}10, PTF^{+}15, WHIN11]. Larus [DD13]. Latency [MV16, ETR^{+}15, JH11]. lawn [CH17]. laws [DMS11]. Layer [OTR^{+}18, SKKR11, Den18]. layered [RCR^{+}14]. lazy [TD15]. Leading [MSS10, PGA18]. leak [SS14, XR13, YSCX17]. Leaks [And14, RW17, VB18]. LeakSpot [RW17]. lean [BRGG12, SV15b]. Learn [RT14]. Learning [Dan18, JJCO19, PSJ18, Pau14, RT14, BSAL18, CNS13, KC12, Ano15, Teo13]. learnt [GY16]. Legacy [KH18, SVB^{+}17, CDTM10]. Legally [Sam12]. length [SMP10]. Less [BNE16]. Lessons [URJ18]. Level [AC16, MGI14, SWU^{+}15, YXS^{+}19, EKUR10, Hos12, IHWN12, KBL14, LWC17, MGI17, RFBJ14, TTD^{+}11, VWJB10, WCG14]. leveraging [WCST19]. Lexical [GN16]. Lexicon [TAF^{+}18]. Libraries [BK12, RDCP12, BIvdS17, Cho14, EKR^{+}12, PMTL14, PLR18, TTD^{+}11]. Library [CH17, CWGA17, NBB18, OCFLI14, TAF^{+}18, WN10, dJM18, CMM17, PMP^{+}16, PQTGS17, Pos19, TFPB14, TGZ17]. License [GD12]. Life [Esq11]. LIFT [BTR^{+}13]. Light [MvH15]. Light-weight [MvH15]. Lightweight [BW12, KBL14, KKK^{+}17, RO12]. like [BDGS13, BCD13, DJLP10, PMTL14, SZ10, VGS14, OW16]. Lime [ABCR10]. line [SV17]. linearizability [LTZ14]. lines [BTR^{+}13, KATS12]. linguistic [UR15]. Linux [Ric14]. Linux-basierte [Ric14]. Listener [JH11]. little [Han15]. liveness [LDL14]. Load [GMNC13, PDPM^{+}16]. loaders [SM12]. loading [WGF11]. Local [NBB18, DDDF17]. localised [SP10b]. locality [HJH10, OJ12]. localize [ZZK13]. location [NCS10]. Locators [SDM12]. Lock [FC11, NM10a, NFV15, UMP10]. Lock-free [FC11, NFV15]. Locking [GGRSY17, JTO12, GGRSY14, GGRSY15, YKA^{+}19]. locks [SPS17]. Logging [CJ19, CJ17]. Logic [ZLNP18, GMS12, Pha18, SD16b]. loop [DD13, HWI^{+}12, PLR18]. Loops [RD15, LLL13]. loss [WHIN11]. Low [ETR^{+}15, GM12, SWU^{+}15, WCG14, ZHCB15, ZFK^{+}16, BCR13, XMA^{+}10].

Low-Budget [GM12]. Low-latency [ETR^{+}15]. Low-level [WCG14]. Low-overhead [ZHCB15, ZFK^{+}16]. low-utility [XMA^{+}10]. lunch [DTLM14].

m [MZC10b]. m-JGRIM [MZC10b]. M2M [Pau14]. Machine [JJCO19, LYBB14, Ame13, CBLFD12, KS13, KC12, McM11, Piz17, SSMGD10, WGF11, WHV^{+}13, BZD17, Cle16, LYBB13a, LYBB13b, LTK17, PTHH14, RRB19, SSB^{+}14a, Sch13, Set13, SMSB11, SGV12, SSB01, SSB14b, UR15]. Machine-Learning [JJCO19]. Machines [AGR12, GTS^{+}15, JK13, KRCH14, NK10]. macros [DFHF15]. Magic [SP10b]. Magic-sets [SP10b]. Magnitude [BNE16]. major [Ano12]. Making [Loc13, Sta10, PS11]. malformed [SHU16]. Malicious [KCD12, AFGM15]. malleable [MZC10a]. malware [CSK17]. Managed [MAHK16, NFN^{+}18, NWB^{+}18, BM14, CBGM12, GTL^{+}10, ZIvdS17]. Managed-Language [MAHK16]. Management [OTR^{+}18, Pau14, YPMM12, AHK^{+}15, BVGV14a, BGS^{+}13, EKUR10, HB13, KCP^{+}17, KB17, MLM17, Nil12b, PCL14, SWB^{+}15, Tar11, WGW^{+}11]. manipulating [KRR19, YS10]. Manipulation [MS14]. manual [KCP^{+}17, KPP^{+}18]. many [GTSS11, SV18]. Map [BBB^{+}17]. mapped [SV15b]. Mapping [LTD^{+}12, UR15]. MapReduce [LZYP16, RFRS14, SKBL11]. maps [NFV15, SV18]. mashup [ETR12]. Masses [BMSV18, BSO18, IvdS16]. Massive [BMSV18]. Massively [NBB18]. mastering [Sub11]. Math.Js [dJM18]. Mathematical [BW12]. Mathematics [dJM18]. MATLAB [Alt12, FBH17, PMTL14, VF10, Has12]. MATLAB-like [PMTL14]. matrix [HD17, TGZ17]. matters [DJB16]. Maxine [WHV^{+}13]. MCAPL [Den18]. me [LCW18, CMM^{+}10, GM12, XHH12].

ME-Based [GM12]. mean [Rub14]. Means [SS19]. Measurement [YPMM12, YW13]. Measurement-Based [YPMM12]. Measuring [DW10, DTLM14, Gra15, JH11]. mechanical [ZZK13]. mechanised [BCF^{+}14]. Mechanising [Loc18]. Media [Bro12]. meets [KHL^{+}13]. Memento [CPST15]. memoization [TPG15]. memories [ASME18]. Memory [BG17, JYKS12, MSM^{+}16, NWB^{+}18, OTR^{+}18, SS14, ST15, WZL^{+}18, AHK^{+}11, AHK^{+}15, AGGZ10, BSMB16, BFS^{+}18, CWW13, DLZ^{+}13, DVL13, FC11, FF10, GYB^{+}11, HHB^{+}14, HB13, KHL^{+}17, KCP^{+}17, KB17, Loc13, MSM^{+}10, MLM17, Nil12b, OMK^{+}10, RW17, SMS^{+}12, SEPV19, SMN^{+}12, SWB^{+}15, SV15a, Tar11, TVD10, VB18, WGW^{+}11, XR13, YSCX17, ZP14, ZHCB15, ZBB17]. memory-performance [SEPV19]. MemSAT [TVD10]. merge [ABC18]. Mergesort [LL15]. merging [SZZ^{+}19, TLX17]. Message [KF11, ETTD12, TRTD11, TTD12, UR15]. message-passing [ETTD12, TRTD11, TTD12, UR15]. messages [eBH11]. meta [MD15, SZ10]. meta-circular [SZ10]. meta-compilation [MD15]. metadata [DVL13, WCST19]. MetaFJig [SZ10]. metaheuristics [DDDF17]. metaprogramming [PS11]. Method [AC16, BVGVEAFG11, BA19, GD12, AST12, AJL16, HMDE12, SS19, SS16, VBMDP16, ZYY^{+}19]. Method-Level [AC16]. Methods [MM16, Pau14, VBZ^{+}18, Bra14, GRF11, LSBV16, LSBV17, SSL18]. Metrics [KB11, JK11, SSK13, Sch13]. Metriken [Sch13]. Microscopic [RXK^{+}17]. Microservices [KH18, LSCPE18]. Microsoft [Ano13]. Middleware [RTE^{+}13, AdSCdR^{+}19, HOKO14, HWLM11, MZC10b]. middleweight [IF16, MT14]. midstream [SSG^{+}14]. Migrating [AST^{+}16, CDTM10, FGB^{+}19]. Migration [OwKPM15, Fee16]. migrations [TFPB14].

Miniboxing [UTO13]. minimal [CNS13]. mining [DRN14]. Mint [WRI^{+}10]. minute [DHS15]. minutes [BTR^{+}13]. misconfigurations [MCC17]. Mismatch [YCYC12]. misses [IN12]. Missions [WCB16]. Mistakes [BA17]. Mitigating [BGS^{+}13, KC12]. mixed [CL17]. Mobile [GM12, GPT12, LYM^{+}18, MV16, XHH12, GMNC13, GGC18, KF11, MZC10b, MTSH16]. Mock [SABB19]. Model [Bul18, CSF^{+}16, CDG^{+}17, CCA^{+}12, DLR16, FSK12, JJCO19, JYKS12, Loc18, MSM^{+}16, MCC17, MV16, BVGVEA11a, FGB^{+}19, CHM13, CWW13, CV14, CS12, CSKB12, DLZ^{+}13, FLZ^{+}18, GY16, HAW13, Loc13, LSSD14, MLT17, MSM^{+}10, PSW11, RR14, RBV16, RAS16, RDF15, SMN^{+}12, SSG^{+}14, SS19, Tai13, VWJB10, ZP14, ZXL16]. Model-Aware [JYKS12]. Model-based [MCC17, PSW11]. model-driven [FGB^{+}19, CHM13]. Modeling [GBC12, JC10, KSPK12, LDL14, Rey13, SM12, CRAT^{+}12, SKR17, TLX17, ZIvdS17]. Modelling [CSZ17]. Models [CC15, PE11, ZLCW14, AGR17, HHB^{+}14, TVD10, ZBB17]. Modern [LMZP19, FIF^{+}15, Hav11, JK13, KB17, Mor18, Teo13, WGW^{+}11, ZDK^{+}19]. Modernization [KH18, Nil12a]. modes [BP19]. Modified [GT10a]. Modular [IvdS16, LN15, RDCP12, AACR18, MRA^{+}17, RO12]. Modularisation [SDM12]. modularity [Del13, SPAK10]. module [KR12, dARPH^{+}19]. Modules [Bla18, PiLCH11]. monad [GSD^{+}15]. MongoDB [Guo17]. monitor [STA18]. Monitoring [AGR12, MRF18, CMM^{+}10, DJLP10, ES14, KF11]. Monitors [BLH12, HM12]. Morgan [Ano18]. mori [CPST15]. movement [NCS10]. MPI [RAS16, SZ11, VGRS16]. MPI-based [SZ11]. MPJ [JQJ^{+}16, TTD12]. MrCrypt [TLMM13]. MS [FH16]. Multi [GSS^{+}18, JTO12, RTE^{+}13, BGS^{+}13, DSEE13, Fee16, FC11, GSS^{+}16, IHWN12, MS10, Puf13, SE12, SKBL11, SV18, TRTD11, Tar11, WRI^{+}10, YKA^{+}19]. Multi-Core [RTE^{+}13, MS10, TRTD11]. multi-cores [SKBL11]. multi-engine [Tar11]. multi-granularity [YKA^{+}19]. Multi-Language [GSS^{+}18, Fee16, GSS^{+}16]. multi-level [IHWN12]. multi-maps [SV18]. multi-processor [Puf13]. multi-stage [WRI^{+}10]. Multi-threaded [JTO12, DSEE13, SE12, Taf13]. multi-threats [BGS^{+}13]. multi-version [FC11]. Multicore [ASV^{+}16, CCH11, MKG^{+}17, SE12, SSMGD10, TWX^{+}10]. multilevel [JK13]. multiphase [GvRN^{+}11]. Multiplatform [ZKB^{+}16]. Multiple [AF12, ASF17, HLSK13, CSV15, DD13]. multiplexing [BVGVEAFG11]. Multiprocessing [VGS14]. multiprocessor [PS10b, PWA13, SPS17]. Multiprocessors [KW11, RS12]. Multithreaded [KKW14, Loc18, SR14a, BNS12, DJLP10, Fer13]. Multithreading [CCH11]. multivariate [AÖ11]. multiway [YKA^{+}19]. Mungo [KDPG18]. MuscalietJS [RCR^{+}14]. Mutagenic [YCYC12]. mutants [FRC^{+}17]. Mutation [MMP15, KPP^{+}18]. mutators [AHK^{+}11]. MySQL [Ano15].

Names [SRTR17]. Naming [STST12].

Native

[JQJ^{+}16, LT11, LT14, KFBK^{+}15, STS^{+}13]. Natural [LL15]. naturalness [HBG^{+}16]. NDetermin [BENS12]. nested [CHM16, ZLB^{+}13]. Netflix [Liu14]. Network [CC15, GGC18, GGC19, RR14]. Networking [Hol12]. Networks [AFGG11, ETR^{+}15, ZYY^{+}19]. neural [ZYY^{+}19]. neuromorphic [HNTL12]. Neutral [WBHN18]. Next [YWW^{+}18, CRJ^{+}10, CMM^{+}10]. Next-Generation [YWW^{+}18]. NG2C [BOF17]. NGS [YWW^{+}18]. NGS-FC [YWW^{+}18]. Nicolai [Bla18]. Nixon [Ano15]. No [BVGVEA10]. No-Heap [BVGVEA10]. NoCs [PWA13]. Node [HC11, BJBK12]. Node.js [BSMB16, MTL15, Ano14]. nodes [DRN14]. Nominal [BO13]. Non [BVGVEA11b, BSOG12, GGZ^{+}15, TD17, WZL^{+}18, YKM17, MZC10a, OMK^{+}10, SSL18, ZP14]. Non-Adequate [GGZ^{+}15]. non-cache-coherent [ZP14]. non-cloned [SSL18]. Non-equivocation [TD17]. Non-functional [BVGVEA11b]. non-intrusively [MZC10a]. Non-Java [YKM17, OMK^{+}10]. Non-termination [BSOG12]. Non-volatile [WZL^{+}18]. Non-Volatility [WZL^{+}18]. Nonblocking [RTET15, SP10a]. Nondeterministic [RB15, BENS12]. noninterference [IF16]. Nopol [XMD^{+}17]. Normalization [ADJG19]. NoSQL [DFR13]. Notation [Sev12a]. Novel [NK10, MZC10b]. November [Hol12]. Novice [BA17]. Novices [RT14]. null [AT16]. NullPointerExceptions [BSOG12]. NUMA [GTS^{+}15]. NumaGiC [GTS^{+}15]. number [PPMH15, SLF14]. Numbers [Jaf13, AJL16, Wal12]. Numerical [HCLH18, KS15, KFBK^{+}15, PQTGS17]. NXT [SWF12].

Obfuscated [KCD12]. obfuscation [CCFB15]. obfuscations [CSK17]. Object [CSGT17, GLGA19, GS11, KB11, LZ12, NWB^{+}15, PTHH14, PiLCH11, RC17, Sev12a, SW12, AST^{+}16, BZD17, DDDF17, FMBH15, IvdS16, KRR19, LX19, MME14, MHBO13, RDF15, UJR14, VM10, WM10, ZCdSOvdS15, Zha12, ZDS14, hEYJD12]. Object-Bounded [NWB^{+}15]. object-constraint [FMBH15]. object-manipulating [KRR19]. Object-Oriented [GS11, KB11, RC17, PTHH14, AST^{+}16, DDDF17, MHBO13, VM10, ZDS14, hEYJD12]. object-sensitive [LX19]. Objective [Sta10]. Objective-C [Sta10]. Objects [BS12, RKN^{+}18, AFGM15, MHL15, SK13, SABB19, WXR16, BVGVEA10]. Observations [AAB^{+}10]. OCaml [Cle16]. OCaml-Java [Cle16]. OCTET [BKC^{+}13]. odeToJava [KS15]. offloading [ZHL^{+}12]. on-demand [ZHL^{+}12]. On-Stack [WBHN18]. On-the-fly [URJ18, UJR14]. once [WSH^{+}19]. one [SV18]. ones [AST^{+}16]. Online [NG13, GGC18, GGC19, HCV17, NK10]. only [NM10a]. Ontology [KSPK12]. OoOJava [JhED11]. Open [BSA14, GD12, ABC18, CJ17, CJ19, EKUR10, JK11, Tai13, VGRS16]. Open-Source [BSA14, ABC18, Tai13]. OpenJDK [BFS^{+}18, CHM16, dGRdB^{+}15]. OpenMP [VGS14]. OpenMP-like [VGS14]. operating [HDK^{+}11]. operation [KKW11]. operations [MHN19, TABS12, TGZ17]. Operator [PQD12]. opportunities [TPG15]. Optimal [AD16, JCMM19, SK12, ELW15]. optimale [Sch13]. optimally [BGS^{+}13]. optimisation [PPS16]. optimistic [WGF11]. Optimization [LTD^{+}12, RRB19, YKM17, AFG^{+}11, BDB11, DDDF17, JMO14, KS13, KC12, NG12]. Optimizations [DR10, BB17, CPST15, DS16, NG13, SAdB^{+}16]. Optimized [PKPM19]. Optimizing [LYM^{+}18, SV15b, WZK^{+}19, YRHBL13, HWW^{+}15, KRH16, MD15, ZLBF14]. optional [CMS^{+}12]. Oracle [LMS^{+}12, Sam12]. ORB [OUY^{+}13]. Orchestrating [CDBD18]. Order [SGD15, SMP19, JhED11, KT15, SPKT18, TD15]. ordering [KC12]. Orders [BNE16]. ordinary [MZC10a]. O’Reilly [Ano15, Bro12]. Oriented [ABMV12, BH10, GLGA19, GS11, KB11, LYM^{+}18, RC17, AST^{+}16, DDDF17, EABVGV14, MHBO13, PTHH14, RVP11, VM10, VBAM10b, WBA^{+}11, ZDS14, hEYJD12]. OSck [HDK^{+}11]. OSGi [BVGVEA13, GD10, Del13]. OSS [ZMM^{+}16]. other [EKUR10, KS13]. out-of-order [JhED11]. out-of-thin-air [OD18]. output [KM10]. Over-exposed [VBDPM16]. overhead [BCR13, ZHCB15, ZFK^{+}16]. overlap [ADJG19]. overlay [CDTM10]. Overloading [PQD12]. overview [Nil12b]. own [MPM^{+}15]. Ownership [ZPL^{+}10, BDGS13, DDM11].

PaaS [ZLHD15]. Package [SLS^{+}12, CRAT^{+}12, MB12, OW16, AK13]. Packages [PiLCH11]. PackedObjects [YKA^{+}19]. panic [Ano12]. Paper [DDDF17, PDPM^{+}16, Cha18, SV15a]. paperback [Ano18]. Papers [DVL13, HL13, LMK16, Puf13]. Parallel [DS16, Esq11, LLL13, LHR19, MKG^{+}17, NKH16, NBB18, QSaS^{+}16, RD15, RSI12, AACR18, BP10, BBP13, BSMB16, CRP^{+}10, MGS19, NG12, NG13, PPMH15, Sie10, SZ11, TTD12, Taf13, VYY10, BKP16, WN10]. Parallelisation [GS11]. Parallelism [NKH16, BENS12, HHSS13, MZC10a, RHSD15, TWL12, ZLB^{+}13]. Parallelization [DTM^{+}18, SS16, YRHBL13]. parallelize [LPA13]. Parallelizing [NKH16, hEYJD12]. parameters [GBS14]. Parametric [AGGZ10, PULO16, UTO13]. Parlog [Bla18]. Part [KP15]. ParTejas [MKG^{+}17]. Partial [CSK17, JB12, SGD15, SMP19, BS13, LX19, MD15, TD15, WGF11, WWH^{+}17]. Partial-Order [SGD15, TD15]. Partially [BLH12, BCR11]. Partitioning [AD16, BS12]. party [FOPZ14, LVG10]. passing [ETTD12, TRTD11, TTD12, UR15]. Path [SGD15, AZLY18, DD13, HHSS13, SMP10]. path-based [AZLY18]. path-length [SMP10]. Path-Sensitive [SGD15]. pathfinder [KPP12, CS12, MPR12, NNTK17, NNT^{+}19, PdMG12, SM12, vdMvdMV12, Den18, RR14, SMP19]. patient [EKUR10]. patient-level [EKUR10]. pattern [GSD^{+}15, SAdB^{+}16]. Patterns [RC17, BVGVEA11b, Del13, Ste10]. PayPal [Ano14]. PCR [YCYC12]. PCR-RFLP [YCYC12]. PE [JB12]. PE-KeY [JB12]. perceptible [JH11]. Perfect [SLE^{+}17]. Perform [LMZP19]. Performance [AACR18, CSZ17, CCH11, DR10, GBC12, Hol12, HJ12, MSM^{+}16, Oak14, OCFLI14, QSaS^{+}16, RVT18, TRE^{+}13, TPG15, THC^{+}14, URJ18, VP16, WWG^{+}18, WN10, ACS^{+}14, AAB^{+}10, BRGG12, BRWA14, CBGM12, Dei11, GSS^{+}16, HWI^{+}12, IRJ^{+}12, JH11, Ngo12, ODL15, PSNS14, SE12, SEPV19, TTD^{+}11, TWX^{+}10, WFF18, WHIN11, WWH^{+}17, Zak10]. performance-guided [PSNS14]. permission [HBT12, SNS^{+}14]. permits [PPS16]. Persistence [LZ12]. Perspective [YHY13]. Pert [LZ12]. pervasive [MHM10]. pgs [Ano18]. PHALANX [VYY10]. phase [KC12]. phase-ordering [KC12]. phone [GMNC13]. phoneME [RDCP12]. Phosphor [BK14]. PHP [Ano15, Ott18, TTS^{+}10]. Phynx [EKUR10]. Physics [Zak18, JEC^{+}12]. pickler [MHBO13]. pickles [MHBO13]. PINs [MTSH16]. pipeline [LPA13]. pipelines [CRP^{+}10]. Pivot [AD16, MRF18]. PL [FGB^{+}19]. PL/SQL [FGB^{+}19]. place [DVL13]. Plan [DLZ^{+}13]. Platform [AFGG11, PE11, WBHN18, BD17, CRJ^{+}10, CMM^{+}10, GD10, GMC^{+}13, MKZ^{+}14, PWA13, YP10]. Platforms [DR10, Has12, BP10, JMO14, KSR14]. PLDI [FLL^{+}13]. pluggable [MME^{+}10]. Point [Jaf13, AJL16]. Pointer [LHR19, LTMS18, LX19, TL17]. Pointers [RKN^{+}18, AT16]. Points [BK12, SDC^{+}12, BSAL18, DHS15, SBK13, TLX17]. Points-To [SDC^{+}12, BSAL18, DHS15, SBK13, TLX17]. Policies [FHSR12, MPS12, BVGV14a]. policing [DW10]. Policy [YPMM12, JK13]. Policy-Driven [YPMM12]. polyglot [EV13]. Polymorphic [Zha12]. polymorphism [GMT14, PULO16, UTO13]. polynomial [Pos19]. POPL [BCR13]. Popular [Has12, SRB18]. Popular-but-Seemingly-Dissimilar [Has12]. portable [BM18, LTK17, RGM13]. portal [MCY^{+}10]. Power [MV16, Pau14, BRGG12, CBGM12, Kie13, THC^{+}14]. pp. [Bro12]. PQL [RSI12]. Practical [AMT17, JACS10, SLES15, VS10, WBHN18, WWH^{+}17, FGB^{+}19, FIF^{+}15, WT10]. Practice [HGCA11, AS14, EKUR10, LWC17, TRE^{+}13]. practices [CJ17, YW13]. Pragmatic [Ano18, RO12]. pre [SBK13]. pre-processing [SBK13]. Precise [PIR17, TN19, XR13, BHSB14, CVG^{+}17, HyG12, PLR18, PG12, RGM13, TLX17, WFF18]. Precise-Yet-Efficient [TN19]. Precision [LTMS18, LX19, RSB^{+}14]. Precision-guided [LTMS18]. Precision-preserving [LX19]. Predicate [PL12]. predictable [LTK17]. Predicting [BSA14, RVK15, RVK19, AZLY18]. prediction [ZWZ^{+}14]. predictive [RGB18]. Presence [RK19, ZBB15]. preserving [AK13, LX19]. pressure [DTLM14]. pretenuring [BOF17]. Preventing [AFGM15, MSSK16]. prevention [VS11]. Price [Ano18]. Primer [YCYC12]. primitives [BJBK12]. Principles [HGCA11, JEC^{+}12, VM10]. Printf [ALB^{+}19]. Printing [AJL16]. Prioritization [LMZP19, MT13]. Prioritized [NGB16]. Priority [ASV^{+}16, HM12]. Privacy [And14]. Proactive [CL17, BGS^{+}13]. PROB [YP10]. Probabilistic [RBV16, GY16, ZWZ^{+}14]. Problem [YHY13, ZW13, J^{+}12, KC12]. problem-solution [J^{+}12]. problems [TPG15]. Proceedings [Hol12, KP15]. Process [SK12, AGR17, GT10a]. Processes [BMDK15, CDBD18]. Processing [LLL13, WN10, SBK13, SSG^{+}14, UJR14, ZDK^{+}19]. Processor [TKL^{+}15, Puf13, SPPH10, SMN^{+}12]. Processors [ASV^{+}16, MKG^{+}17]. producers [DAA13]. product [BTR^{+}13, KATS12, KvRHA14, SV17]. product-based [KvRHA14]. production [RGM13]. professionals [JACS10]. profile [Ott18, VSG17, WKJ17]. profile-guided [Ott18]. profiler [DTLM14]. profilers [MDHS10]. profiling [DD13, JH11, KRH16, NK10, RCB17, SSB^{+}14a, STY^{+}14, THC^{+}14, WLL19, XR13, ZBB15]. Program [BGK17, JJCO19, KKW14, RVK15, RT14, WCG^{+}18, ZKB^{+}16, AZLY18, AÖ11, DS16, GMS12, HCN14, JJL17, JWMC15, KM10, KMZN16, MKZ^{+}14, NS13, RKHN18, RVK19, Sch10a, SPY^{+}16, Tai13, TABS12, UPR^{+}18, WGF11, ZMG^{+}14]. Programmable [OA17, AYZI10]. Programmers [Esq11, RLMM15, Rau14]. Programming [AFGG11, ABMV12, BCR11, Bro12, BA17, DLPT14, GLGA19, HWM11, HGCA11, Köl10, KSPK12, LM15, McK16, NM10b, OAC18, PTML11, RSI12, RB15, SS13, Sub11, Alt12, AMWW15, BCvC^{+}13, BMR14, BSMB16, BRWA14, CL17, ECG12, EV13, FMBH15, Han15, HA13, Hav11, Lew13, MSM^{+}10, MGS19, MvH15, OW16, PTF^{+}15, RVP11, RFBJ14, SNS^{+}14, SGG^{+}17, TB14, UFM15, VWJB10, VBAM10b, Wam11, WRI^{+}10, WBA^{+}11, ZWSS15]. Programs [AGR12, BH17, BR12, BMOG12, CYWD19, GS11, JB12, LTD^{+}12, PSJ18, STST12, SS12, SDM12, SR17, TN19, XMD^{+}17, ZLCW14, AACR18, ASdMGM14, AZMT18, AdCGGH16, BA12, BNS12, CDBD18, DJLP10, ECS15, ES14, EP14, Fer13, HL13, IN12, KRR19, LKP19, LMS^{+}13, LO15, LPA13, MHN19, MRMV12, MCW19, NG12, OJ12, PL12, RR14, RAS16, RLBV10, SMS^{+}12, SZ11, SJPS10, SHU16, Taf13, WCST19, YS10, YSCX17, dCMMN12, hEYJD12]. Progress [WCG^{+}18, Sie17, ZHCB15]. Project [Wan11]. Projects [LMZP19, ZMM^{+}16, ABC18, CJ17]. Projekte [Ric14]. Prolog [CMM17, Tar11]. promises [AZMT18, MLT17]. promising [KHL^{+}17]. Proof [LL15]. Proofs [BMOG12]. propagation [IvdS16, PQTGS17]. Properties [BO11, CYWD19, RVK15, SS12, AZLY18, FWDL15, RVK19, SD16b, YS10]. Protecting [MPS12]. Protein [YHY13]. Protocol [GM12, FGR12]. protocols [KDPG18, PS10a]. prototyping [PWA13]. Provably [AdCGGH16, DJLP10, PS10a]. providing [OW16]. proving [AGH^{+}17, Taf13]. Proxies [VM10, Eug13, KT14]. PSE [KS15]. pseudorandom [PPMH15, SLF14]. PT [MGS19]. Published [Ano18, LSBV17]. pure [SS16]. Purely [RSI12, NFV15, SV18]. Purely-Declarative [RSI12]. purely-functional [NFV15]. Purity [NSDD17, HMDE12]. purpose [AdSCdR^{+}19]. Putting [BNP^{+}18]. PYE [TN19]. Python [Ric14, Dan18].

QoS [YPMM12]. qualitas [TMVB13]. Qualitas.class [TMVB13]. Quality [BNP11, CCFB15, WKJ17]. Quantitative [CPV15, GYB^{+}11, MRA^{+}17, PMTP12]. queries [GK15, MRA^{+}17, SGG^{+}17]. query [FWDL15]. query- [FWDL15]. questions [KM10]. Quicksort [AD16].

R [CH17, Dan18, KMMV14, NL14, SLS^{+}12, Vit14]. Race [BH10, EP14, RD15, AMT17, BGOS18, EQT10, HHB^{+}14, RGB18, WFF18]. race-aware [EQT10]. RacerD [BGOS18]. races [FF10, WCG14, XXZ13]. Racket [YK14]. racy [SRJ15]. Rady [Teo12]. Rails [Teo12]. Range [BS12]. Ranged [FSK12]. rapid [PWA13]. rationing [ASME18]. raw [HH13]. rays [SBF^{+}10]. RCDC [DNB^{+}12]. RDMA [ETR^{+}15, IRJ^{+}12]. RDMA-based [IRJ^{+}12]. RDMA-enabled [ETR^{+}15]. re [NCS10]. re-location [NCS10]. Reachability [NS13]. React [HOSC16]. reaction [SRB18]. reactive [BCvC^{+}13, MvH15]. read [NM10a]. read-only [NM10a]. Reading [Jaf13]. ready [RHSD15]. Real [BVEAGVA10, BBB^{+}17, Fox17b, HTW14, KW11, Nil12a, Pau14, SLES15, SLE^{+}17, VK12, BCR13, BVGVEA10, BVGVEA11a, BVGVEA11b, BVGVEA13, BVGV14a, BVGV14b, CRAJ10, DW10, EABVGV14, Fox17a, GMC^{+}13, HTLC10, KHM^{+}11, KPHV11, KvGS^{+}14, KW10, KPP^{+}18, KSR14, LTK17, MDS^{+}17, PS10b, PZM^{+}10, PSW11, Puf13, RHT13, SP10a, Sie10, SPS17]. Real-Time [BVEAGVA10, BBB^{+}17, Fox17b, HTW14, KW11, Pau14, SLES15, SLE^{+}17, VK12, Nil12a, BCR13, BVGVEA10, BVGVEA11a, BVGVEA11b, BVGVEA13, BVGV14a, BVGV14b, CRAJ10, DW10, EABVGV14, Fox17a, GMC^{+}13, HTLC10, KHM^{+}11, KPHV11, KvGS^{+}14, KW10, KSR14, LTK17, PS10b, PZM^{+}10, PSW11, Puf13, RHT13, SP10a, Sie10, SPS17]. realtime [OUY^{+}13]. Reasoning [LN15, Sun18, ABK^{+}16, MLT17, RKHN18]. Recaf [BIvdS17]. recipes [J^{+}12]. recommendations [LMS^{+}13]. recompilation [NED^{+}13]. Reconfigurable [OUY^{+}13, STY^{+}14, OIA^{+}13]. reconstruction [LSWM16]. Recovering [CRAJ10]. Reducing [MV16, WHIN11]. Reduction [BO12, MSS19, SMP19, TD15]. redundant [HLO15]. reengineering [FGB^{+}19]. Refactoring [AS14, STST12, VBZ^{+}18, ZHL^{+}12, FMM^{+}11, FM13, SZZ^{+}19]. refactoring-aware [SZZ^{+}19]. Reference [Sch14, Sun18, UJR14, HMDE12]. refinement [GY16, JLP^{+}14, KSW^{+}14, MCW19, SNCM19, ZMG^{+}14, ZFK^{+}16]. Reflexes [SPP^{+}10]. region [Ott18]. region-based [Ott18]. regions [AC10]. register [ZYZ^{+}12]. register-based [ZYZ^{+}12]. Regression [MM12]. Regular [CYWD19, NM10b, PIR17]. reification [RRB17]. Reified [GBS14]. Reim [HMDE12]. ReImInfer [HMDE12]. relation [TD15]. relational [MLGA11]. relationship [LSBV16, LSBV17, SH12]. relaxed [DNB^{+}12, KHL^{+}17, PPS16]. relaxed-memory [KHL^{+}17]. Release [Ano14]. reliability [HWLM11]. reliable [LMS^{+}13]. relying [IN12]. Remodularizing [OJ12]. Remote [BVGVEA10, BVGV14a, BJBK12, GSD^{+}15, SS19, BVGVEAFG11]. removal [MRMV12, WGF11]. removing [PLR14]. rename [FM13]. Repair [SEK^{+}19, XMD^{+}17, ZLNP18, MDS^{+}17, SHU16]. repeatability [Vit14]. Replacement [WBHN18, BCD13]. Replay [BH12]. Replaying [WKG17]. replication [CJ17, UIY10]. replication-based [UIY10]. report [CBLFD12, Sch10a]. Reports [OW16]. repository [HC10]. representation [AZLY18]. reproducibility [Vit14]. reproduction [SR14b]. Requirements [MSS19, AGGZ10]. ResAna [KvGS^{+}14]. Research [SR17, TRE^{+}13, CRJ^{+}10, CBLFD12, EKUR10, Rub14, VBMDP16, Vit14]. RESOLVE [Sun18]. Resource [BVGV14a, WZK^{+}19, YPMM12, ADI13, ES14, KvGS^{+}14, KSR14, SGV12]. resource-aware [SGV12]. resource-based [ADI13]. responsive [SPP^{+}10]. responsiveness [PSNS14]. restart [CNS13]. Restructuring [RC17]. results [OD18]. Retention [ZMM^{+}16]. Rethinking [LHR19, Xue12, RCR^{+}14]. retrofitted [TTS^{+}10]. retrofitting [LPGK14]. Reusability [Tai13]. reusable [HC10, MME14]. reuse [WR10]. Reusing [PKPM19]. Reverse [CCA^{+}12, MLM17, MLM19]. Review [Ano15, Ano18, Bro12, Del13, Gve13, Kie13, Ngo12, Teo12, Teo13, EKUR10]. Revisited [Mei14, Gon11]. rewriting [GGC19, HLO15]. rewriting-based [GGC19]. RFID [AYZI10]. RFLP [YCYC12]. richer [CV14]. rigor [Vit14]. Rigorous [AGR17]. rings [Pos19, Pos19]. Rise [DiP18a]. risk [MPM^{+}15]. River [HHSS13]. RJ [OW16]. RMI [SS19]. Road [RXK^{+}17, SWU^{+}15]. Robin [Ano15]. Robotic [DiP18b, LM15]. Robots [SWF12]. Robust [VM15, VDV17, MKZ^{+}14, SGV12, VM10]. Rod [Teo12]. ROM [MLM19]. row [Lei17]. row-typed [Lei17]. RTSJ [ZW10]. Rubah [PVH14]. Ruby [Teo12]. Rule [YPMM12, QLBS17]. Rules [CCA^{+}12, HLO15, PTRV18]. run [WAB^{+}11]. run-time [WAB^{+}11]. Running [HC11, TWX^{+}10, YK14]. runs [FIF^{+}15]. Runtime [BLH12, CMM^{+}10, GSS^{+}18, MAHK16, MSS10, NWB^{+}15, OCFLI14, XMA^{+}14, BRGG12, EQT10, GTL^{+}10, GSS^{+}16, LMK16, MS10, OOK^{+}10, PKC^{+}13, RO12, STY^{+}14, TWSC10, VBAM10a, WLL19, YRHBL13, dCMMN12]. runtimes [BM14, CSV15, RCR^{+}14, WWH^{+}17]. ryu [ALB^{+}19].

S [Gve13]. Safe [Eug13, GvRN^{+}11, JTO12, Loc18, MPS12, RSF^{+}15, SWB^{+}15, WAB^{+}11, HJS^{+}10, HAW13, KHR11, KMLS15, KCP^{+}17, Loc13, RDP16, WWS13]. Safety [MCW19, RS12, SDH^{+}17, WCB16, ZLCW14, AGR17, EKUR10, GMC^{+}13, Nil12b, PG12, SD16b, Taf13, YS10, ZBvdB^{+}19, CWW13, HL13, LWC17, WK12]. Safety-Critical [MCW19, WCB16, ZLCW14, RS12, SDH^{+}17, AGR17, CWW13, LWC17].

Salespoint [ZDS14]. Salt [Hol12]. SAM [BO13]. San [KP15]. Sane [MPS12]. sanitizer [VS11]. Sapphire [URJ18]. SAT [UPR^{+}18]. Satin [VWJB10]. SAW [CFH^{+}13]. Scaffolding [RT14]. Scala [SMS^{+}12, AT16, Hin13, Lew13, PTML11, Pos19, SMSB11, SMS^{+}12]. Scala-Based [PTML11]. Scala.js [DS16]. Scalability [CCH11, VP16, WZK^{+}19, AAB^{+}10, DSEE13, BFS^{+}18, GTSS11]. Scalable [BBB^{+}17, BS12, DFR13, GGRSY17, HC11, JQJ^{+}16, RXK^{+}17, RTE^{+}13, XMA^{+}14, XXCL19, ETTD12, FC11, GGRSY15, NFV15, PIR17, PLR18, RTET15, TTD12]. ScalaLab [PTML11, PMTL14]. scalar [PQTGS17]. Scale [BA17, PE11, CJ19, DHS15, LO15, MDS^{+}17, MCY^{+}10, PTF^{+}15, WHIN11]. SCEL [DLPT14]. scenarios [AMWW15, Sch13]. Scheduler [QSaS^{+}16, IF16, TWL12]. scheduler-independent [IF16]. Scheduling [ASV^{+}16, BVEAGVA10, KPHV11, EP14, EABVGV14, ZW10]. scheme [XHH12, YKA^{+}19]. SCHISM [PZM^{+}10]. Science [HWM11, VF10, SGV12]. sciences [NL14]. Scientific [Esq11, PTML11, TAF^{+}18, WN10, FRGPLF^{+}12, PMTL14]. scientists [Bra14]. SCJ [MCW19]. SCJ-Circus [MCW19]. SCORM [HC10]. Scrap [ZCdSOvdS15]. Script [MSSK16]. Scripting [CSGT17, KKK^{+}17, HBT12, KRR^{+}14, PMTL14, Zha12]. SE [LYBB14]. Seamless [OwKPM15]. Search [NBB18, SED14, WCG^{+}18, XXCL19, DDDF17]. searching [ETR12]. Second [HD17]. secrets [Alt12]. section [DTLM14]. sections [NM10a]. Secure [AFGM15, GMPS12, GM12, ABFM12, LMS^{+}12, LMS^{+}13, TLMM13, WA19]. securely [SFR^{+}14]. securing [CDMR19]. Security [CDG^{+}17, Gon11, HBS16, JWMC15, MCC17, PS10a, STA18]. Seemingly [Has12]. selection [WHIN11].

Self [MPS12, SEPV19, YXS^{+}19, hED12, AHK^{+}11, AGH^{+}17, CBLFD12, HWW^{+}15, MD15]. Self-adaptive [SEPV19]. Self-Admitted [YXS^{+}19]. self-collecting [AHK^{+}11]. self-composition [AGH^{+}17]. self-hosted [CBLFD12]. self-optimizing [HWW^{+}15, MD15]. Self-stabilizing [hED12]. Semantic [GGRSY17, RvB14, BNS12, GGRSY14, GGRSY15, MKK^{+}12, MKK^{+}13, OA17]. Semantics [BO12, BR15, Kri12, LKP19, LML17, SPY^{+}16, AK13, FBH17, FZ17, KHL^{+}17, Mil13, MT14, PSR15, PPS16, ZHCB15]. Semantics-based [SPY^{+}16]. Semantics-driven [LKP19]. semantics-preserving [AK13]. Semi [FM13, SEK^{+}19, ABC18, MRMV12]. semi-automated [MRMV12]. Semi-automatic [FM13]. Semi-Autonomic [SEK^{+}19]. semi-structured [ABC18]. Sensitive [SGD15, HWM13, KRR19, LMK16, LX19, STA18]. sensitivity [HB13, LTMS18, LX19, PLR18]. Sensor [AFGG11, MTSH16]. separability [WRI^{+}10]. Separating [DDM11, AC10]. Separation [ZLNP18, Pha18, TWSC10]. Sequence [NBB18, ZWZ^{+}14]. Sequencing [YWW^{+}18]. Sequent [FFF17]. sequential [BENS12, DMS11]. serialization [MHBO13]. Seriously [Kie10]. Server [HC11, KRH16, D’H12, Dei11, GGC19, HWLM11, R^{+}13]. Server-Side [HC11, KRH16, D’H12]. Service [BVEAGVA10, SDM12, CSKB12, EABVGV14, GD10, HWLM11, KF11]. service-oriented [EABVGV14]. services [MZC10b]. session [KDPG18, FGR12]. Set [SBK13, Lon10a, Lon10b]. Set-based [SBK13, Lon10a, Lon10b]. sets [SP10b]. setters [Mil13]. setting [BDGS13]. Settings [GM12]. Seven [ST15]. SGX [CDMR19]. Shadow [NNTK17, NNT^{+}19].

ShadowVM [MKZ^{+}14]. shalt [LCW18]. shape [GMT14]. Shared [BG17, FBG17, BSMB16]. Shared-Memory [BG17, BSMB16]. sharing [PKO^{+}15]. Sherlock [ADJG19]. Short [AHK^{+}11, Cha18, SV15a, Zak12]. Short-term [AHK^{+}11]. shortcut [MLM19, CSGT17]. Side [Bul18, HC11, OBPM17, D’H12, KRH16]. Side-Channel [Bul18]. SIGCSE [Wal12]. Signatures [DR10]. significance [FMS^{+}11]. Similarity [ADJG19]. simpA [RVP11]. Simple [BO11, BO12, KCP^{+}17, BVGV14b, MSM^{+}10]. Simplicity [Dei11]. Simplifying [Mor18, Ano18]. Simulating [LM15]. Simulation [HWLM11, FLZ^{+}18, KKW11, Rim12, ZXL16]. Simulation-based [HWLM11]. simulations [MCY^{+}10]. Simulator [MKG^{+}17, RXK^{+}17]. single [JK13]. Sinking [CDG^{+}17]. site [CPST15, SSB^{+}14a]. sites [OOK^{+}10]. size [AST12, UTO13]. sizing [CSV15]. SJL [MvH15]. sketching [HZZK19]. skills [JACS10]. Skip [WBHN18]. Skyway [NFN^{+}18]. slices [YSCX17]. Slicing [XMA^{+}14]. Slimming [WGF11]. SLOC [LSBV16, LSBV17]. Smaller [GS12]. smalltalk [FIF^{+}15, HKVG14]. smart [BL15, GMPS12]. Smartcard [RBL12]. SMArtOp [TGZ17]. Smartphones [RT14]. SMARTS [RXK^{+}17]. snapshots [AST12]. Snippets [SWU^{+}15]. SNP [YCYC12]. SoC [TKL^{+}15]. social [GGC18, GGC19]. Socket [WA19]. Soft [WZK^{+}19, JACS10]. Software [BSA14, CC15, KH18, LMZP19, PBM^{+}19, RC17, Wan11, YQTR15, YMHB19, BMSZ17, BTR^{+}13, CBGM12, CFH^{+}13, CJ17, CJ19, CDMR19, DVL13, EKUR10, FRGPLF^{+}12, FC11, GT10a, HBG^{+}16, JhED11, JK11, LPA13, MHR^{+}12, NGB16, OIA^{+}13, PLL^{+}18, PBB19, RAS16, SZZ^{+}19, SV17, XR13, YRHBL13, ZZK13, ZHCB15, ZDS14, CKS18]. Solidity [Dan17]. Solution [KS15, EKUR10, J^{+}12]. Solving [SED14, FMBH15, UPR^{+}18]. Sorting [BKP16]. soul [McM11]. Sound [BO13, BGK17, LE16, SMP19, BHSB14, ELW15, PPMH15, RGB18]. soundly [BS13]. Source [ADJG19, BSA14, GD12, MM16, RLMM15, SRTR17, SED14, ABC18, AK13, CJ17, CJ19, DRN14, EKUR10, FMS^{+}11, JK11, MKK^{+}12, MKK^{+}13, OJ12, PMP^{+}16, SSK13, Tai13, ZWSS15]. source-code [MKK^{+}12, MKK^{+}13]. source-to-source [AK13]. sources [IN12]. Spark [LXP18]. sparse [TGZ17]. sparse-matrix [TGZ17]. spatial [MLGA11]. Speaking [Rau14, Sam12]. Special [DVL13, Fox17a, HL13, HGCA11, Puf13, HTLC10, RHT13, HTW14, VK12]. specialization [KRR^{+}14, SV15a]. specific [CSdL16, EEK^{+}13, HWW^{+}15, Kie13]. Specification [GJS^{+}13, GJS^{+}14, IF16, KW11, LN15, LYBB13a, LYBB13b, LYBB14, MCW19, TWNH12, BVGVEA11a, BCF^{+}14, KR12, KW10, MRA^{+}17, YP10, dCMMN12]. specifications [BSAL18, BENS12, PS10a, TVD10, UPR^{+}18]. specified [BCR11]. Specifying [BNS12, HL13]. Speculation [AC16, MGI14, MGI17]. speculative [BB17, YRHBL13]. speed [HRS^{+}17, SBF^{+}10, UTO13]. SPF’s [PSJ18]. Spi [PS10a]. SPIN [ASdMGM14]. SPL [BTR^{+}13]. splittable [SLF14]. SPOON [PMP^{+}16]. spot [LMK16]. SPUR [BBF^{+}10]. SQL [FGB^{+}19, KMLS15]. SqueakJS [FIF^{+}15]. SSNTDs [VSG17]. Stability [BSA14, LL15]. stabilizing [hED12]. Stack [WBHN18, CDBD18, KRCH14, Xue12]. stack-based [KRCH14]. stage [WRI^{+}10]. staged [SC16]. staging [RO12]. Stakeholders [YMHB19]. Standard [WKG17, LMS^{+}12]. Standardization [TWNH12]. StarL [LM15]. start [WSH^{+}19]. State [AGR12, BLH12, MvDL12, MS14, GN16, YP10]. state- [YP10]. statecharts [MS13]. Statement [XMD^{+}17, PLR14, ZWSS15]. statements [PLR14]. Static [BGK17, BNE16, JC10, LMZP19, MTL15, ODL15, PiLCH11, PLR18, RD15, SW12, Spo16, SBE^{+}19, SNCM19, SH12, AM14, BGOS18, CGJ^{+}16, Fer13, FLL^{+}13, IF16, KSW^{+}14, LS11, MHR^{+}12, PIR17, TLMM13]. statically [BTR^{+}13, NED^{+}13]. statistical [Bra14, ZFK^{+}16]. statistically [PPMH15]. statistics [HCN14]. stealing [KFB^{+}12, TWL12]. STM [CHM16, Sub11]. STM/HTM [CHM16]. StMungo [KDPG18]. stochastic [CRAT^{+}12]. stock [PVH14]. Stop [LWB^{+}15]. stops [BNP^{+}18]. Storage [Hol12, VDV17]. Store [BS12, Sta10]. stores [DFR13]. Story [Ano14]. strategic [BMR14]. strategy [PDPM^{+}16]. Stream [CWGA17, KBPS17, MV16, BRWA14, SSG^{+}14, ZDK^{+}19]. streaming [MRA^{+}17, STCG13]. StreamJIT [BRWA14]. StreamQRE [MRA^{+}17]. streams [SGG^{+}17, UFM15]. Strength [KCD12]. String [HOKO14, CSK17]. Strings [HWM11, HWM10, LSSD14]. strong [UMP10, ZHCB15, ZBB17]. Structure [ZLNP18, LO15, PLL^{+}18, UMP10]. structured [ABC18, LSWM16]. Structures [GT10b, CDTM10, XMA^{+}10]. studies [EKUR10]. Studio [RT14, FH16]. Studio-Based [RT14]. Study [BF18, KB11, LMZP19, OBPM17, RVT18, RLMM15, WZK^{+}19, ZMM^{+}16, BRGG12, CCFB15, CJ17, ECS15, HTBB16, JK11, KFBK^{+}15, MHN19, MHR^{+}12, NCS10, OMK^{+}10, PTF^{+}15, SSL18, SH12, TFPB14, VBDPM16, WXR16, YW13]. studying [CJ19]. style [UFM15]. substitute [PPMH15]. substrate [GTL^{+}10]. subtypes [HL13]. Subtyping [LN15]. Suite [MSS19, SMSB11, BB12]. Suites [GGZ^{+}15]. Summaries [BH17, BA19].

Summarization [MM16, RLMM15]. Superblock [KS13]. Supercharged [Cec11, GBS13]. Superposition [HD17]. supertype [RRB17]. supervenience [Rez12]. Support [ALB^{+}19, CSGT17, KKK^{+}17, RKN^{+}18, BVGVEA13, Cha18, DVL13, GMC^{+}13, Hos12, NGB16, SMN^{+}12]. supported [FMM^{+}11]. Supporting [LVG10, EKUR10]. Surfaces [FBG17]. Surgical [RSB^{+}14]. surprises [FMBH15]. Survey [AGM^{+}17, OAC18, RVT18, BCvC^{+}13, GD10]. SurveyMan [TB14]. surveys [TB14]. suspension [TWL12]. SV [CKS18]. SV-COMP [CKS18]. sweeping [KBL14]. Sweeten [DFHF15]. Swift [ZYZ^{+}12]. SWIM [Sch10a]. symbol [Tar11]. Symbolic [Bul18, CYWD19, NNTK17, NNT^{+}19, PMTP12, SWMV17, MMP^{+}12, Rim12]. synchrobench [Gra15]. Synchronisation [CHMY19, CHMY15, WBM^{+}10]. synchronization [DHM^{+}12, Gra15, Sub11]. Synchronized [BG17]. Synchronized-by-Default [BG17]. synchronizing [DTM^{+}18]. Synchronous [BVEAGVA10, SK12, MvH15]. syntactic [LE16, MKK^{+}12, MKK^{+}13, QLBS17]. Syntax [SS13, KMMV14, SSK13]. synthesis [SR14a, STR16, SS16]. synthesizable [ABCR10]. synthesizer [OUY^{+}13]. Synthesizing [GK15, SRJ15, LWH^{+}10]. Synthetic [PSJ18]. System [BO13, KCD12, MAHK16, ACS^{+}14, AYZI10, AGR17, BDB11, ELW15, HA13, HDK^{+}11, HWLM11, KR12, MS10, STY^{+}14, TLL11, Nil12a]. systematic [TD15]. Systems [BG17, BSA14, BNE16, CCH11, DLPT14, Fox17b, HTW14, JMB12, LM15, LMZP19, MRF18, NFN^{+}18, NWB^{+}18, RTE^{+}13, SLES15, SLE^{+}17, AT16, CJ19, DW10, FH16, Fox17a, HdM17, HWI^{+}12, HTLC10, LPGK14, LTK17, MHR^{+}12, MAH12, MvH15, OIA^{+}13, PLL^{+}18, PdMG12, PBB19, PDPM^{+}16, RHT13, SDH^{+}17, SSMGD10, SABB19, SH12, TTD12, TWX^{+}10, THC^{+}14, UIY10, Vit14, YRHBL13, VK12].

T [HD17]. T-matrix [HD17]. table [Tar11]. Tableau [FFF17]. Tagged [RKN^{+}18]. Tailoring [LZ12]. Take [Kie10]. Taking [SWU^{+}15]. Tales [Sew12]. talk [Piz17, Sie17]. Taming [TLL11, SC16]. Tardis [BM14]. target [Cle16]. Task [RRB19, Fee16, TWL12, ZLB^{+}13]. TaskLocalRandom [PPMH15]. Tasks [PWSG17, PWSG19, ST15, HAW13, PPMH15, SPP^{+}10]. Taurus [MAHK16]. Taxonomy [SS14]. Teaching [GLGA19, HA13, SWF12, CHM13, ZDS14]. teasing [LBF12]. Technical [YXS^{+}19]. technique [SZZ^{+}19, SSK13]. Techniques [LMZP19, RD15, EV13, KS13]. Technologies [Fox17b, HTW14, VK12, Fox17a, HTLC10, KFBK^{+}15, NL14, RHT13]. technology [NED^{+}13]. TeJaS [LPGK14]. Template [MME14, HJS^{+}10]. templates [FOPZ14, AK13]. ten [HTBB16]. term [AHK^{+}11]. Terminating [FFF17]. Termination [BMOG12, RDCP12, BSOG12, SMP10]. Test [AGM^{+}17, BB12, BM18, GGZ^{+}15, LMZP19, MSS19, Pha18, Rim12, SPKT18, ST15, MT13, PSNS14, SR14a, SKR17]. Test-driven [BM18]. tested [Mil13]. Testing [Ame13, BR12, Hin13, McN19, MM12, MMP15, MMP^{+}12, CSS^{+}16, CNS13, KPP^{+}18, Ler10, SABB19, Teo12, TD15, ZBvdB^{+}19]. tests [AÖ11, NYCS12, SRJ15]. Textbooks [BNP11]. their [RDP16]. theorem [SSH17]. There [Esq11]. thin [OD18, PPS16]. thin-air [PPS16]. things [McK16]. Think [WR10]. Third [Ano15, FOPZ14, LVG10]. third-party [FOPZ14, LVG10]. THOR [TWX^{+}10]. Thoth [KB17]. Thou [LCW18]. Thread [MGI14, BKC^{+}13, CRAJ10, MGI17, PCL14, PG12, SS10, WLL19, YDFF15].

Thread-Level [MGI14, MGI17]. threaded [DSEE13, JTO12, SE12, Taf13]. threads [UR15, WLL19]. threat [BGS^{+}13]. threats [BGS^{+}13]. Three [ZMM^{+}16, Vit14]. Tier [WZK^{+}19]. TigerQuoll [BBP13]. Tim [Teo13]. Time [BVEAGVA10, BBB^{+}17, BLH12, DLR16, Fox17b, HTW14, JMB12, Kie10, KW11, PKPM19, Pau14, SLES15, SLE^{+}17, TN19, VK12, BCR13, BM14, BVGVEA10, BVGVEA11a, BVGVEA11b, BVGVEA13, BVGV14a, BVGV14b, CRAJ10, DW10, EABVGV14, Fox17a, GMNC13, GMC^{+}13, HTLC10, KHM^{+}11, KPHV11, KHL^{+}13, KvGS^{+}14, KW10, KSR14, LMK16, LTK17, MGI17, Nil12a, PS10b, PZM^{+}10, PSW11, Puf13, RHT13, SP10a, SPPH10, Sie10, SPS17, SH12, TTS^{+}10, WSH^{+}19, WAB^{+}11]. time-travel [BM14]. time-triggered [EABVGV14]. timed [LKP19]. Times [BKP16, DW10]. timing [AGH^{+}17, LS11]. TIMP [SLS^{+}12]. tiny [Xue12]. To-many [SV18]. to-one [SV18]. Tolerance [RK19]. tolerant [PZM^{+}10]. Tool [FMM^{+}11, NBB18, PQD12, SW12, SSK13, ABFM12, CRAT^{+}12, ETR12, KSR14, LS11, TWX^{+}10]. Tool-supported [FMM^{+}11]. toolchain [KDPG18, SMN^{+}18]. Toolkit [FBG17]. Tools [Bro12, CSZ17, CS12, CKS18, ABK^{+}16, KPP^{+}18, VBAM10b]. toolset [KvGS^{+}14]. top [RVP11, SGG^{+}17, ZMNY14]. top- [SGG^{+}17]. top-down [ZMNY14]. Topics [Hor11, Jen12]. topology [DDM11]. touch [MTSH16]. TouchSignatures [MTSH16]. Toy [DiP18b]. Trace [HWM14, PiLCH11, SR14b, BBF^{+}10, HWM13, HWI^{+}12, IHWN12, WHIN11]. trace-based [BBF^{+}10, HWM14, HWI^{+}12, IHWN12]. Traceability [CSKB12]. tracer [CZ14]. Traces [WKG17, BA12, RGM13]. Tracing [BP10, DLR14, DLR16, MAK19, MRF18, MD15]. track [VSG17]. TrackEtching [VSG17]. Tracking [OAC18, RLMM15, SDC^{+}12, WLL19, KHL^{+}13, OOK^{+}10]. Tracks [RGM13]. tradeoff [UTO13]. Traffic [RXK^{+}17]. Trail [HHSS13]. Train [MSSK16]. training [KMZN16]. trait [BCD13, VM15]. traits [BDGS13, BD17]. Transactional [URJ18, DVL13, FC11, ZHCB15]. Transactions [DcSG12, CHM16, DFR13]. transfer [BL15]. transformation [AST^{+}16, PDDD17]. transformations [AK13, MHM10, PMP^{+}16, TL17]. Transforming [dMRH12]. transitioning [HWM14]. Translating [RFRS14]. Translation [BO12, LSWM16, LXP18, TJLL18]. translations [UTO13]. translator [LZYP16]. Translators [WWG^{+}18]. Transmission [PE11, BVGVEA11b, BJBK12]. transparent [BDB11]. transpiler [STA18]. travel [BM14]. traversals [ODL15]. Treble [YMHB19]. Tree [Lyo12, HLO15, KMMV14, SSK13, YKA^{+}19]. trees [RBV16]. Trends [CC15, MSS10, SR17]. trie [SV17]. trie-based [SV17]. tries [SV15a, SV15b, SV18]. triggered [EABVGV14]. triggers [FGB^{+}19]. TRINI [PDPM^{+}16]. Trusted [TWNH12, BCF^{+}14]. TUIOFX [FBG17]. tuning [AAB^{+}10, BVGVEAFG11, SKBL11]. Turf [CH17]. Turing [Gri17]. Turn [HOSC16]. Tutorial [Jen12, Nil12b, PBM^{+}19, Taf13, Zak12]. TV [JMO14]. Twitter [Guy14]. Two [Has12]. Type [BO13, CGJ^{+}16, KSW^{+}14, KATS12, Lei17, Loc18, RKN^{+}18, SGD15, WT11, ACS^{+}14, AT16, BS13, CMS^{+}12, CVG^{+}17, DLM10, FH16, GBS14, HyG12, KMLS15, KRR^{+}14, KRH16, KvRHA14, KDPG18, LPGK14, LE16, MHR^{+}12, SV18, SH12, TLL11, Zha12, eBH11]. Type-Based [SGD15]. type-dependent [LE16]. type-heterogeneous [SV18]. Type-Safe [Loc18, KMLS15]. Typechecking [KDPG18, CL17]. Typed [BO13, KKK^{+}17, MHL15, CMS^{+}12, KRCH14, Lei17, RDP16]. Types [BO13, RvB14, SPAK10, BDGS13, CHJ12, DDM11, HH13, MME^{+}10, YDFF15]. TypeScript [Cho14, FH16, RSF^{+}15]. Typing [FZ17, RSF^{+}15, Sie17, SFR^{+}14, TSD^{+}12]. typy [OA17].

Ubiquitous [MCY^{+}10]. UDP [RR14]. ulfjack [ALB^{+}19]. ulfjack/ryu [ALB^{+}19]. ULS [FOPZ14]. ultimate [BL15]. UML [CSF^{+}16]. unbounded [LSSD14, RGB18]. uncertain [McK16]. Unchangeable [RK19]. Understandable [MSM^{+}16]. Understanding [ABC18, FRM^{+}15, MKTD17, NWB^{+}18, PCL14, QLBS17, Set13, TABS12, VBMDP16, LWB^{+}15, Nil12b, OD18]. Undocumented [Alt12, MHR^{+}12]. Unified [LM15]. uniform [AH10, Eug13]. Unifying [Has12, MKK^{+}12, MKK^{+}13]. union [KT15]. uniprocessors [KPHV11]. Units [LLL13]. universe [DDM11]. Unix [PVB17]. Unobtrusive [MGS19]. Unpicking [LBF12]. Unrestricted [WWS13]. unsafe [MPM^{+}15]. unsound [AT16]. Updates [YMHB19, PKC^{+}13]. Upper [SW12]. Upsortable [SGG^{+}17]. uptrees [HB13]. USA [Hol12, KP15]. usability [FH16, MHR^{+}12, WA19]. Usage [OAC18, RC17, PTF^{+}15, QLBS17]. Use [BGK17, Guy14, MPM^{+}15, AMWW15, MKTD17, PBMH13, Sch13]. use-case [AMWW15]. used [XR10]. useless [FRC^{+}17]. User [Liu14, MvDL12, RKHN18, SLS^{+}12, DAA13, FMS^{+}11, MTSH16, PSNS14]. user-defined [FMS^{+}11]. User-guided [RKHN18]. Using [ASdMGM14, BS12, BSA14, BNE16, DLM10, GLGA19, HCN14, KFBK^{+}15, KH18, MV16, MSSK16, NBB18, Pau14, PQD12, RC17, SDM12, SLE^{+}17, UMP10, Wan11, WKG17, WCG^{+}18, XMA^{+}14, YCYC12, Zak18, BB17, Dan18, DDDF17, Del13, FH16, FOPZ14, GBS14, IvdS16, KMLS15, KT14, KC12, LVG10, Lew13, LDL14, MT13, PIR17, PLR18, Pha18, RKHN18, RAS16, SAdB^{+}16, SSK13, SSH17, SHU16, SS19, VGS14, WLL19, WBM^{+}10, WRI^{+}10, XR13, ZLNP18, vdMvdMV12]. UT [Hol12]. utility [CSV15, XMA^{+}10]. utilization [BCR13].

v [Sam12]. V8 [MGI17]. Validating [HLSK13]. Validation [SSB14b, CSdL16, GMNC13, HCV17, SSB01]. Value [BBB^{+}17, DFR13, SNCM19, YSCX17]. value-flow [YSCX17]. variable [CDTM10]. variables [NS13]. VDM [TJLL18]. Verifiable [FHSR12]. Verification [CYWD19, CHMY19, CKS18, KKW14, KP15, RAS16, SS12, SSB14b, CHMY15, DLM10, HCV17, PSW11, SMN^{+}18, SZ11, SJPS10, SSH17, SSB01, dCMMN12]. verification-validation [HCV17]. Verified [HM12, Loc18, JLP^{+}14, WFF18]. VerifiedFT [WFF18]. Verifier [BDT10, Rey13]. verifiers [SPY^{+}16]. Verifying [LM15, YS10, vdMvdMV12, SD16b]. Veritesting [SWMV17]. Version [FLZ^{+}18, FC11, HD17, SM12, TMVB13, ZXL16]. vertical [BFS^{+}18, STY^{+}14]. via [Bul18, DMS11, GGRSY15, GGRSY17, Hos12, HB13, JWMC15, LSWM16, MTSH16, Rim12, SS16, TD17]. Video [PBM^{+}19]. view [Guy14]. violations [LTZ14, PG12, RDF15]. Virtual [BZD17, Cle16, LYBB13a, LYBB13b, LYBB14, LTK17, PTHH14, PQD12, RRB19, SSB^{+}14a, Sch13, Set13, SMSB11, SGV12, SSB01, SSB14b, UR15, Ame13, CBLFD12, KRCH14, McM11, NK10, Piz17, RCB17, SSMGD10, WGF11, WHV^{+}13]. virtualized [HOKO14, MHM10]. virus [RBL12]. Vision [DiP18b, HCV17]. Vision-Guided [DiP18b]. visitors [DRN14]. Visual [FH16]. Visualization [TAF^{+}18, JEC^{+}12, JJL17, MCY^{+}10]. visualizing [DSEE13, KS14, MPR12]. vital [EV13]. VM [AFGM15, LBF12, WBHN18, YKM17]. VM/application [LBF12]. VMKit [GTL^{+}10]. VMs [SEPV19]. volatile [WZL^{+}18]. Volatility [WZL^{+}18]. volume [Gve13]. Vroom [BMDK15]. vs [BA17, GBC12, MD15, SRTR17, SK12, SH12, WKJ17]. Vulnerabilities [MS14, GGC18]. vulnerability [MLM19, Sve14].

Wampler [Bro12]. wanted [Gra15].

watering [Ano13]. wave [PQTGS17]. way [Ker15, PLR18, WGF11]. ways [Kie13]. weak [WRI^{+}10]. Weakening [WRPP19]. Weakly [KRR19]. Weapon [Nil12a]. weaving [VBMA11]. web [AMT17, EKUR10, ETR12, HRS^{+}17, HCLH18, HCN14, KFBK^{+}15, MCC17, MCY^{+}10, RHSD15, RCR^{+}14, Ryu16, VB18, WGW^{+}11, DAA13, HLSK13, Kri12, LYM^{+}18, MGI14, MvDL12, MMP15, NL14, OwKPM15, RFBJ14, RPP19, Sch10b, VP16, YW13, Zak18]. web-based [EKUR10]. web-portal [MCY^{+}10]. WebAssembly [HRS^{+}17, WRPP19]. WebCL [KFBK^{+}15]. Websites [KCD12]. weight [MvH15]. weighted [PLL^{+}18]. Weka [SS19]. well [EV13]. well-grounded [EV13]. WETSUIT [ETR12]. Whalesong [YK14]. whole [DS16]. whole-program [DS16]. Widening [KKW14]. Wild [RPP19, MPM^{+}15, Ryu16, STS^{+}13]. wildcards [AS14, TLL11]. WIP [Cha18]. Wireless [AFGG11]. Wise [SEPV19]. Withers [Lyo12]. without [FMBH15, IN12, KFB^{+}12, SS12, Sta10, WHIN11]. Word [SRTR17]. Work [KFB^{+}12, PKO^{+}15, TWL12]. Work-stealing [KFB^{+}12, TWL12]. workarounds [UPR^{+}18]. workbench [CFH^{+}13]. Workers [VP16]. Working [ST15]. workshop [Fox17a]. world [CIAD13, McK16, STS^{+}13]. worms [GGC19]. Worst [SPPH10, dGRdB^{+}15]. Worst-case [SPPH10]. would [Han15]. wrap [FOPZ14]. Wrappers [MPS12]. Wright [Teo13]. Write [ASME18, HJH10]. Write-rationing [ASME18]. Writing [HOSC16, Jaf13, Mor18].

x [MSM^{+}16]. X10 [TWL12]. Xbase [EEK^{+}13]. XIR [TWSC10]. XML [NL14]. XSS [GGC18, GGC19, MSSK16, VS11]. Xtraitj [BD17].

yang [CBGM12]. years [BTR^{+}13, HTBB16]. yieldpoint [LWB^{+}15]. yin [CBGM12].

Z [SBF^{+}10]. Z-rays [SBF^{+}10]. Zero [ZW13].

References

Altman:2010:OTJ

[AAB^{+}10] E. Altman, M. Arnold, R. Bordawekar, R. M. Delmonico, N. Mitchell, and P. F. Sweeney. Observations on tuning a Java enterprise application for performance and scalability. IBM Journal of Research and Development, 54(5):2:1–2:12, ???? 2010. CODEN IB-MJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Acar:2018:PCM

[AACR18] Umut A. Acar, Vitaly Aksenov, Arthur Charguéraud, and Mike Rainey. Performance challenges in modular parallel programs. ACM SIGPLAN Notices, 53(1): 381–382, January 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Accioly:2018:USS

[ABC18] Paola Accioly, Paulo Borba, and Guilherme Cavalcanti. Understanding semi-structured merge conflict characteristics in open-source Java projects. Empirical Software Engineering, 23(4): 2051–2085, August 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/article/10.1007/s10664-
017-9586-1.

Auerbach:2010:LJC

[ABCR10] Joshua Auerbach, David F. Bacon, Perry Cheng, and Rodric Rabbah. Lime: a Java-compatible and synthesizable language for heterogeneous architectures. ACM SIGPLAN Notices, 45 (10):89–108, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avvenuti:2012:JTC

[ABFM12] Marco Avvenuti, Cinzia Bernardeschi, Nicoletta De Francesco, and Paolo Masci. JCSI: a tool for checking secure information flow in Java Card applications. The Journal of systems and software, 85(11):2479–2493, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121212001513.

Abanades:2016:DAR

[ABK^{+}16] Miguel Abánades, Francisco Botana, Zoltán Kovács, Tomás Recio, and Csilla Sólyom-Gecse. Development of automatic reasoning tools in GeoGebra. ACM Communications in Computer Algebra, 50(3):85–88, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ansaloni:2012:DAO

[ABMV12] Danilo Ansaloni, Walter Binder, Philippe Moret, and Alex Villazón. Dynamic aspect-oriented programming in Java: The Hot-Wave experience. Lecture Notes in Computer Science, 7271:92–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/

                                        10.1007/978-3-642-35551-
                                        6_3/.

Akai:2010:EAS

[AC10] Shumpei Akai and Shigeru Chiba. Extending AspectJ for separating regions. ACM SIGPLAN Notices, 45(2):45–54, February 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anjo:2016:DML

[AC16] Ivo Anjo and João Cachopo. Design of a method-level speculation framework for boosting irregular JVM applications. Journal of Parallel and Distributed Computing, 87(??):13–25, January 2016. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0743731515001720.

Ahn:2014:IJP

[ACS^{+}14] Wonsun Ahn, Jiho Choi, Thomas Shull, María J. Garzarán, and Josep Torrellas. Improving JavaScript performance by deconstructing the type system. ACM SIGPLAN Notices, 49(6):496–507, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aumuller:2016:OPD

[AD16] Martin Aumüller and Martin Dietzfelbinger. Optimal partitioning for dual-pivot Quicksort. ACM Transactions on Algorithms, 12(2): 18:1–18:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amighi:2016:PCC

[AdCGGH16] Afshin Amighi, Pedro de Carvalho Gomes, Dilian Gurov, and Marieke Huisman. Provably correct control flow graphs from Java bytecode programs with exceptions. International Journal on Software Tools for Technology Transfer (STTT), 18(6): 653–684, November 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s10009-015-0375-
0;   http://link.springer.
com/article/10.1007/s10009-
015-0375-0.

Autili:2013:HAR

[ADI13] Marco Autili, Paolo Di Benedetto, and Paola Inverardi. A hybrid approach for resource-based comparison of adaptable Java applications. Science of Computer Programming, 78(8):987– 1009, August 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642312000147.

Allyson:2019:SOI

[ADJG19] F. B. Allyson, M. L. Danilo, S. M. José, and B. C. Giovanni. Sherlock N-overlap: Invasive normalization and overlap coefficient for the similarity analysis between source code. IEEE Transactions on Computers, 68 (5):740–751, May 2019. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Almeida:2019:GPD

[AdSCdR^{+}19] André Luís Barroso Almeida, Leonardo de Souza Cimino, José Estevão Eugênio de Resende, Lucas Henrique Moreira Silva, Samuel Queiroz Souza Rocha, Guilherme Aparecido Gregorio, Gustavo Silva Paiva, Saul Delabrida, Haroldo Gambini Santos, Marco Antonio Moreira de Carvalho, Andre Luiz Lins Aquino, and Joubert de Castro Lima. A general-purpose distributed computing Java middleware. Concurrency and Computation: Practice and Experience, 31(7):e4967:1– e4967:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Austin:2012:MFD

[AF12] Thomas H. Austin and Cormac Flanagan. Multiple facets for dynamic information flow. ACM SIG-PLAN Notices, 47(1):165– 178, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:2011:AOJ

[AFG^{+}11] Matthew Arnold, Stephen Fink, David Grove, Michael Hind, and Peter F. Sweeney. Adaptive optimization in the Jalapeno JVM. ACM SIGPLAN Notices, 46(4): 65–83, April 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aiello:2011:JBA

[AFGG11] Francesco Aiello, Giancarlo Fortino, Raffaele Gravina, and Antonio Guerrieri. A Java-based agent platform for programming wireless sensor networks. The Computer Journal, 54(3):439– 454, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/439.
full.pdf+html.

Anikeev:2015:SGC

[AFGM15] Maxim Anikeev, Felix C.

Freiling, Johannes Götzfried, and Tilo Müller. Secure garbage collection: Preventing malicious data harvesting from deallocated Java objects inside the Dalvik VM. Journal of Information Security and Applications (JISA), 22(??):81–86, June 2015. CODEN ???? ISSN 2214-2126. URL http:/

/www.sciencedirect.com/
science/article/pii/S2214212614001367.

Albert:2010:PIM

[AGGZ10] Elvira Albert, Samir Genaim, and Miguel Gómez-Zamalloa. Parametric inference of memory requirements for garbage collected languages. ACM SIGPLAN Notices, 45 (8):121–130, August 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antonopoulos:2017:DIS

[AGH^{+}17] Timos Antonopoulos, Paul Gazzillo, Michael Hicks, Eric Koskinen, Tachio Terauchi, and Shiyi Wei. Decomposition instead of selfcomposition for proving the absence of timing channels. ACM SIGPLAN Notices, 52(6):362–375, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andreasen:2017:SDA

[AGM^{+}17] Esben Andreasen, Liang Gong, Anders Møller, Michael Pradel, Marija Selakovic, Koushik Sen, and Cristian-Alexandru Staicu. A survey of dynamic analysis and test generation for JavaScript. ACM Computing Surveys, 50(5):66:1–66:??, November 2017. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Arcaini:2012:CCM

[AGR12] Paolo Arcaini, Angelo Gargantini, and Elvinia Riccobene. CoMA: Conformance monitoring of Java programs by abstract state machines. Lecture Notes in Computer Science, 7186: 223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-29860-
8_17/.

Arcaini:2017:RDP

[AGR17] Paolo Arcaini, Angelo Gargantini, and Elvinia Riccobene. Rigorous development process of a safetycritical system: from ASM models to Java code. International Journal on Software Tools for Technology Transfer (STTT), 19 (2):247–269, April 2017. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

http://link.springer.

                                        com/accesspage/article/
                                        10.1007/s10009-015-0394-
                                        x;   http://link.springer.
                                        com/article/10.1007/s10009-
                                        015-0394-x.

Apel:2010:CUF

[AH10] Sven Apel and Delesley Hutchins. A calculus for uniform feature composition. ACM Transactions on Programming Languages and Systems, 32(5):19:1– 19:33, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Aigner:2011:STM

[AHK^{+}11] Martin Aigner, Andreas Haas, Christoph M. Kirsch, Michael Lippautz, Ana Sokolova, Stephanie Stroka, and Andreas Unterweger. Short-term memory for selfcollecting mutators. ACM SIGPLAN Notices, 46(11): 99–108, November 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’11 conference proceedings.

Aigner:2015:AJE

[AHK^{+}15] Martin Aigner, Thomas Hütter, Christoph M. Kirsch, Alexander Miller, Hannes Payer, and Mario Preishuber. ACDC-JS: explorative benchmarking of JavaScript memory management. ACM SIGPLAN Notices, 50(2):67–78, February 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andrysco:2016:PFP

[AJL16] Marc Andrysco, Ranjit Jhala, and Sorin Lerner. Printing floating-point numbers: a faster, always correct method. ACM SIG-PLAN Notices, 51(1):555– 567, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Axelsen:2013:PTD

[AK13] Eyvind W. Axelsen and Stein Krogdahl. Package Templates: a definition by semantics-preserving source-to-source transformations to efficient Java code. ACM SIGPLAN Notices, 48(3):50–59, March 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:2019:URP

[ALB^{+}19] Ulf Adams, Stephan T.

Lavavej, Alexander Bolz, Vinnie Falco, David Tolnay, Mitchell Blank, Jr., Mara Bos, Caleb Spare, and Alexander Iljin. ulfjack/ryu: Printf support. Web site, August 12, 2019.

URL  https://github.
com/ulfjack/ryu/tree/

                                        v2.0;   https://zenodo.
                                        org/record/3366212.

Altman:2012:USM

[Alt12] Yair M. Altman. Undocumented secrets of MATLAB– Java programming. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-6904-9 (e-book), 1-4398-6903-0 (hardback), 1-4398-6903-0. xxi + 663 + 16 pp. LCCN QA297 .A544 2012.

Andreasen:2014:DSA

[AM14] Esben Andreasen and Anders Møller. Determinacy in static analysis for jQuery. ACM SIGPLAN Notices, 49(10):17–31, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ament:2013:ATG

[Ame13] John Ament. Arquillian Testing Guide: get familiarized with the Arquillian framework and its capabilities to carry out integration and functional testing on a Java virtual machine. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78216-070-1. v + 224 pp. LCCN QA76.73.J3.

                                        URL http://proquest.
                                        tech.safaribooksonline.
                                        de/9781782160700.

Adamsen:2017:PIR

[AMT17] Christoffer Quist Adamsen, Anders Møller, and Frank Tip. Practical initialization race detection for JavaScript web applications. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):66:1–66:??, October 2017. CODEN ???? ISSN 2475-1421.

Ashrov:2015:UCB

[AMWW15] Adiel Ashrov, Assaf Marron, Gera Weiss, and Guy Wiener. A use-case for behavioral programming: an architecture in JavaScript and Blockly for interactive applications with crosscutting scenarios. Science of Computer Programming, 98 (part 2)(??):268–292, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S016764231400032X.

Andersen:2014:PLJ

[And14] J. S. Andersen. Privacy leaks in Java classes. Ada User Journal, 35(3):191–??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AMJ

[Ano12] Anonymous. Another major Java panic. Network Security, 2012(9):19–20, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1353485812700843.

Anonymous:2013:FAM

[Ano13] Anonymous. Facebook, Apple and Microsoft hit by Java exploit at watering hole. Network Security, 2013(3):1–2, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1353485813700357.

Anonymous:2014:RKS

[Ano14] Anonymous. Release the kraken: a story of Node.js in the enterprise (PayPal). ACM Queue: Tomorrow’s Computing Today, 12(2):80, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2015:BRL

[Ano15] Anonymous. Book review: Learning PHP, MySQL, JavaScript, CSS & HTML5, by Robin Nixon. Third Edition. O’Reilly. ISBN 978-1-4919-4946-7. Network Security, 2015(1):4, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1353485815700061.

Anonymous:2018:BRS

[Ano18] Anonymous. Book review:

Simplifying JavaScript, by Joe Morgan. Published by Pragmatic Bookshelf. ISBN: 978-1-68050-288-6. Price: $39.95, 282 pgs, paperback. Network Security, 2018(5): 5, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1353485818300412.

Arslan:2011:JPM

[AÖ11] Güvenç Arslan and Ilknur Özmen. A JAVA program for the multivariate Z_{p} and C_{p} tests and its application. Journal of Computational and Applied Mathematics, 235(16):4729–4735, June 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271000484X.

Altidor:2014:RJG

[AS14] John Altidor and Yannis Smaragdakis. Refactoring Java generics by inferring wildcards, in practice. ACM SIGPLAN Notices, 49(10):271–290, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adalid:2014:USA

[ASdMGM14] Damián Adalid, Alberto Salmerón, María del Mar Gallardo, and Pedro Merino. Using SPIN for automated debugging of infinite executions of Java programs. The Journal of systems and software, 90(??):61–75, April 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121213002641.

Austin:2017:MFD

[ASF17] Thomas H. Austin, Tommy Schmitz, and Cormac Flanagan. Multiple facets for dynamic information flow with exceptions. ACM Transactions on Programming Languages and Systems, 39 (3):10:1–10:??, July 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Akram:2018:WRG

[ASME18] Shoaib Akram, Jennifer B. Sartor, Kathryn S. McKinley, and Lieven Eeckhout. Write-rationing garbage collection for hybrid memories. ACM SIGPLAN Notices, 53 (4):62–77, April 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Afek:2012:ISJ

[AST12] Yehuda Afek, Nir Shavit, and Moran Tzafrir. Interrupting snapshots and the Java size method. Journal of Parallel and Distributed Computing, 72(7): 880–888, July 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373151200072X.

Alshara:2016:MLO

[AST^{+}16] Zakarea Alshara, Abdelhak-Djamel Seriai, Chouki Tibermacine, Hinde Lilia Bouziane, Christophe Dony, and Anas Shatnawi. Migrating large object-oriented applications into componentbased ones: instantiation and inheritance transformation. ACM SIGPLAN Notices, 51(3):55–64, March 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Akram:2016:BPG

[ASV^{+}16] Shoaib Akram, Jennifer B.

Sartor, Kenzo Van Craeynest, Wim Heirman, and Lieven Eeckhout. Boosting the priority of garbage: Scheduling collection on heterogeneous multicore processors. ACM Transactions on Architecture and Code Optimization, 13(1):4:1–4:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Amin:2016:JST

[AT16] Nada Amin and Ross Tate. Java and Scala’s type systems are unsound: the existential crisis of null pointers. ACM SIGPLAN Notices, 51(10):838–848, October 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ali:2010:DJB

[AYZI10] Mohammed F. M. Ali, Mohammed I. Younis, Kamal Z. Zamli, and Widad Ismail. Development of Java based RFID application programmable interface for heterogeneous RFID system. The Journal of systems and software, 83(11): 2322–2331, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alon:2018:GPB

[AZLY18] Uri Alon, Meital Zilberstein, Omer Levy, and Eran Yahav. A general pathbased representation for predicting program properties. ACM SIGPLAN Notices, 53(4):404–419, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alimadadi:2018:FBP

[AZMT18] Saba Alimadadi, Di Zhong, Magnus Madsen, and Frank Tip. Finding broken promises in asynchronous JavaScript programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 162:1–162:26, October 2018.

URL  https://dl.acm.
org/doi/abs/10.1145/3276532.

Bradel:2012:ITJ

[BA12] Bory J. Bradel and Tarek S.

Abdelrahman. Inlining with traces in Java programs. International Journal of Computer Systems Science and Engineering, 27(4):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Brown:2017:NJP

[BA17] Neil C. C. Brown and Amjad Altadmri. Novice Java programming mistakes: Large-scale data vs. educator beliefs. ACM Transactions on Computing Education, 17(2):7:1–7:??, June 2017. CODEN ???? ISSN 1946-6226.

Berglund:2019:MSJ

[BA19] Lasse Berglund and Cyrille Artho. Method summaries for JPF. ACM SIG-SOFT Software Engineering Notes, 44(4):16, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://

dl.acm.org/doi/10.1145/
3364452.33644560.

Boland:2012:JCC

[BB12] Tim Boland and Paul E. Black. Juliet 1.1 C/C++ and Java Test Suite. Computer, 45(10):88–90, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bonetta:2017:FJF

[BB17] Daniele Bonetta and Matthias Brantner. FAD.js: fast JSON data access using JIT-based speculative optimizations. Proceedings of the VLDB Endowment, 10(12):1778–1789, August 2017. CODEN ???? ISSN 2150-8097.

Basin:2017:KKV

[BBB^{+}17] Dmitry Basin, Edward Bortnikov, Anastasia Braginsky, Guy Golan-Gueta, Eshcar Hillel, Idit Keidar, and Moshe Sulamy. KiWi: a key–value map for scalable real-time analytics. ACM SIGPLAN Notices, 52(8):357–369, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bebenita:2010:STB

[BBF^{+}10] Michael Bebenita, Florian Brandner, Manuel Fahndrich, Francesco Logozzo, Wolfram Schulte, Nikolai Tillmann, and Herman Venter. SPUR: a tracebased JIT compiler for CIL. ACM SIGPLAN Notices, 45(10):708–725, October 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonetta:2013:TPE

[BBP13] Daniele Bonetta, Walter Binder, and Cesare Pautasso. TigerQuoll: parallel event-based JavaScript. ACM SIGPLAN Notices, 48 (8):251–260, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Bu:2013:BAD

[BBXC13] Yingyi Bu, Vinayak Borkar, Guoqing Xu, and Michael J. Carey. A bloat-aware design for big data applications. ACM SIGPLAN Notices, 48(11):119–130, November 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’13 conference proceedings.

Bettini:2013:FDT

[BCD13] Lorenzo Bettini, Sara Capecchi, and Ferruccio Damiani. On flexible dynamic trait replacement for Java-like languages. Science of Computer Programming, 78(7): 907–932, July 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167642312002092.

Bodin:2014:TMJ

[BCF^{+}14] Martin Bodin, Arthur Chargueraud, Daniele Filaretti, Philippa Gardner, Sergio Maffeis, Daiva Naudziuniene, Alan Schmitt, and Gareth Smith. A trusted mechanised JavaScript specification. ACM SIG-PLAN Notices, 49(1):87– 100, January 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). POPL ’14 conference proceedings.

Bergenti:2011:PPS

[BCR11] F. Bergenti, L. Chiarabini, and G. Rossi. Programming with partially specified aggregates in Java. Computer Languages, Systems and Structures, 37(4):178– 192, October 2011. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1477842411000169.

Bacon:2013:PRT

[BCR13] David F. Bacon, Perry Cheng, and V. T. Rajan. POPL 2003: a realtime garbage collector with low overhead and consistent utilization. ACM SIG-PLAN Notices, 48(4S):58– 71, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bainomugisha:2013:SRP

[BCvC^{+}13] Engineer Bainomugisha, Andoni Lombide Carreton, Tom van Cutsem, Stijn Mostinckx, and Wolfgang de Meuter. A survey on reactive programming. ACM Computing Surveys, 45(4): 52:1–52:??, August 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bettini:2017:XTJ

[BD17] Lorenzo Bettini and Ferruccio Damiani. Xtraitj: Traits for the Java platform. The Journal of systems and software, 131(??):419–441, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121216301297.

Bala:2011:DTD

[BDB11] Vasanth Bala, Evelyn Duesterwald, and Sanjeev Banerjia. Dynamo: a transparent dynamic optimization system. ACM SIG-PLAN Notices, 46(4):41– 52, April 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bettini:2013:CTB

[BDGS13] Lorenzo Bettini, Ferruccio Damiani, Kathrin Geilmann, and Jan Schäfer. Combining traits with boxes and ownership types in a Java-like setting. Science of Computer Programming, 78(2):218–247, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642311001833.

Barbuti:2010:AIA

[BDT10] Roberto Barbuti, Nicoletta De Francesco, and Luca Tesei. An abstract interpretation approach for enhancing the Java Bytecode Verifier. The Computer Journal, 53 (6):679–700, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/679;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/679;   http:
//www.oxfordjournals.
org/our_journals/computer_
journal/wilkes_award.

html. This article is the winner of The Computer Journal Wilkes Award for 2010.

Burnim:2012:NIN

[BENS12] Jacob Burnim, Tayfun Elmas, George Necula, and Koushik Sen. NDetermin: inferring nondeterministic sequential specifications for parallelism correctness. ACM SIG-PLAN Notices, 47(8):329– 330, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Bruno:2018:SGC

[BF18] Rodrigo Bruno and Paulo Ferreira. A study on garbage collection algorithms for big data environments. ACM Computing Surveys, 51(1):20:1– 20:??, April 2018. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bruno:2018:DVM

[BFS^{+}18] Rodrigo Bruno, Paulo Ferreira, Ruslan Synytsky, Tetiana Fydorenchyk, Jia Rao, Hang Huang, and Song Wu. Dynamic vertical memory scalability for OpenJDK cloud applications. ACM SIGPLAN Notices, 53(5): 59–70, May 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Battig:2017:SDC

[BG17] Martin Bättig and Thomas R. Gross. Synchronized-by-default concurrency for shared-memory systems. ACM SIGPLAN Notices, 52 (8):299–312, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:2017:EUS

[BGK17] Lewis Berman, Keith Gallagher, and Suzanne Kozaitis. Evaluating the use of sound in static program comprehension. ACM Transactions on Applied Perception, 15(1):7:1–7:??, November 2017. CODEN ???? ISSN 1544-3558 (print), 1544-3965 (electronic).

Blackshear:2018:RCS

[BGOS18] Sam Blackshear, Nikos Gorogiannis, Peter W. O’Hearn, and Ilya Sergey. RacerD: compositional static race detection. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 144:1–144:28, October 2018.

                                        URL  https://dl.acm.
                                        org/doi/abs/10.1145/3276514.

Bedi:2013:MMT

[BGS^{+}13] Punam Bedi, Vandana Gandotra, Archana Singhal, Himanshi Narang, and Sumit Sharma. Mitigating multithreats optimally in proactive threat management. ACM SIGSOFT Software Engineering Notes, 38(1):1– 7, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bodden:2010:AOR

[BH10] E. Bodden and K. Havelund.

Aspect-oriented race detection in Java. IEEE Transactions on Software Engineering, 36(4):509–527, July/August 2010. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
5406531.

Barbu:2012:ARA

[BH12] Guillaume Barbu and Philippe Hoogvorst. Applicationreplay attack on Java cards: When the garbage collector gets confused. Lecture Notes in Computer Science, 7159:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-28166-
2_1/.

Badihi:2017:CAG

[BH17] Sahar Badihi and Abbas Heydarnoori. CrowdSummarizer: Automated generation of code summaries for Java programs through crowdsourcing. IEEE Software, 34(2):71–80, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        so/2017/02/mso2017020071-
                                        abs.html.

Biswas:2014:DES

[BHSB14] Swarnendu Biswas, Jipeng Huang, Aritra Sengupta, and Michael D. Bond. DoubleChecker: efficient sound and precise atomicity checking. ACM SIGPLAN Notices, 49(6):28–39, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Biboudis:2017:RJD

[BIvdS17] Aggelos Biboudis, Pablo Inostroza, and Tijs van der Storm. Recaf: Java dialects as libraries. ACM SIGPLAN Notices, 52(3):2– 13, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burdette:2012:ECJ

[BJBK12] Philip F. Burdette, William F. Jones, Brian C. Blose, and Gregory M. Kapfhammer. An empirical comparison of Java remote communication primitives for intranode data transmission. ACM SIGMETRICS Perform. Eval. Rev., 39(4):2– 11, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baar:2012:DEP

[BK12] Thomas Baar and Philipp Kumar. Detecting entry points in Java libraries. Lecture Notes in Computer Science, 7162: 42–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-29709-
0_6/.

Bell:2014:PID

[BK14] Jonathan Bell and Gail Kaiser. Phosphor: illuminating dynamic data flow in commodity JVMs. ACM SIGPLAN Notices, 49(10): 83–101, October 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2013:OCC

[BKC^{+}13] Michael D. Bond, Milind Kulkarni, Man Cao, Minjia Zhang, Meisam Fathi Salmi, Swarnendu Biswas, Aritra Sengupta, and Jipeng Huang. OCTET: capturing and controlling crossthread dependences efficiently. ACM SIGPLAN Notices, 48(10):693–712, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Brooks:2016:CST

[BKP16] Andrew Brooks, Laura Krebs, and Brandon Paulsen. A comparison of sorting times between Java 8 and Parallel Colt: an exploratory experiment. ACM SIGSOFT Software Engineering Notes, 41(4):1– 5, July 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bouffard:2015:UCF

[BL15] Guillaume Bouffard and Jean-Louis Lanet. The ultimate control flow transfer in a Java based smart card. Computers & Security, 50 (??):33–46, May 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016740481500005X.

Black:2018:NPJ

[Bla18] N. Black. Nicolai Parlog on Java 9 modules. IEEE Software, 35(3):101– 104, May/June 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bodden:2012:PEF

[BLH12] Eric Bodden, Patrick Lam, and Laurie Hendren. Partially evaluating finite-state runtime monitors ahead of time. ACM Transactions on Programming Languages and Systems, 34(2):7:1– 7:??, June 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barr:2014:TAT

[BM14] Earl T. Barr and Mark Marron. Tardis: affordable time-travel debugging in managed runtimes. ACM SIGPLAN Notices, 49(10): 67–82, October 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouraqadi:2018:TDD

[BM18] Noury Bouraqadi and Dave Mason. Test-driven development for generated portable Javascript apps. Science of Computer Programming, 161(??):2–17, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642318300595.

Bell:2015:VFB

[BMDK15] Jonathan Bell, Eric Melski, Mohan Dattatreya, and Gail E. Kaiser. Vroom: Faster build processes for Java. IEEE Software, 32(2): 97–104, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.org/
csdl/mags/so/2015/02/
mso2015020097-abs.html.

Brockschmidt:2012:ATP

[BMOG12] Marc Brockschmidt, Richard Musiol, Carsten Otto, and Jürgen Giesl. Automated termination proofs for Java programs with cyclic data. Lecture Notes in Computer Science, 7358: 105–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-31424-
7_13/.

Balland:2014:ESP

[BMR14] Emilie Balland, Pierre-Etienne Moreau, and Antoine Reilles. Effective strategic programming for Java developers. Software— Practice and Experience, 44 (2):129–162, February 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boldi:2018:BMC

[BMSV18] Paolo Boldi, Andrea Marino, Massimo Santini, and Sebastiano Vigna. BUbiNG: Massive crawling for the masses. ACM Transactions on the Web (TWEB), 12(2): 12:1–12:26, June 2018. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/citation.cfm?
doid=3176641.3160017.

Bliudze:2017:ECC

[BMSZ17] Simon Bliudze, Anastasia Mavridou, Radoslaw Szymanek, and Alina Zolotukhina. Exogenous coordination of concurrent software components with JavaBIP. Software—Practice and Experience, 47(11): 1801–1836, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:2016:HBS

[BNE16] Fraser Brown, Andres Nötzli, and Dawson Engler. How to build static checking systems using orders of magnitude less code. Operating Systems Review, 50(2):143–157, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borstler:2011:QEI

[BNP11] Jürgen Börstler, Marie Nordström, and James H. Paterson. On the quality of examples in introductory Java textbooks. ACM Transactions on Computing Education, 11(1):3:1–3:??, February 2011. CODEN ???? ISSN 1946-6226.

Baxter:2018:PAS

[BNP^{+}18] Samuel Baxter, Rachit Nigam, Joe Gibbs Politz, Shriram Krishnamurthi, and Arjun Guha. Putting in all the stops: execution control for JavaScript. ACM SIGPLAN Notices, 53(4): 30–45, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burnim:2012:SCS

[BNS12] Jacob Burnim, George Necula, and Koushik Sen. Specifying and checking semantic atomicity for multithreaded programs. ACM SIGPLAN Notices, 47(4): 79–90, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bellia:2011:PJS

[BO11] Marco Bellia and M. Eugenia Occhiuto. Properties of Java simple closures. Fundamenta Informaticae, 109 (3):237–253, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2012:ERT

[BO12] Marco Bellia and M. Eugenia Occhiuto. The equivalence of reduction and translation semantics of Java simple closures. Fundamenta Informaticae, 119(3–4):249–264, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2013:JST

[BO13] Marco Bellia and M. Eugenia Occhiuto. Java SAM typed closures: A sound and complete type inference system for nominal types. Fundamenta Informaticae, 128 (1–2):17–33, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bruno:2017:NPG

[BOF17] Rodrigo Bruno, Luís Picciochi Oliveira, and Paulo Ferreira. NG2C: pretenuring garbage collection with dynamic generations for HotSpot big data applications. ACM SIGPLAN Notices, 52(9):2–13, September 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barabash:2010:TGC

[BP10] Katherine Barabash and Erez Petrank. Tracing garbage collection on highly parallel platforms. ACM SIGPLAN Notices, 45(8):1– 10, August 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bender:2019:FJC

[BP19] John Bender and Jens Palsberg. A formalization of Java’s concurrent access modes. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):142:1–142:28, October 2019. URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3360568.

Bluemke:2012:DTJ

[BR12] Ilona Bluemke and Artur Rembiszewski. Dataflow testing of Java programs with DFC. Lecture Notes in Computer Science, 7054: 215–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-28038-
                                        2_17/.

Bogdanas:2015:KJC

[BR15] Denis Bogdanas and Grigore Rosu. K-Java: a complete semantics of Java. ACM SIGPLAN Notices, 50 (1):445–456, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brandt:2014:DAS

[Bra14] Siegmund Brandt. Data analysis: statistical and computational methods for scientists and engineers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2014. ISBN 3-319-03762-5 (ebook). ???? pp. LCCN QA273; QA273.

Bhattacharya:2012:DLI

[BRGG12] Suparna Bhattacharya, Karthick Rajamani, K. Gopinath, and Manish Gupta. Does lean imply green?: a study of the power performance implications of Java runtime bloat. ACM SIGMETRICS Perform. Eval. Rev., 40(1): 259–270, June 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brown:2012:BRF

[Bro12]  Neil Brown.  Book re-

view: Functional Programming for Java Developers — Tools for Better Concurrency, Abstraction, and Agility, By Dean Wampler, O’Reilly Media, July 2011, ISBN-13: 978-1-4493-1103-2, 90 pp. Journal of Functional Programming, 22(6):853–854, November 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://

www.cambridge.org/core/
product/02755329E5B068C2D403EE6722FCBD06.

Bosboom:2014:SCC

[BRWA14] Jeffrey Bosboom, Sumanaruban Rajadurai, Weng-Fai Wong, and Saman Amarasinghe. StreamJIT: a commensal compiler for high-performance stream programming. ACM SIG-PLAN Notices, 49(10):177– 195, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bedla:2012:SSJ

[BS12] Mariusz Bedla and Krzysztof Sapiecha. Scalable store of Java objects using range partitioning. Lecture Notes in Computer Science, 7054: 84–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-28038-
                                        2_7/.

Balatsouras:2013:CHC

[BS13] George Balatsouras and Yannis Smaragdakis. Class hierarchy complementation: soundly completing a partial type graph. ACM SIG-PLAN Notices, 48(10):515– 532, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Bouktif:2014:PSO

[BSA14] Salah Bouktif, Houari Sahraoui, and Faheem Ahmed. Predicting stability of open-source software systems using combination of Bayesian classifiers. ACM Transactions on Management Information Systems (TMIS), 5(1): 3:1–3:??, April 2014. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bastani:2018:ALP

[BSAL18] Osbert Bastani, Rahul Sharma, Alex Aiken, and Percy Liang. Active learning of points-to specifications. ACM SIGPLAN Notices, 53(4):678–692, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonetta:2016:GSM

[BSMB16] Daniele Bonetta, Luca Salucci, Stefan Marr, and Walter Binder. GEMs: shared-memory parallel programming for Node.js. ACM SIGPLAN Notices, 51 (10):531–547, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brachthauser:2018:EHM

[BSO18] Jonathan Immanuel Brachthäuser, Philipp Schuster, and Klaus Ostermann. Effect handlers for the masses. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 111:1–111:27, October 2018.

                                        URL  https://dl.acm.
                                        org/doi/abs/10.1145/3276481.

Brockschmidt:2012:ADN

[BSOG12] Marc Brockschmidt, Thomas Ströder, Carsten Otto, and Jürgen Giesl. Automated detection of non-

                                        termination and NullPointerExceptions

for Java bytecode. Lecture Notes in Computer Science, 7421:123–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.
                                        com/chapter/10.1007/978-
                                        3-642-31762-0_9/.

Bodden:2013:SLS

[BTR^{+}13] Eric Bodden, Társis Tolêdo, Márcio Ribeiro, Claus Brabrand, Paulo Borba, and Mira Mezini. SPL LIFT: statically analyzing software product lines in minutes instead of years. ACM SIGPLAN Notices, 48(6):355–364, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bultan:2018:SCA

[Bul18] Tevfik Bultan. Side-channel analysis via symbolic execution and model counting. ACM SIGSOFT Software Engineering Notes, 43(4): 55, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Basanta-Val:2010:SSS

[BVEAGVA10] Pablo Basanta-Val, Iria Estevez-Ayres, Marisol Garcia-Valls, and Luis Almeida. A synchronous scheduling service for distributed realtime Java. IEEE Transactions on Parallel and Distributed Systems, 21(4):506– 519, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Basanta-Val:2014:RMP

[BVGV14a] Pablo Basanta-Val and Marisol García-Valls. Resource management policies for real-time Java remote invocations. Journal of Parallel and Distributed Computing, 74(1):1930–1944, January 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0743731513001378.

Basanta-Val:2014:SDG

[BVGV14b] Pablo Basanta-Val and Marisol García-Valls. A simple distributed garbage collector for distributed real-time Java. The Journal of Supercomputing, 70(3):1588–1616, December 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s11227-014-1259-
                                        x.

Basanta-Val:2010:NHR

[BVGVEA10] Pablo Basanta-Val, Marisol García-Valls, and Iria Estévez-Ayres. No-Heap Remote Objects for distributed realtime Java. ACM Transactions on Embedded Computing Systems, 10(1):7:1– 7:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Basanta-Val:2011:ECM

[BVGVEA11a] P. Basanta-Val, M. García-Valls, and I. Estévez-Ayres. Extending the concurrency model of the real-time specification for Java. Concurrency and Computation: Practice and Experience, 23(14):1623–1645, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Basanta-Val:2011:NFI

[BVGVEA11b] P. Basanta-Val, M. Garcia-Valls, and I. Estevez-Ayres. Non-functional information transmission patterns for distributed real-time Java. Software—Practice and Experience, 41(12):1409–1435, ???? 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basanta-Val:2013:JRA

[BVGVEA13] P. Basanta-Val, M. García-Valls, and I. Estévez-Ayres. Enhancing OSGi with realtime Java support. Software —Practice and Experience, 43(1):33–65, January 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basanta-Val:2011:FTM

[BVGVEAFG11] Pablo Basanta-Val, Marisol García-Valls, Iria Estévez-Ayres, and Jorge Fernández-González. Fine tuning of the multiplexing facilities of Java’s Remote Method Invocation. Concurrency and Computation: Practice and Experience, 23(11):1236– 1260, August 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bourdykine:2012:LAM

[BW12] Pavel Bourdykine and Stephen M. Watt. Lightweight abstraction for mathematical computation in Java. Lecture Notes in Computer Science, 7442: 47–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-32973-
9_5/.

Briggs:2017:COI

[BZD17] Kim T. Briggs, Baoguo Zhou, and Gerhard W. Dueck. Cold object identification in the Java Virtual Machine. Software—Practice and Experience, 47(1): 79–95, January 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carlisle:2011:WCB

[Car11] Martin C. Carlisle. Why I came back to Ada. ACM SIGADA Ada Letters, 31 (3):37–38, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cao:2012:YYP

[CBGM12] Ting Cao, Stephen M. Blackburn, Tiejun Gao, and Kathryn S. McKinley. The yin and yang of power and performance for asymmetric hardware and managed software. ACM SIGARCH Computer Architecture News, 40(3):225– 236, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Chevalier-Boisvert:2012:BSH

[CBLFD12] Maxime Chevalier-Boisvert, Erick Lavoie, Marc Feeley, and Bruno Dufour. Bootstrapping a self-hosted research virtual machine for JavaScript: an experience report. ACM SIGPLAN Notices, 47(2):61–72, February 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaikalis:2015:FJS

[CC15] T. Chaikalis and A. Chatzigeorgiou. Forecasting Java software evolution trends employing network models. IEEE Transactions on Software Engineering, 41(6): 582–602, June 2015. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
6985636.

Cosentino:2012:MDR

[CCA^{+}12] Valerio Cosentino, Jordi Cabot, Patrick Albert, Philippe Bauquel, and Jacques Perronnet. A model driven reverse engineering framework for extracting business rules out of a Java application. Lecture Notes in Computer Science, 7438:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-32689-
9_3/.

Ceccato:2015:LSE

[CCFB15] Mariano Ceccato, Andrea Capiluppi, Paolo Falcarin, and Cornelia Boldyreff. A large study on the effect of code obfuscation on the quality of Java code. Empirical Software Engineering, 20(6):1486– 1524, December 2015. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s10664-014-9321-
0.

Chen:2011:MJP

[CCH11] Kuo-Yi Chen, J. Morris Chang, and Ting-Wei Hou. Multithreading in Java: Performance and scalability on multicore systems. IEEE Transactions on Computers, 60(11): 1521–1534, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=5661769.

Christophe:2018:ODA

[CDBD18] Laurent Christophe, Coen De Roover, Elisa Gonzalez Boix, and Wolfgang De Meuter. Orchestrating dynamic analyses of distributed processes for full-stack JavaScript programs. ACM SIG-PLAN Notices, 53(9):107– 118, November 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3393934.3278135.

Chisnall:2017:CJS

[CDG^{+}17] David Chisnall, Brooks Davis, Khilan Gudka, David Brazdil, Alexandre Joannou, Jonathan Woodruff, A. Theodore Markettos, J. Edward Maste, Robert Norton, Stacey Son, Michael Roe, Simon W. Moore, Peter G. Neumann, Ben Laurie, and Robert N. M. Watson. CHERI JNI: Sinking the Java security model into the C. ACM SIGARCH Computer Architecture News, 45(1):569– 583, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Coppolino:2019:CAE

[CDMR19] Luigi Coppolino, Salvatore D’Antonio, Giovanni Mazzeo, and Luigi Romano. A comparative analysis of emerging approaches for securing Java software with Intel SGX. Future Generation Computer Systems, 97(??):620–633, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315942.

Ceccato:2010:MLD

[CDTM10] Mariano Ceccato, Thomas Roy Dean, Paolo Tonella, and Davide Marchignoli. Migrating legacy data structures based on variable overlay to Java. Journal of Software Maintenance and Evolution: Research and Practice, 22(3):211–237, April 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Cecco:2011:SJG

[Cec11]  Raffaele Cecco.  Super-

charged JavaScript Graphics. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-9363-2. 400 (est.) pp. LCCN ????

Carter:2013:SSA

[CFH^{+}13] Kyle Carter, Adam Foltzer, Joe Hendrix, Brian Huffman, and Aaron Tomb. SAW: the software analysis workbench. ACM SIGADA Ada Letters, 33(3):15–18, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chandra:2016:TIS

[CGJ^{+}16] Satish Chandra, Colin S. Gordon, Jean-Baptiste Jeannin, Cole Schlesinger, Manu Sridharan, Frank Tip, and Youngil Choi. Type inference for static compilation of JavaScript. ACM SIG-PLAN Notices, 51(10):410– 429, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chamberlain:2017:PLR

[CH17] Scott Chamberlain and Jeffrey W. Hollister. lawn: An R client for the Turf Javascript Library for Geospatial Analysis. Journal of Open Source Software, 2 (11):194:1, March 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00194.

Chadha:2018:JAS

[Cha18] Gaurav Chadha. JSCore:

architectural support for accelerating JavaScript execution (short WIP paper). ACM SIGPLAN Notices, 53(6):104–108, June 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chugh:2012:DTJ

[CHJ12] Ravi Chugh, David Herman, and Ranjit Jhala. Dependent types for JavaScript. ACM SIGPLAN Notices, 47 (10):587–606, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carro:2013:MDA

[CHM13] Manuel Carro, Ángel Herranz, and Julio Mariño. A model-driven approach to teaching concurrency. ACM Transactions on Computing Education, 13(1):5:1– 5:??, January 2013. CO-DEN ???? ISSN 1946-6226. Chapman:2016:HSH

[CHM16] Keith Chapman, Antony L. Hosking, and J. Eliot B. Moss. Hybrid STM/HTM for nested transactions on OpenJDK. ACM SIG-PLAN Notices, 51(10):660– 676, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cogumbreiro:2015:DDV

[CHMY15] Tiago Cogumbreiro, Raymond Hu, Francisco Martins, and Nobuko Yoshida. Dynamic deadlock verification for general barrier synchronisation. ACM SIG-PLAN Notices, 50(8):150– 160, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cogumbreiro:2019:DDV

[CHMY19] Tiago Cogumbreiro, Raymond Hu, Francisco Martins, and Nobuko Yoshida. Dynamic deadlock verification for general barrier synchronisation. ACM Transactions on Programming Languages and Systems, 41 (1):1:1–1:??, March 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3229060.

Chong:2014:CCT

[Cho14] Stephen Chong. Checking correctness of TypeScript interfaces for JavaScript libraries. ACM SIGPLAN Notices, 49(10):1–16, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campbell:2013:ICC

[CIAD13] Bill Campbell, Swami Iyer, and Bahar Akbal-Delibas. Introduction to compiler construction in a Java world. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4398-6088-2 (hardcover). ???? pp. LCCN QA76.73.J38 C363 2013.

Chen:2017:CLP

[CJ17] Boyuan Chen and Zhen Ming (Jack) Jiang. Characterizing logging practices in Java-based open source software projects — a replication study in Apache Software Foundation. Empirical Software Engineering, 22 (1):330–374, February 2017. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s10664-016-9429-
5.

Chen:2019:ESL

[CJ19] Boyuan Chen and Zhen Ming (Jack) Jiang. Extracting and studying the logging–code-issue-introducing changes in Java-based largescale open source software systems. Empirical Software Engineering, 24(4):2285– 2322, August 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10664-019-09690-
0.

Cordeiro:2018:BJV

[CKS18] Lucas C. Cordeiro, Daniel Kroening, and Peter Schrammel. Benchmarking of Java verification tools at the Software Verification Competition (SV-COMP). ACM SIGSOFT Software Engineering Notes, 43(4): 56, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Canino:2017:PAE

[CL17] Anthony Canino and Yu David Liu. Proactive and adaptive energy-aware programming with mixed typechecking. ACM SIGPLAN Notices, 52(6):217–232, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clerc:2016:OJJ

[Cle16] Xavier Clerc. OCaml-Java:

The Java Virtual Machine as the target of an OCaml compiler. Journal of Functional Programming, 26:e7, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://

www.cambridge.org/core/
journals/journal-of-functional-
programming/article/ocamljava-
the-java-virtual-machine-
as-the-target-of-an-ocaml-
compiler/EDA83983550B6025B3705E2F8D97EB81.

Costa:2010:RMN

[CMM^{+}10] Gabriele Costa, Fabio Martinelli, Paolo Mori, Christian Schaefer, and Thomas Walter. Runtime monitoring for next generation Java ME platform. Computers & Security, 29(1):74– 87, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167404809000790.

Castro:2017:JLC

[CMM17] Sergio Castro, Kim Mens, and Paulo Moura. JPC: a library for categorising and applying interlanguage conversions between Java and Prolog. Science of Computer Programming, 134(??):75–99, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167642315004049.

Chang:2012:IOT

[CMS^{+}12] Mason Chang, Bernd Mathiske, Edwin Smith, Avik Chaudhuri, Andreas Gal, Michael Bebenita, Christian Wimmer, and Michael Franz. The impact of optional type information on JIT compilation of dynamically typed languages. ACM SIGPLAN Notices, 47(2):13–24, February 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Celik:2019:DIA

[CNRG19] Ahmet Celik, Pengyu Nie, Christopher J. Rossbach, and Milos Gligoric. Design, implementation, and application of GPU-based Java bytecode interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 177:1–177:28, October 2019.

                                        URL  https://dl.acm.
                                        org/doi/abs/10.1145/3360603.

Choi:2013:GGT

[CNS13] Wontae Choi, George Necula, and Koushik Sen. Guided GUI testing of Android apps with minimal restart and approximate learning. ACM SIG-PLAN Notices, 48(10):623– 640, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Clifford:2014:AFB

[CPST14] Daniel Clifford, Hannes Payer, Michael Starzinger, and Ben L. Titzer. Allocation folding based on dominance. ACM SIGPLAN Notices, 49(11):15–24, November 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clifford:2015:MMD

[CPST15] Daniel Clifford, Hannes Payer, Michael Stanton, and Ben L. Titzer. Memento mori: dynamic allocation-site-based optimizations. ACM SIGPLAN Notices, 50(11):105–117, November 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatterjee:2015:QIA

[CPV15] Krishnendu Chatterjee, Andreas Pavlogiannis, and Yaron Velner. Quantitative interprocedural analysis. ACM SIGPLAN Notices, 50(1):539–551, January 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Curley:2010:RDT

[CRAJ10] Edward Curley, Binoy Ravindran, Jonathan Anderson, and E. Douglas Jensen. Recovering from distributable thread failures in distributed realtime Java. ACM Transactions on Embedded Computing Systems, 10(1):8:1– 8:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cote:2012:JPS

[CRAT^{+}12] Marco Cote, German Riano, Raha Akhavan-Tabatabaei, Juan Fernando Perez, Andres Sarmiento, and Julio Goez. jMarkov package: a stochastic modeling tool. ACM SIGMETRICS Perform. Eval. Rev., 39(4): 48, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chalin:2010:TIG

[CRJ^{+}10] Patrice Chalin, Robby, Perry R. James, Jooyong Lee, and George Karabotsos. Towards an industrial grade IVE for Java and next generation research platform for JML. International Journal on Software Tools for Technology Transfer (STTT), 12(6): 429–446, November 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1433-2779&volume=
12&issue=6&spage=429.

Chambers:2010:FEE

[CRP^{+}10] Craig Chambers, Ashish Raniwala, Frances Perry, Stephen Adams, Robert R. Henry, Robert Bradshaw, and Nathan Weizenbaum. FlumeJava: easy, efficient data-parallel pipelines. ACM SIGPLAN Notices, 45(6):363–375, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ceccarello:2012:TGC

[CS12] Matteo Ceccarello and Nastaran Shafiei. Tools to generate and check consistency of model classes for Java PathFinder. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cordoba-Sanchez:2016:ADS

[CSdL16] Irene Córdoba-Sánchez and Juan de Lara. Ann: a domain-specific language for the effective design and validation of Java annotations. Computer Languages, Systems and Structures, ?? (??):164–190, ???? 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1477842416300318.

Chavez:2016:ACC

[CSF^{+}16] H. M. Chavez, W. Shen, R. B. France, B. A. Mechling, and G. Li. An approach to checking consistency between UML class model and its Java implementation. IEEE Transactions on Software Engineering, 42(4):322–344, April 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7294689.

Choi:2017:SAS

[CSGT17] Jiho Choi, Thomas Shull, Maria J. Garzaran, and Josep Torrellas. Short-Cut: Architectural support for fast object access in scripting languages. ACM SIGARCH Computer Architecture News, 45(2):494– 506, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chawdhary:2017:PES

[CSK17] Aziem Chawdhary, Ranjeet Singh, and Andy King. Partial evaluation of string obfuscations for Java malware detection. Formal Aspects of Computing, 29 (1):33–55, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s00165-016-0357-
3;   http://link.springer.
com/article/10.1007/s00165-
016-0357-3.

Chanda:2012:TBS

[CSKB12] Jayeeta Chanda, Sabnam Sengupta, Ananya Kanjilal, and Swapan Bhattacharya. Traceability between service component and class: a model based approach. ACM SIGSOFT Software Engineering Notes, 37(6):1– 5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2016:CDD

[CSS^{+}16] Yuting Chen, Ting Su, Chengnian Sun, Zhendong Su, and Jianjun Zhao. Coverage-directed differential testing of JVM implementations. ACM SIG-PLAN Notices, 51(6):85– 99, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:2015:JFE

[CSV15] Callum Cameron, Jeremy Singer, and David Vengerov. The judgment of FORSETI: economic utility for dynamic heap sizing of multiple runtimes. ACM SIG-PLAN Notices, 50(11):143– 156, November 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casale:2017:PEJ

[CSZ17] Giuliano Casale, Giuseppe Serazzi, and Lulai Zhu. Performance evaluation with Java modelling tools: a hands-on introduction. ACM SIGMETRICS Perform. Eval. Rev., 45(3):246–247, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cazzola:2014:JBR

[CV14] Walter Cazzola and Edoardo Vacchi. @Java: Bringing a richer annotation model to Java. Computer Languages, Systems and Structures, 40 (1):2–18, April 2014. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1477842414000037.

Chaudhuri:2017:FPT

[CVG^{+}17] Avik Chaudhuri, Panagiotis Vekris, Sam Goldman, Marshall Roch, and Gabriel Levi. Fast and precise type checking for JavaScript. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):48:1–48:??, October 2017. CODEN ???? ISSN 2475-1421.

Chan:2017:DSL

[CWGA17] Y. Chan, A. Wellings, I. Gray, and N. Audsley. A distributed stream library for Java 8. IEEE Transactions on Big Data, 3(3):262– 275, September 2017. ISSN 2332-7790.

Cavalcanti:2013:SCJ

[CWW13] Ana Cavalcanti, Andy Wellings, and Jim Woodcock. The Safety-Critical Java memory model formalised. Formal Aspects of Computing, 25(1):37–57, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s00165-012-0253-
4.

Chen:2019:SVR

[CYWD19] Zhenbang Chen, Hengbiao Yu, Ji Wang, and Wei Dong. Symbolic verification of regular properties for Java programs. ACM SIGSOFT Software Engineering Notes, 44(4):17, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://

dl.acm.org/doi/10.1145/
3364452.33644561.

Caserta:2014:JTJ

[CZ14] Pierre Caserta and Olivier Zendra. JBInsTrace: a tracer of Java and JRE classes at basic-block granularity by dynamically instrumenting bytecode. Science of Computer Programming, 79(??):116–125, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642312000299.

Diaz:2013:LEU

[DAA13] Oscar Díaz, Cristóbal Arellano, and Maider Azanza. A language for end-user Web augmentation: Caring for producers and consumers alike. ACM Transactions on the Web (TWEB), 7(2): 9:1–9:??, May 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dannen:2017:IES

[Dan17] Chris Dannen. Introducing Ethereum and Solidity. Apress, Brooklyn, NY, 2017. ISBN 1-4842-2535-X. xxi + 185 pp. URL

http://link.springer.
com/book/10.1007/978-1-
4842-2535-6.

Dan:2018:LJE

[Dan18] Toomey Dan. Learning Jupyter 5: explore interactive computing using Python, Java, JavaScript, R, Julia, and JupyterLab. Packt Publishing, Birmingham, 2018. ISBN 1-78913-740-3, 1-78913-744-

6. 282 pp. LCCN Q183.9; QA76.9.I52 .T666 2018.

Resenes:2019:JJM

[dARPH^{+}19] Jonas de Abreu Resenes, Willingthon Pavan, Carlos Amaral Hölbig, José Maurício Cunha Fernandes, Vakhtang Shelia, Cheryl Porter, and Gerrit Hoogenboom. jDSSAT: a JavaScript module for DSSAT-CSM integration. SoftwareX, 10 (??):Article 100271, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S235271101930158X.

daCosta:2012:JSL

[dCMMN12] Umberto Souza da Costa, Anamaria Martins Moreira, Martin A. Musicante, and Plácido A. Souza Neto. JCML: a specification language for the runtime verification of Java Card programs. Science of Computer Programming, 77(4):

533–550, April 1, 2012. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167642310000596.

Dhawan:2012:EJT

[DcSG12] Mohan Dhawan, Chung chieh Shan, and Vinod Ganapathy. Enhancing JavaScript with transactions. Lecture Notes in Computer Science, 7313: 383–408, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-31057-
                                        7_18/.

DElia:2013:BLP

[DD13] Daniele Cono D’Elia and Camil Demetrescu. Ball– Larus path profiling across multiple loop iterations. ACM SIGPLAN Notices, 48 (10):373–390, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

DeBeukelaer:2017:ECP

[DDDF17] Herman De Beukelaer, Guy F. Davenport, Geert De Meyer, and Veerle Fack. Extended conference paper: JAMES: an objectoriented Java framework for discrete optimization using local search metaheuristics. Software—Practice and Experience, 47(6):921– 938, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dietl:2011:SOT

[DDM11] Werner Dietl, Sophia Drossopoulou, and Peter Müller. Separating ownership topology and encapsulation with generic universe types. ACM Transactions on Programming Languages and Systems, 33 (6):20:1–20:??, December 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Deitcher:2010:JEJ

[Dei10] Avi Deitcher. JSormdb — an embedded JavaScript database. Linux Journal, 2010(192):4:1–4:??, April 2010. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Deitcher:2011:SPJ

[Dei11] Avi Deitcher. Simplicity and performance: JavaScript on the server. Linux Journal, 2011(204):3:1–3:??, April 2011. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

DelRa:2013:BRJ

[Del13] William Del Ra III. Book review: Java application architecture: modularity patterns with examples using OSGi by Kirk Knoernschild. ACM SIGSOFT Software Engineering Notes, 38(1): 55, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dennis:2018:MFI

[Den18] Louise A. Dennis. The MCAPL framework including the Agent Infrastructure Layer and Agent Java Pathfinder. Journal of Open Source Software, 3(24): 617:1–617:2, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00617.

Disney:2015:SYJ

[DFHF15] Tim Disney, Nathan Faubion, David Herman, and Cormac Flanagan. Sweeten your JavaScript: hygienic macros for ES5. ACM SIG-PLAN Notices, 50(2):35–44, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dey:2013:STA

[DFR13] Akon Dey, Alan Fekete, and Uwe Röhm. Scalable transactions across heterogeneous NoSQL key-value data stores. Proceedings of the VLDB Endowment, 6(12):1434–1439, August 2013. CODEN ???? ISSN 2150-8097.

deGouw:2015:OJU

[dGRdB^{+}15] Stijn de Gouw, Jurriaan Rot, Frank S. de Boer, Richard Bubel, and Reiner

Hähnle. OpenJDK’s Java.utils.Collection.sort()

is broken: The good, the bad and the worst case. In Kroening and Păsăreanu [KP15], pages 273–289. ISBN 3-319-21689-9. URL http://

envisage-project.eu/wp-
content/uploads/2015/
02/sorting.pdf;   http:
//link.springer.com/chapter/
10.1007/978-3-319-21690-
4_16.

DHondt:2012:ISS

[D’H12] Theo D’Hondt. An interpreter for server-side HOP. ACM SIGPLAN Notices, 47(2):1–12, February 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolby:2012:DCA

[DHM^{+}12] Julian Dolby, Christian Hammer, Daniel Marino, Frank Tip, Mandana Vaziri, and Jan Vitek. A datacentric approach to synchronization. ACM Transactions on Programming Languages and Systems, 34 (1):4:1–4:48, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dietrich:2015:GSE

[DHS15] Jens Dietrich, Nicholas Hollingum, and Bernhard Scholz. Giga-scale exhaustive points-to analysis for Java in under a minute. ACM SIGPLAN Notices, 50 (10):535–551, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DiPierro:2018:RJ

[DiP18a] Massimo DiPierro. The rise of JavaScript. Computing in Science and Engineering, 20(1):9–10, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2018:TVG

[DiP18b] Massimo DiPierro. Toy vision-guided 3D robotic arm in JavaScript. Computing in Science and Engineering, 20(1):43–49, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        document/8291773/.

Dietrich:2016:WJD

[DJB16] Jens Dietrich, Kamil Jezek, and Premek Brada. What Java developers know about compatibility, and why this matters. Empirical Software Engineering, 21(3):

1371–1396, June 2016. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s10664-015-9389-
1.

Dam:2010:PCI

[DJLP10] Mads Dam, Bart Jacobs, Andreas Lundblad, and Frank Piessens. Provably correct inline monitoring for multithreaded Java-like programs. Journal of Computer Security, 18(1):37–59, ???? 2010. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

deJong:2018:MJA

[dJM18] Jos de Jong and Eric Mansfield. Math.Js: An advanced mathematics library for JavaScript. Computing in Science and Engineering, 20(1):20–32, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291769/.

DeFrancesco:2010:UAI

[DLM10] Nicoletta De Francesco, Giuseppe Lettieri, and Luca Martini. Using abstract interpretation to add type checking for interfaces in Java bytecode verification. Theoretical Computer Science, 411(22–24):

2174–2201, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeNicola:2014:FAA

[DLPT14] Rocco De Nicola, Michele Loreti, Rosario Pugliese, and Francesco Tiezzi. A formal approach to autonomic systems programming: The SCEL language. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(2): 7:1–7:??, July 2014. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dissegna:2014:TCA

[DLR14] Stefano Dissegna, Francesco Logozzo, and Francesco Ranzato. Tracing compilation by abstract interpretation. ACM SIGPLAN Notices, 49(1):47–59, January 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). POPL ’14 conference proceedings.

Dissegna:2016:AIB

[DLR16] Stefano Dissegna, Francesco Logozzo, and Francesco Ranzato. An abstract interpretation-based model of tracing just-in-time compilation. ACM Transactions on Programming Languages and Systems, 38 (2):7:1–7:??, January 2016.

CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Demange:2013:PBB

[DLZ^{+}13] Delphine Demange, Vincent Laporte, Lei Zhao, Suresh Jagannathan, David Pichardie, and Jan Vitek. Plan B: a buffered memory model for Java. ACM SIG-PLAN Notices, 48(1):329– 342, January 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deMol:2012:GTJ

[dMRH12] Maarten de Mol, Arend Rensink, and James J. Hunt. Graph transforming Java data. Lecture Notes in Computer Science, 7212: 209–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-28872-
2_15/.

Duarte:2011:ICS

[DMS11] Rafael Duarte, Alexandre Mota, and Augusto Sampaio. Introducing concurrency in sequential Java via laws. Information Processing Letters, 111(3):129–134, January 1, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devietti:2012:RRC

[DNB^{+}12] Joseph Devietti, Jacob Nelson, Tom Bergan, Luis Ceze, and Dan Grossman. RCDC: a relaxed consistency deterministic computer. ACM SIGPLAN Notices, 47(4):67–78, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dietrich:2010:POD

[DR10] K. Dietrich and F. Röck. Performance optimizations for DAA signatures on Java enabled platforms. J.UCS: Journal of Universal Computer Science, 16 (4):519–??, ???? 2010. CODEN ???? ISSN 0948-6968. URL http://

                                        www.jucs.org/jucs_16_4/
                                        performance_optimizations_
                                        for_daa.

Dyer:2014:DVE

[DRN14] Robert Dyer, Hridesh Rajan, and Tien N. Nguyen. Declarative visitors to ease fine-grained source code mining with full history on billions of AST nodes. ACM SIGPLAN Notices, 49(3): 23–32, March 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Doeraene:2016:PIW

[DS16] Sébastien Doeraene and Tobias Schlatter. Parallel incremental whole-program optimizations for Scala.js. ACM SIGPLAN Notices, 51 (10):59–73, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bois:2013:BGV

[DSEE13] Kristof Du Bois, Jennifer B. Sartor, Stijn Eyerman, and Lieven Eeckhout. Bottle graphs: visualizing scalability bottlenecks in multi-threaded applications. ACM SIG-PLAN Notices, 48(10):355– 372, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

David:2014:CMC

[DTLM14] Florian David, Gael Thomas, Julia Lawall, and Gilles Muller. Continuously measuring critical section pressure with the free-lunch profiler. ACM SIGPLAN Notices, 49(10):291–307, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Daloze:2018:PDL

[DTM^{+}18] Benoit Daloze, Arie Tal, Stefan Marr, Hanspeter Mössenböck, and Erez Petrank. Parallelization of dynamic languages: synchronizing built-in collections. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):108:1–108:30, October 2018. URL

https://dl.acm.org/doi/
abs/10.1145/3276478.

Dias:2013:SIP

[DVL13] Ricardo J. Dias, Tiago M. Vale, and João M. Lourenço. Special issue papers: Efficient support for in-place metadata in Java software transactional memory. Concurrency and Computation: Practice and Experience, 25(17):2394–2411, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DosSantos:2010:MPB

[DW10] Osmar Marchi Dos Santos and Andy Wellings. Measuring and policing blocking times in real-time systems. ACM Transactions on Embedded Computing Systems, 10(1):2:1–2:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Estevez-Ayres:2014:CSS

[EABVGV14] Iria Estévez-Ayres, Pablo Basanta-Val, and Marisol García-Valls. Composing and scheduling serviceoriented applications in time-triggered distributed real-time Java environments. Concurrency and Computation: Practice and Experience, 26(1):152–193, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

elBoustani:2011:ITE

[eBH11] Nabil el Boustani and Jurriaan Hage. Improving type error messages for generic Java. Higher-Order and Symbolic Computation, 24 (1–2):3–39, June 2011. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1388-3690&volume=
24&issue=1&spage=3.

Emerick:2012:CP

[ECG12] Chas Emerick, Brian Carper, and Christophe Grand. Clojure programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2012. ISBN 1-4493-9470-1. xviii + 607 pp. LCCN QA76.73.C565 E538 2012. URL http:

//www.loc.gov/catdir/

                                        enhancements/fy1211/2012405367-
                                        b.html;   http://www.
                                        loc.gov/catdir/enhancements/
                                        fy1211/2012405367-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1211/2012405367-
                                        t.html.

Ebert:2015:ESE

[ECS15] Felipe Ebert, Fernando Castor, and Alexander Serebrenik. An exploratory study on exception handling bugs in Java programs. The Journal of systems and software, 106(??):82–101, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121215000862.

Efftinge:2013:XID

[EEK^{+}13] Sven Efftinge, Moritz Eysholdt, Jan Köhnlein, Sebastian Zarnekow, Robert von Massow, Wilhelm Hasselbring, and Michael Hanus. Xbase: implementing domain-specific languages for Java. ACM SIG-PLAN Notices, 48(3):112– 121, March 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Erdweg:2012:GLE

[EKR^{+}12] Sebastian Erdweg, Lennart C. L. Kats, Tillmann Rendel, Christian Kästner, Klaus Ostermann, and Eelco Visser. Growing a language environment with editor libraries. ACM SIG-PLAN Notices, 47(3):167– 176, March 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Egbring:2010:POS

[EKUR10] Marco Egbring, Gerd A.

Kullak-Ublick, and Stefan Russmann. Phynx: an open source software solution supporting data management and web-based patient-level data review for drug safety studies in the general practice research database and other health care databases. Pharmacoepidemiology and Drug Safety, 19(1):38–44, 2010. ISSN 1053-8569 (print), 1099-1557 (electronic).

Erdweg:2015:SOI

[ELW15] Sebastian Erdweg, Moritz Lichter, and Manuel Weiel. A sound and optimal incremental build system with dynamic dependencies. ACM SIGPLAN Notices, 50(10):89–106, October 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eslamimehr:2014:RDS

[EP14] Mahdi Eslamimehr and Jens Palsberg. Race directed scheduling of concurrent programs. ACM SIG-PLAN Notices, 49(8):301– 314, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elmas:2010:GRA

[EQT10] Tayfun Elmas, Shaz Qadeer, and Serdar Tasiran. Goldilocks: a race-aware Java runtime. Communications of the ACM, 53(11):85–92, November 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erdweg:2014:FEL

[ER14] Sebastian Erdweg and Felix Rieger. A framework for extensible languages. ACM SIGPLAN Notices, 49(3):3– 12, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eichelberger:2014:FRM

[ES14] Holger Eichelberger and Klaus Schmid. Flexible resource monitoring of Java programs. The Journal of systems and software, 93(??):163–186, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121214000533.

Esquembre:2011:TPL

[Esq11] Francisco Esquembre. There is parallel life for Java scientific programmers! Computing in Science and Engineering, 13(4):6–10, July/ August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Endrullis:2012:WEM

[ETR12] Stefan Endrullis, Andreas Thor, and Erhard Rahm. WETSUIT: an efficient mashup tool for searching and fusing web entities. Proceedings of the VLDB Endowment, 5(12): 1970–1973, August 2012. CODEN ???? ISSN 2150-8097.

Exposito:2015:LLJ

[ETR^{+}15] Roberto R. Expósito, Guillermo L.

Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Low-latency Java communication devices on RDMA-enabled networks. Concurrency and Computation: Practice and Experience, 27(17):4852–4879, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Exposito:2012:DSJ

[ETTD12] Roberto R. Expósito, Guillermo L.

Taboada, Juan Touriño, and Ramón Doallo. Design of scalable Java messagepassing communications over InfiniBand. The Journal of Supercomputing, 61 (1):141–165, July 2012. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
61&issue=1&spage=141.

Eugster:2013:SUP

[Eug13] Patrick Eugster. Safe uniform proxies for Java. Science of Computer Programming, 78(9):1490–1520, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642312002043.

Evans:2013:WGJ

[EV13] Benjamin J. Evans and Martijn Verburg. The well-grounded Java developer: vital techniques of Java 7 and polyglot programming. Manning Publications, Greenwich, CT, USA, 2013. ISBN 1-61729-006-8. xxx + 462 pp. LCCN QA76.73.J38 E93 2013. URL http:

//www.loc.gov/catdir/
enhancements/fy1304/2012288194-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy1304/2012288194-d.html.

Foreword by Heinz Kabutz.

Fetter:2017:TJT

[FBG17] Mirko Fetter, David Bimamisa, and Tom Gross. TUIOFX: a JavaFX toolkit for shared interactive surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):10:1–10:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/doi/
10.1145/3095812.

Foley-Bourgon:2017:EIC

[FBH17] Vincent Foley-Bourgon and Laurie Hendren. Efficiently implementing the copy semantics of MATLAB’s arrays in JavaScript. ACM SIGPLAN Notices, 52(2): 72–83, February 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fernandes:2011:LFS

[FC11] Sérgio Miguel Fernandes and João Cachopo. Lockfree and scalable multiversion software transactional memory. ACM SIG-PLAN Notices, 46(8):179– 188, August 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Feeley:2016:CML

[Fee16] Marc Feeley. Compiling for multi-language task migration. ACM SIGPLAN Notices, 51(2):63–77, February 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrara:2013:GSA

[Fer13] P. Ferrara. A generic static analyzer for multithreaded Java programs. Software —Practice and Experience, 43(6):663–684, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Flanagan:2010:AMD

[FF10] Cormac Flanagan and Stephen N. Freund. Adversarial memory for detecting destructive races. ACM SIGPLAN Notices, 45(6): 244–254, June 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrari:2017:JJF

[FFF17] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. JTabWb: a Java framework for implementing terminating sequent and tableau calculi. Fundamenta Informaticae, 150 (1):119–142, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Candel:2019:DMD

[FGB^{+}19] Carlos Javier Fernández Candel, Jesús García Molina, Francisco Javier Bermúdez Ruiz, Jose Ramón Hoyos Barceló, Diego Sevilla Ruiz, and Benito José Cuesta Viera. Developing a modeldriven reengineering approach for migrating PL/ SQL triggers to Java: a practical experience. The Journal of systems and software, 151(??):38–64, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121219300214.

Femminella:2012:EJC

[FGR12] Mauro Femminella, Francesco Giacinti, and Gianluca Reali. An extended Java call control for the Session Initiation Protocol. IEEE MultiMedia, 19(4):60–71, October/December 2012. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Fogus:2011:JC

[FH11] Michael Fogus and Chris Houser. The joy of Clojure. Manning Publications, Greenwich, CT, USA, 2011. ISBN 1-935182-64-1 (paperback). xxxi + 328 pp. LCCN QA76.62 .F64 2011.

Fischer:2016:EIE

[FH16] Lars Fischer and Stefan Hanenberg. An empirical investigation of the effects of type systems and code completion on API usability using TypeScript and JavaScript in MS Visual Studio. ACM SIG-PLAN Notices, 51(2):154– 167, February 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forth:2012:RAA

[FHP^{+}12] Shaun Forth, Paul Hovland, Eric Phipps, Jean Utke, and Andrea Walther, editors. Recent Advances in Algorithmic Differentiation, volume 87 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ????

URL http://link.springer.
com/book/10.1007/978-
3-642-30023-3;   http:
//www.springerlink.com/
content/978-3-642-30023-

3. Proceedings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Fontaine:2012:VCF

[FHSR12] Arnaud Fontaine, Samuel Hym, and Isabelle Simplot-Ryl. Verifiable control flow policies for Java bytecode. Lecture Notes in Computer Science, 7140: 115–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-29420-
4_8/.

Freudenberg:2015:SMP

[FIF^{+}15] Bert Freudenberg, Dan H. H. Ingalls, Tim Felgentreff, Tobias Pape, and Robert Hirschfeld. SqueakJS: a modern and practical smalltalk that runs in any browser. ACM SIGPLAN Notices, 50(2): 57–66, February 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2013:PES

[FLL^{+}13] Cormac Flanagan, K. Rustan M. Leino, Mark Lillibridge, Greg Nelson, James B. Saxe, and Raymie Stata. PLDI 2002: Extended static checking for Java. ACM SIGPLAN Notices, 48(4S):22–33, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fan:2018:VCJ

[FLZ^{+}18] Linyu Fan, Jianwei Liao, Junsen Zuo, Kebo Zhang, Chao Li, and Hailing Xiong. Version 4.0 of code Java for 3D simulation of the CCA model. Computer Physics Communications, 228(??): 290–292, July 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518300869.

Feldthaus:2013:SAR

[FM13] Asger Feldthaus and Anders Møller. Semi-automatic rename refactoring for JavaScript. ACM SIG-PLAN Notices, 48(10):323– 338, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Felgentreff:2015:CBC

[FMBH15] Tim Felgentreff, Todd Millstein, Alan Borning, and Robert Hirschfeld. Checks and balances: constraint solving without surprises in object-constraint programming languages. ACM SIG-PLAN Notices, 50(10):767– 782, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldthaus:2011:TSR

[FMM^{+}11] Asger Feldthaus, Todd Millstein, Anders Møller, Max Schäfer, and Frank Tip. Tool-supported refactoring for JavaScript. ACM SIG-PLAN Notices, 46(10):119– 138, October 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Frantzeskou:2011:SUD

[FMS^{+}11] Georgia Frantzeskou, Stephen G.

MacDonell, Efstathios Stamatatos, Stelios Georgiou, and Stefanos Gritzalis. The significance of user-defined identifiers in Java source code authorship identification. International Journal of Computer Systems Science and Engineering, 26 (2):??, March 2011. CO-DEN CSSEEI. ISSN 0267-6192.

Fu:2014:FDC

[FOPZ14] Yupeng Fu, Kian Win Ong, Yannis Papakonstantinou, and Erick Zamora. Forward: data-centric ULS using declarative templates that efficiently wrap thirdparty JavaScript components. Proceedings of the VLDB Endowment, 7(13): 1649–1652, August 2014. CODEN ???? ISSN 2150-8097.

Fox:2017:ESI

[Fox17a] Geoffrey Fox. Editorial: Special issue on 12th international workshop on Java technologies for realtime and embedded systems (JTRES2014). Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fox:2017:EJT

[Fox17b] Geoffrey Fox. Editorials: Java Technologies for Real-Time and Embedded Systems (JTRES2013). Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fernandes:2017:AUM

[FRC^{+}17] Leonardo Fernandes, Márcio Ribeiro, Luiz Carvalho, Rohit Gheyi, Melina Mongiovi, André Santos, Ana Cavalcanti, Fabiano Ferrari, and José Carlos Maldonado. Avoiding useless mutants. ACM SIGPLAN Notices, 52(12):187–198, December 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fdez-Riverola:2012:JAF

[FRGPLF^{+}12] F. Fdez-Riverola, D. Glez-Peña, H. López-Fernández, M. Reboiro-Jato, and J. R. Méndez. A Java application framework for scientific software development. Software—Practice and Experience, 42(8):1015–1036, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fan:2015:UCC

[FRM^{+}15] Hua Fan, Aditya Ramaraju, Marlon McKenzie, Wojciech Golab, and Bernard Wong. Understanding the causes of consistency anomalies in Apache Cassandra. Proceedings of the VLDB Endowment, 8(7):810–813, February 2015. CODEN ???? ISSN 2150-8097.

Fournet:2013:FAC

[FSC^{+}13] Cedric Fournet, Nikhil Swamy, Juan Chen, Pierre-Evariste Dagand, Pierre-Yves Strub, and Benjamin Livshits. Fully abstract compilation to JavaScript. ACM SIGPLAN Notices, 48 (1):371–384, January 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Funes:2012:RMC

[FSK12] Diego Funes, Junaid Haroon Siddiqui, and Sarfraz Khurshid. Ranged model checking. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feng:2015:EQD

[FWDL15] Yu Feng, Xinyu Wang, Isil Dillig, and Calvin Lin. EXPLORER: query- and demand-driven exploration of interprocedural control flow properties. ACM SIG-PLAN Notices, 50(10):520– 534, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fritz:2017:TSA

[FZ17] Eric Fritz and Tian Zhao. Typing and semantics of asynchronous arrows in JavaScript. Science of Computer Programming, 141–142(??):1–39, July/ August 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642317300527.

Gherardi:2012:JVC

[GBC12] Luca Gherardi, Davide Brugali, and Daniele Comotti. A Java vs. C++ performance evaluation: a 3D modeling benchmark. Lecture Notes in Computer Science, 7628:161– 172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-34327-
8_17/.

Gerakios:2013:FIS

[GBS13] Prodromos Gerakios, Aggelos Biboudis, and Yannis Smaragdakis. Forsaking inheritance: supercharged delegation in DelphJ. ACM SIGPLAN Notices, 48(10): 233–252, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Gerakios:2014:RTP

[GBS14] Prodromos Gerakios, Aggelos Biboudis, and Yannis Smaragdakis. Reified type parameters using Java annotations. ACM SIG-PLAN Notices, 49(3):61– 64, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gama:2010:SAA

[GD10] Kiev Gama and Didier Donsez. A survey on approaches for addressing dependability attributes in the OSGi service platform. ACM SIG-SOFT Software Engineering Notes, 35(3):1–8, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

German:2012:MOS

[GD12] Daniel German and Massimiliano Di Penta. A method for open source license compliance of Java applications. IEEE Software, 29 (3):58–63, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gupta:2018:HDB

[GGC18] Shashank Gupta, B. B. Gupta, and Pooja Chaudhary. Hunting for DOMbased XSS vulnerabilities in mobile cloud-based online social network. Future Generation Computer Systems, 79 (part 1)(??): 319–336, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311068.

Gupta:2019:CSJ

[GGC19] Shashank Gupta, B. B. Gupta, and Pooja Chaudhary. A client-server JavaScript code rewritingbased framework to detect the XSS worms from online social network. Concurrency and Computation: Practice and Experience, 31(21):e4646:1– e4646:??, November 10, 2019. CODEN CCPEBO.

ISSN 1532-0626 (print), 1532-0634 (electronic).

Golan-Gueta:2014:ASL

[GGRSY14] Guy Golan-Gueta, G. Ramalingam, Mooly Sagiv, and Eran Yahav. Automatic semantic locking. ACM SIGPLAN Notices, 49(8): 385–386, August 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Golan-Gueta:2015:ASA

[GGRSY15] Guy Golan-Gueta, G. Ramalingam, Mooly Sagiv, and Eran Yahav. Automatic scalable atomicity via semantic locking. ACM SIG-PLAN Notices, 50(8):31– 41, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Golan-Gueta:2017:ASA

[GGRSY17] Guy Golan-Gueta, G. Ramalingam, Mooly Sagiv, and Eran Yahav. Automatic scalable atomicity via semantic locking. ACM Transactions on Parallel Computing (TOPC), 3 (4):21:1–21:??, March 2017. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Gligoric:2015:GCB

[GGZ^{+}15] Milos Gligoric, Alex Groce, Chaoqiang Zhang, Rohan Sharma, Mohammad Amin Alipour, and Darko Marinov. Guidelines for coverage-based comparisons of non-adequate test suites. ACM Transactions on Software Engineering and Methodology, 24(4): 22:1–22:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Gosling:2013:JLS

[GJS^{+}13] James Gosling, Bill Joy, Guy L. Steele Jr., Gilad Bracha, and Alex Buckley. The Java Language Specification. Addison-Wesley, Reading, MA, USA, Java SE 7 edition, 2013. ISBN 0-13-326022-4 (paperback). xxvii + 644 pp. LCCN QA76.73.J38 G68 2013.

Gosling:2014:JLS

[GJS^{+}14] James Gosling, Bill Joy, Guy L. Steele Jr., Gilad Bracha, and Alex Buckley. The Java Language Specification. Addison-Wes-ley, Addison-Wesley, Java SE 8 edition, 2014. ISBN 0-13-390069-X (paperback). xxii + 758 pp. LCCN QA76.73.J38 G68 2014.

Gvero:2015:SJE

[GK15] Tihomir Gvero and Viktor Kuncak. Synthesizing Java expressions from freeform queries. ACM SIG-PLAN Notices, 50(10):416– 432, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabaruk:2019:TJO

[GLGA19] J. Gabaruk, D. Logofatu, D. Großkreutz, and C. Andersson. On teaching Java and object oriented programming by using children board games. In 2019 IEEE Global Engineering Education Conference (EDUCON), pages 601–606. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2019.

Gejibo:2012:CIE

[GM12] Samson Gejibo and Federico Mancini. Challenges in implementing an end-to-end secure protocol for Java ME-based mobile data collection in low-budget settings. Lecture Notes in Computer Science, 7159: 38–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-28166-
2_5/.

Gonzalez:2013:HBP

[GMC^{+}13] Apolinar Gonzalez, Walter Mata, Alfons Crespo, Miguel Masmano, José Félix, and Alvaro Aburto. A hypervisor based platform to support real-time safety critical embedded Java applications. International Journal of Computer Systems Science and Engineering, 28(3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192. Gadyatskaya:2013:LTC

[GMNC13] Olga Gadyatskaya, Fabio Massacci, Quang-Huy Nguyen, and Boutheina Chetali. Load time code validation for mobile phone Java cards. Journal of Information Security and Applications (JISA), 18(2–3): 108–129, September 2013. CODEN ???? ISSN 2214-2126. URL http:/

/www.sciencedirect.com/
science/article/pii/S2214212613000318.

Gadyatskaya:2012:JCA

[GMPS12] Olga Gadyatskaya, Fabio Massacci, Federica Paci, and Sergey Stankevich. Java card architecture for autonomous yet secure evolution of Smart Cards applications. Lecture Notes in Computer Science, 7127: 187–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-27937-
9_13/.

Gardner:2012:TPL

[GMS12] Philippa Anne Gardner, Sergio Maffeis, and Gareth David Smith. Towards a program logic for JavaScript. ACM SIGPLAN Notices, 47(1): 31–44, January 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenman:2014:GFB

[GMT14] Ben Greenman, Fabian Muehlboeck, and Ross Tate. Getting F-bounded polymorphism into shape. ACM SIGPLAN Notices, 49(6): 89–99, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2016:LSA

[GN16] Kartik Gupta and V. Krishna Nandivada. Lexical state analyzer for JavaCC grammars. Software—Practice and Experience, 46(6): 751–765, June 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gong:2011:JSA

[Gon11] Li Gong. Java security architecture revisited. Communications of the ACM, 54 (11):48–52, November 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grossschadl:2012:EJI

[GPT12] Johann Großschädl, Dan Page, and Stefan Tillich. Efficient Java implementation of elliptic curve cryptography for J2ME-enabled mobile devices. Lecture Notes in Computer Science, 7322: 189–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-30955-
7_17/.

Gramoli:2015:MTY

[Gra15] Vincent Gramoli. More than you ever wanted to know about synchronization: synchrobench, measuring the impact of the synchronization on concurrent algorithms. ACM SIG-PLAN Notices, 50(8):1–10, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grech:2011:JGE

[GRF11] Neville Grech, Julian Rathke, and Bernd Fischer. JEqualityGen: generating equality and hashing methods. ACM SIGPLAN Notices, 46 (2):177–186, February 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grigore:2017:JGT

[Gri17] Radu Grigore. Java generics are Turing complete. ACM SIGPLAN Notices, 52(1): 73–85, January 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giacaman:2011:OOP

[GS11] Nasser Giacaman and Oliver Sinnen. Object-oriented parallelisation of Java desktop programs. IEEE Software, 28(1):32–38, January/ February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gil:2012:SFJ

[GS12] Joseph Gil and Yuval Shimron. Smaller footprint for Java collections. Lecture Notes in Computer Science, 7313:356– 382, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-31057-
7_17/.

Gill:2015:RMD

[GSD^{+}15] Andy Gill, Neil Sculthorpe, Justin Dawson, Aleksander Eskilson, Andrew Farmer, Mark Grebe, Jeffrey Rosenbluth, Ryan Scott, and James Stanton. The remote monad design pattern. ACM SIGPLAN Notices, 50(12):59–70, December 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grimmer:2016:HPC

[GSS^{+}16] Matthias Grimmer, Chris Seaton, Roland Schatz, Thomas Würthinger, and Hanspeter Mössenböck. Highperformance cross-language interoperability in a multilanguage runtime. ACM SIGPLAN Notices, 51(2): 78–90, February 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grimmer:2018:CLI

[GSS^{+}18] Matthias Grimmer, Roland Schatz, Chris Seaton, Thomas Würthinger, and Mikel Luján. Cross-language interoperability in a multilanguage runtime. ACM Transactions on Programming Languages and Systems, 40(2):8:1–8:??, June 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gill:2010:MDP

[GT10a] Nasib Singh Gill and Pradeep Tomar. Modified development process of component-based software engineering. ACM SIG-SOFT Software Engineering Notes, 35(2):1–6, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goodrich:2010:DSA

[GT10b] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, fifth edition, 2010. ISBN 0-470-38326-7 (hardcover), 0-470-39880-9 (paperback). xxii + 714 pp. LCCN QA76.73.J38 G66 2010.

Geoffray:2010:VSM

[GTL^{+}10] Nicolas Geoffray, Gaël Thomas, Julia Lawall, Gilles Muller, and Bertil Folliot. VMKit: a substrate for managed runtime environments. ACM SIG-PLAN Notices, 45(7):51– 62, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gidra:2015:NGC

[GTS^{+}15] Lokesh Gidra, Gaël Thomas, Julien Sopena, Marc Shapiro, and Nhan Nguyen. NumaGiC: a garbage collector for big data on big NUMA machines. ACM SIG-PLAN Notices, 50(4):661– 673, April 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gidra:2011:ASG

[GTSS11] Lokesh Gidra, Gaël Thomas, Julien Sopena, and Marc Shapiro. Assessing the scalability of garbage collectors on many cores. Operating Systems Review, 5(3):15–19, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gunther:2014:ACC

[Gun14] John C. Gunther. Algorithm 938: Compressing circular buffers. ACM Transactions on Mathematical Software, 40(2):17:1–17:12, February 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Guo:2017:MJF

[Guo17] Robert Guo. MongoDB’s JavaScript fuzzer. Communications of the ACM, 60 (5):43–47, May 2017. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        cacm.acm.org/magazines/
                                        2017/5/216320/fulltext.

Guyer:2014:UJT

[Guy14] Samuel Z. Guyer. Use of the JVM at Twitter: a bird’s eye view. ACM SIGPLAN Notices, 49(11): 1, November 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gvero:2013:BRC

[Gve13]  Igor Gvero.  Book re-

view: Core Java volume I: fundamentals, 9th edition by Cay S. Horstmann and Gary Cornell. ACM SIG-SOFT Software Engineering Notes, 38(3):33, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gampe:2011:SMB

[GvRN^{+}11] Andreas Gampe, Jeffery von Ronne, David Niedzielski, Jonathan Vasek, and Kleanthis Psarris. Safe, multiphase bounds check elimination in Java. Software—Practice and Experience, 41(7):753–788, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grigore:2016:ARG

[GY16] Radu Grigore and Hongseok Yang. Abstraction refinement guided by a learnt probabilistic model. ACM SIGPLAN Notices, 51(1): 485–498, January 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garbervetsky:2011:QDM

[GYB^{+}11] Diego Garbervetsky, Sergio Yovine, Víctor Braberman, Martín Rouaux, and Alejandro Taboada. Quantitative dynamic-memory analysis for Java. Concurrency and Computation: Practice and Experience, 23(14): 1665–1678, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hauswirth:2013:TJP

[HA13] Matthias Hauswirth and Andrea Adamoli. Teaching Java programming with the Informa clicker system. Science of Computer Programming, 78(5):499–520, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642311001468.

Hanenberg:2015:WDW

[Han15] Stefan Hanenberg. Why do we know so little about programming languages, and what would have happened if we had known more? ACM SIGPLAN Notices, 50 (2):1, February 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hasbun:2012:UTP

[Has12] Javier E. Hasbun. Unifying two popular-but-seemingly-dissimilar platforms: Matlab and Java. Computing in Science and Engineering, 14(3):6–7, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haverbeke:2011:EJM

[Hav11] Marijn Haverbeke. Eloquent JavaScript: a modern introduction to programming. No Starch Press, San Francisco, CA, USA, 2011. ISBN 1-59327-282-0. xiv + 205 pp. LCCN QA76.73.J39 HAV 2011; QA76.73.J39 H38 2009. URL http:

//www.loc.gov/catdir/
enhancements/fy1012/2010032246-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy1012/2010032246-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1107/2010032246-
t.html.

Heumann:2013:TEM

[HAW13] Stephen T. Heumann, Vikram S.

Adve, and Shengjie Wang. The tasks with effects model for safe concurrency. ACM SIGPLAN Notices, 48(8): 239–250, August 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Huang:2013:ECS

[HB13] Jipeng Huang and Michael D. Bond. Efficient context sensitivity for dynamic analyses via calling context uptrees and customized memory management. ACM SIGPLAN Notices, 48(10): 53–72, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Hindle:2016:NS

[HBG^{+}16] Abram Hindle, Earl T. Barr, Mark Gabel, Zhendong Su, and Premkumar Devanbu. On the naturalness of software. Communications of the ACM, 59(5): 122–131, May 2016. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        cacm.acm.org/magazines/
                                        2016/5/201595/fulltext.

Hedin:2016:IFS

[HBS16] Daniel Hedin, Luciano Bello, and Andrei Sabelfeld. Information-flow security for JavaScript and its APIs. Journal of Computer Security, 24(2):181–234, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Heidegger:2012:APC

[HBT12] Phillip Heidegger, Annette Bieniusa, and Peter Thiemann. Access permission contracts for scripting languages. ACM SIG-PLAN Notices, 47(1):111– 122, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hsiao:2010:EST

[HC10] Chun-Feng Hsiao and Chih-Ping Chu. Enhancing SCORM through creating a reusable Java class repository. Software—Practice and Experience, 40(10):865– 881, September 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes-Croucher:2011:NRS

[HC11] Tom Hughes-Croucher. Node:

Up and Running: Scalable Server-Side Code with JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-9858-8. ???? pp. LCCN ????

Horstmann:2013:CJF

[HC13] Cay S. Horstmann and Gary Cornell. Core Java: Fundamentals, volume 1. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, ninth edition, 2013. ISBN 0-13-708189-8 (paperback). xxvi + 974 pp. LCCN QA76.73.J38 H6753 2013.

Herrera:2018:NCW

[HCLH18] David Herrera, Hanfeng Chen, Erick Lavoie, and Laurie Hendren. Numerical computing on the web: benchmarking for the future. ACM SIG-PLAN Notices, 53(8):88– 100, October 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3393673.3276968.

Hsiao:2014:UWC

[HCN14] Chun-Hung Hsiao, Michael Cafarella, and Satish Narayanasamy. Using web corpus statistics for program analysis. ACM SIGPLAN Notices, 49(10):49–65, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:2017:VOV

[HCV17] Matthew A. Hammer, Bor-Yuh Evan Chang, and David Van Horn. A vision for online verificationvalidation. ACM SIG-PLAN Notices, 52(3):190– 201, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Halder:2017:JSV

[HD17] Prithish Halder and Himadri Sekhar Das. JaSTA-2: Second version of the Java Superposition Tmatrix Application. Computer Physics Communications, 221(??):421–422, December 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465517302667.

Hofmann:2011:EOS

[HDK^{+}11] Owen S. Hofmann, Alan M.

Dunn, Sangman Kim, Indrajit Roy, and Emmett Witchel. Ensuring operating system kernel integrity with OSck. ACM SIG-PLAN Notices, 46(3):279– 290, March 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanazumi:2017:FAI

[HdM17] Simone Hanazumi and Ana C. V. de Melo. A formal approach to implement Java exceptions in cooperative systems. The Journal of systems and software, 131(??):475–490, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121216301285.

hunEom:2012:SSJ

[hED12] Yong hun Eom and Brian Demsky. Self-stabilizing Java. ACM SIGPLAN Notices, 47(6):287–298, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

hunEom:2012:DDP

[hEYJD12] Yong hun Eom, Stephen Yang, James C. Jenista, and Brian Demsky. DOJ: dynamically parallelizing object-oriented programs. ACM SIGPLAN Notices, 47(8):85–96, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Horspool:2011:PPP

[HGCA11] R. Nigel Horspool, Ralf Gitzel, Giacomo Cabri, and Markus Aleksy. Principles and practice of programming in Java 2008—special issue. Science of Computer Programming, 76(11):969, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hoppe:2013:DDB

[HH13] Michael Hoppe and Stefan Hanenberg. Do developers benefit from generic types?: an empirical comparison of generic and raw types in Java. ACM SIG-PLAN Notices, 48(10):457– 474, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Hower:2014:HRF

[HHB^{+}14] Derek R. Hower, Blake A.

Hechtman, Bradford M. Beckmann, Benedict R. Gaster, Mark D. Hill, Steven K. Reinhardt, and David A. Wood. Heterogeneousrace-free memory models. ACM SIGARCH Computer Architecture News, 42(1): 427–440, March 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Herhut:2013:RTP

[HHSS13] Stephan Herhut, Richard L. Hudson, Tatiana Shpeisman, and Jaswanth Sreeram. River Trail: a path to parallelism in JavaScript. ACM SIG-PLAN Notices, 48(10):729– 744, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Hinojosa:2013:TS

[Hin13] Daniel Hinojosa. Testing in Scala. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2013. ISBN 1-4493-1511-9, 1-4493-6034-3 (e-book), 1-4493-6033-5 (ebook). xiv + 148 pp. LCCN QA76.7 .H555 2013. URL

                                        http://proquest.tech.
                                        safaribooksonline.de/
                                        9781449360313;   http:
                                        //www.loc.gov/catdir/
                                        enhancements/fy1310/2012554271-
                                        b.html;   http://www.
                                        loc.gov/catdir/enhancements/
                                        fy1310/2012554271-d.html.

Hunt:2012:JP

[HJ12] Charlie Hunt and Binu John. Java Performance. The Java series. Addison-Wesley, Addison-Wesley, 2012. ISBN 0-13-714252-8 (paperback). xxi + 693 pp. LCCN QA76.73.J38 H845 2012.

Hellyer:2010:LCW

[HJH10] Laurence Hellyer, Richard Jones, and Antony L. Hosking. The locality of concurrent write barriers. ACM SIGPLAN Notices, 45(8): 83–92, August 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heidenreich:2010:GST

[HJS^{+}10] Florian Heidenreich, Jendrik Johannes, Mirko Seifert, Christian Wende, and Marcel Böhme. Generating safe template languages. ACM SIGPLAN Notices, 45(2): 99–108, February 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hlopko:2014:ISJ

[HKVG14] Marcel Hlopko, Jan Kurs, Jan Vraný, and Claus Gittinger. On the integration of Smalltalk and Java. Science of Computer Programming, 96 (part 1)(??):17–33, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642313002839.

Haddad:2013:SIP

[HL13] Ghaith Haddad and Gary T.

Leavens. Special issue papers: Specifying subtypes in Safety Critical Java programs. Concurrency and Computation: Practice and Experience, 25(16):2290– 2306, November 2013. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hague:2015:DRC

[HLO15] Matthew Hague, Anthony W.

Lin, and C.-H. Luke Ong. Detecting redundant CSS rules in HTML5 applications: a tree rewriting approach. ACM SIGPLAN Notices, 50(10):1–19, October 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herczeg:2013:TFF

[HLSK13] Zoltán Herczeg, Gábor Lóki, Tamás Szirbucz, and Ákos Kiss. Validating JavaScript guidelines across multiple Web browsers. Nordic Journal of Computing, 15(1):18–31, Summer 2013. CODEN NJCOFR. ISSN 1236-6064.

Herranz:2012:VIP

[HM12] Ángel Herranz and Julio Mariño. A verified implementation of priority monitors in Java. Lecture Notes in Computer Science, 7421: 160–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-31762-
                                        0_11/.

Huang:2012:RRC

[HMDE12] Wei Huang, Ana Milanova, Werner Dietl, and Michael D. Ernst. Reim & ReImInfer: checking and inference of reference immutability and method purity. ACM SIGPLAN Notices, 47(10):879–896, October 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hashmi:2012:CNI

[HNTL12] Atif Hashmi, Andrew Nere, James Jamal Thomas, and Mikko Lipasti. A case for neuromorphic ISAs. ACM SIGPLAN Notices, 47(4): 145–158, April 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horie:2014:SDJ

[HOKO14] Michihiro Horie, Kazunori Ogata, Kiyokuni Kawachiya, and Tamiya Onodera. String deduplication for Javabased middleware in virtualized environments. ACM SIGPLAN Notices, 49(7): 177–188, July 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hollingsworth:2012:SPI

[Hol12] Jeffrey Hollingsworth, editor. SC ’12: Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, Salt Lake Convention Center, Salt Lake City, UT, USA, November 10– 16, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0804-8.

Horstmann:2011:CJA

[Hor11] Cay S. Horstmann. Core Java: Advanced Topics, volume 2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, nineth edition, 2011. ISBN 0-13-708160-X. 1088 (est.) pp. LCCN ????

Horstmann:2012:JEC

[Hor12] Cay S. Horstmann. Java for everyone: compatible with Java 5, 6, and 7. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2012. ISBN 1-118-06331-7 (paperback). xxxiii + 589 pp. LCCN QA76.73.J38 H675445 2012.

Hosking:2012:CHL

[Hos12] Tony Hosking. Compiling a high-level language for GPUs: (via language support for architectures and compilers). ACM SIGPLAN Notices, 47(6): 1–12, June 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Hunt:2016:RFF

[HOSC16] Pete Hunt, Paul O’Shannessy, Dave Smith, and Terry Coatta. React: Facebook’s functional turn on writing JavaScript. ACM Queue: Tomorrow’s Computing Today, 14(4):40, July 2016.

CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL

http://queue.acm.org/
detail.cfm?id=2994373.

Haas:2017:BWS

[HRS^{+}17] Andreas Haas, Andreas Rossberg, Derek L. Schuff, Ben L. Titzer, Michael Holman, Dan Gohman, Luke Wagner, Alon Zakai, and JF Bastien. Bringing the web up to speed with WebAssembly. ACM SIG-PLAN Notices, 52(6):185– 200, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holzinger:2016:DSM

[HTBB16] P. Holzinger, S. Triller, A. Bartel, and E. Bodden. An in-depth study of more than ten years of Java exploitation. In ACM, editor, Proceedings of the ACM SIGSAC Conference on Computer and Communications Security, pages 779–790. ACM Press, New York, NY 10036, USA, 2016.

Higuera-Toledano:2010:ISI

[HTLC10] M. Teresa Higuera-Toledano, Doug Locke, and Angelo Corsaro. Introduction to special issue on Java technologies for real-time and embedded systems. ACM Transactions on Embedded Computing Systems, 10(1):

1:1–1:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Higuera-Toledano:2014:EIS

[HTW14] M. Teresa Higuera-Toledano and Andy Wellings. Editorials: Introduction to the Special Issue on Java Technologies for Real-Time and Embedded Systems: JTRES 2012. Concurrency and Computation: Practice and Experience, 26(14): 2405–2406, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hayashizaki:2012:IPT

[HWI^{+}12] Hiroshige Hayashizaki, Peng Wu, Hiroshi Inoue, Mauricio J. Serrano, and Toshio Nakatani. Improving the performance of trace-based systems by false loop filtering. ACM SIGPLAN Notices, 47(4):405–418, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2011:SBA

[HWLM11] Gang Huang, Weihu Wang, Tiancheng Liu, and Hong Mei. Simulation-based analysis of middleware service impact on system reliability: Experiment on Java application server. The Journal of systems and software, 84(7):1160–1170, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Haubl:2010:CES

[HWM10] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Compact and efficient strings for Java. Science of Computer Programming, 75(11):1077– 1094, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Haubl:2011:ECE

[HWM11] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Erratum to “Compact and Efficient Strings for Java” [Science of Computer Programming 75 (2010) 1077–1094]. Science of Computer Programming, 76(11):1073, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Haubl:2013:CST

[HWM13] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Contextsensitive trace inlining for Java. Computer Languages, Systems and Structures, 39(4):123–141, December 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1477842413000146.

Haubl:2014:TTE

[HWM14] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Trace transitioning and exception handling in a trace-based JIT compiler for Java. ACM Transactions on Architecture and Code Optimization, 11(1):6:1–6:??, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Humer:2015:DSL

[HWW^{+}15] Christian Humer, Christian Wimmer, Christian Wirth, Andreas Wöß, and Thomas Würthinger. A domainspecific language for building self-optimizing AST interpreters. ACM SIG-PLAN Notices, 50(3):123– 132, March 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hackett:2012:FPH

[HyG12] Brian Hackett and Shu yu Guo. Fast and precise hybrid type inference for JavaScript. ACM SIG-PLAN Notices, 47(6):239– 250, June 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Hua:2019:EED

[HZZK19] Jinru Hua, Yushan Zhang, Yuqun Zhang, and Sarfraz Khurshid. EdSketch: execution-driven sketching for Java. International Journal on Software Tools for Technology Transfer (STTT), 21 (3):249–265, June 2019. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

https://link.springer.
com/article/10.1007/s10009-
019-00512-8.

Iranmanesh:2016:SSE

[IF16] Zeinab Iranmanesh and Mehran S. Fallah. Specification and static enforcement of scheduler-independent noninterference in a middleweight Java. Computer Languages, Systems and Structures, 46(??):20– 43, November 2016. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842415300300.

Inoue:2012:AML

[IHWN12] Hiroshi Inoue, Hiroshige Hayashizaki, Peng Wu, and Toshio Nakatani. Adaptive multi-level compilation in a trace-based Java JIT compiler. ACM SIGPLAN Notices, 47(10):179–194, October 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inoue:2012:ISC

[IN12] Hiroshi Inoue and Toshio Nakatani. Identifying the sources of cache misses in Java programs without relying on hardware counters. ACM SIGPLAN Notices, 47(11):133–142, November 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’12 conference proceedings.

Islam:2012:HPR

[IRJ^{+}12] N. S. Islam, M. W. Rahman, J. Jose, R. Rajachandrasekar, H. Wang, H. Subramoni, C. Murthy, and D. K. Panda. High performance RDMA-based design of HDFS over InfiniBand. In Hollingsworth [Hol12], pages 35:1–35:12. ISBN 1-4673-0804-8. URL http:

                                        //conferences.computer.
                                        org/sc/2012/papers/1000a058.
                                        pdf.

Insa:2018:AAJ

[IS18] David Insa and Josep Silva. Automatic assessment of Java code. Computer Languages, Systems and Structures, 53(??):59–72, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842417301045.

Inostroza:2016:MIM

[IvdS16] Pablo Inostroza and Tijs van der Storm. Modular interpreters for the masses: implicit context propagation using object algebras. ACM SIGPLAN Notices, 51 (3):171–180, March 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Juneau:2012:JRP

[J^{+}12] Josh Juneau et al. Java 7 recipes: a problem-solution approach. Apress, Berkeley, CA, USA, 2012. ISBN 1-4302-4056-3 (paperback). xxii + 848 pp. LCCN QA76.73.J38 J35946 2012.

Joseph:2010:PII

[JACS10] Damien Joseph, Soon Ang, Roger H. L. Chang, and Sandra A. Slaughter. Practical intelligence in IT: assessing soft skills of IT professionals. Communications of the ACM, 53(2):149–154, February 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jaffer:2013:EAR

[Jaf13] Aubrey Jaffer. Easy accurate reading and writing of floating-point numbers. arXiv.org, ??(??): 1–7, October 28, 2013. CODEN ???? ISSN

???? URL http://arxiv.
org/abs/1310.8121.

Ji:2012:PKP

[JB12] Ran Ji and Richard Bubel. PE-KeY: a partial evaluator for Java programs. Lecture Notes in Computer Science, 7321:283– 295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-30729-
4_20/.

James:2010:FMC

[JC10] Perry R. James and Patrice Chalin. Faster and more complete extended static checking for the Java modeling language. Journal of Automated Reasoning, 44(1–2):??, February 2010. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10817-009-9134-
9.

Jacek:2019:OCW

[JCMM19] Nicholas Jacek, Meng-Chieh Chiu, Benjamin M. Marlin, and J. Eliot B. Moss. Optimal choice of when to garbage collect. ACM Transactions on Programming Languages and Systems, 41(1):3:1– 3:??, March 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jara:2012:NVJ

[JEC^{+}12] Carlos A. Jara, Francisco Esquembre, Wolfgang Christian, Francisco A. Candelas, Fernando Torres, and Sebastián Dormido. A new 3D visualization Java framework based on physics principles. Computer Physics Communications, 183(2):231–244, February 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465511002797.

Jendrock:2012:JET

[Jen12] Eric Jendrock. The Java EE 6 Tutorial: advanced topics, volume II. Addison-Wes-ley, Addison-Wesley, fourth edition, 2012. ISBN 0-13-708186-3 (paperback). ???? pp. LCCN ????

Jovic:2011:LLP

[JH11] Milan Jovic and Matthias Hauswirth. Listener latency profiling: Measuring the perceptible performance of interactive Java applications. Science of Computer Programming, 76(11):1054– 1072, November 1, 2011.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jenista:2011:OSO

[JhED11] James Christopher Jenista, Yong hun Eom, and Brian Charles Demsky. OoOJava: software out-of-order execution. ACM SIGPLAN Notices, 46(8):57–68, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Jeon:2019:MLA

[JJCO19] Minseok Jeon, Sehun Jeong, Sungdeok Cha, and Hakjoo Oh. A machine-learning algorithm with disjunctive model for data-driven program analysis. ACM Transactions on Programming Languages and Systems, 41 (2):13:1–13:??, June 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3293607.

Jayaraman:2017:CVJ

[JJL17] S. Jayaraman, B. Jayaraman, and D. Lessa. Compact visualization of Java program execution. Software—Practice and Experience, 47(2):163–191, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johari:2011:ESE

[JK11] Kalpana Johari and Arvinder Kaur. Effect of software evolution on software metrics: an open source case study. ACM SIG-SOFT Software Engineering Notes, 36(5):1–8, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jantz:2013:ESM

[JK13] Michael R. Jantz and Prasad A. Kulkarni. Exploring single and multilevel JIT compilation policy for modern machines 1. ACM Transactions on Architecture and Code Optimization, 10(4):22:1–22:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jagannathan:2014:ARV

[JLP^{+}14] Suresh Jagannathan, Vincent Laporte, Gustavo Petri, David Pichardie, and Jan Vitek. Atomicity refinement for verified compilation. ACM SIGPLAN Notices, 49(6):27, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jung:2012:EJA

[JMB12] Dong-Heon Jung, Soo-Mook Moon, and Sung-Hwan Bae. Evaluation of a Java ahead-of-time compiler for embedded systems. The Computer Journal, 55(2):232–252, February 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/2/232.
                                        full.pdf+html.

Jung:2014:HCO

[JMO14] Dong-Heon Jung, Soo-Mook Moon, and Hyeong-Seok Oh. Hybrid compilation and optimization for Java-based digital TV platforms. ACM Transactions on Embedded Computing Systems, 13(2s):62:1– 62:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Javed:2016:TSJ

[JQJ^{+}16] Ansar Javed, Bibrak Qamar, Mohsan Jameel, Aamir Shafi, and Bryan Carpenter. Towards scalable Java HPC with hybrid and native communication devices in MPJ express. International Journal of Parallel Programming, 44(6):1142–1172, December 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10766-015-0375-
4.

Johnsen:2012:SLM

[JTO12] Einar Broch Johnsen, Thi Mai Thuong Tran, and Olaf Owe. Safe locking for multi-threaded Java. Lecture Notes in Computer Science, 7141:158– 173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-29320-
7_11/.

Johnson:2015:EES

[JWMC15] Andrew Johnson, Lucas Waye, Scott Moore, and Stephen Chong. Exploring and enforcing security guarantees via program dependence graphs. ACM SIG-PLAN Notices, 50(6):291– 302, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jin:2012:JMM

[JYKS12] Huafeng Jin, Tuba Yavuz-Kahveci, and Beverly A. Sanders. Java memory model-aware model checking. Lecture Notes in Computer Science, 7214: 220–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-28756-
                                        5_16/.

Kossakowski:2012:JED

[KARO12] Grzegorz Kossakowski, Nada Amin, Tiark Rompf, and Martin Odersky. JavaScript as an embedded DSL. Lecture Notes in Computer Science, 7313:409– 434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-31057-
                                        7_19/.

Kastner:2012:TCA

[KATS12] Christian Kästner, Sven Apel, Thomas Thüm, and Gunter Saake. Type checking annotation-based product lines. ACM Transactions on Software Engineering and Methodology, 21 (3):14:1–14:??, June 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kumari:2011:AOO

[KB11] Usha Kumari and Sucheta Bhasin. Application of object-oriented metrics to C++ and Java: a comparative study. ACM SIG-SOFT Software Engineering Notes, 36(2):1–10, March 2011. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Kunjir:2017:TAM

[KB17] Mayuresh Kunjir and Shivnath Babu. Thoth in action: memory management in modern data analytics. Proceedings of the VLDB Endowment, 10(12): 1917–1920, August 2017. CODEN ???? ISSN 2150-8097.

Kim:2014:LBL

[KBL14] Hongjune Kim, Seonmyeong Bak, and Jaejin Lee. Lightweight and block-level concurrent sweeping for JavaScript garbage collection. ACM SIGPLAN Notices, 49(5):155–164, May 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:2017:SFC

[KBPS17] Oleg Kiselyov, Aggelos Biboudis, Nick Palladinos, and Yannis Smaragdakis. Stream fusion, to completeness. ACM SIGPLAN Notices, 52(1):285–299, January 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2012:MCO

[KC12] Sameer Kulkarni and John Cavazos. Mitigating the compiler optimization phaseordering problem using machine learning. ACM SIG-PLAN Notices, 47(10):147– 162, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnaveni:2012:HOJ

[KCD12] R. Krishnaveni, C. Chellappan, and R. Dhanalakshmi. Hybrid obfuscated Javascript strength analysis system for detection of malicious websites. Lecture Notes in Computer Science, 7513: 129–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-35606-
3_15/.

Kedia:2017:SFS

[KCP^{+}17] Piyus Kedia, Manuel Costa, Matthew Parkinson, Kapil Vaswani, Dimitrios Vytiniotis, and Aaron Blankstein. Simple, fast, and safe manual memory management. ACM SIGPLAN Notices, 52(6):233–247, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kouzapas:2018:TPM

[KDPG18] Dimitrios Kouzapas, Ornela Dardha, Roly Perera, and Simon J. Gay. Typechecking protocols with Mungo and StMungo: A session type toolchain for Java. Science of Computer Programming, 155(??):52–75, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642317302186.

Kereki:2015:JAW

[Ker15] Federico Kereki. JavaScript all the way down. Linux Journal, 2015(250):1:1– 1:??, February 2015. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/dl.acm.org/ft_gateway.
cfm?id=2754898.

Kuehnhausen:2011:AJM

[KF11] Martin Kuehnhausen and Victor S. Frost. Application of the Java Message Service in mobile monitoring environments. Journal of Network and Computer Applications, 34(5):1707–1716, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1084804511001159.

Kumar:2012:WSB

[KFB^{+}12] Vivek Kumar, Daniel Frampton, Stephen M. Blackburn, David Grove, and Olivier Tardieu. Workstealing without the baggage. ACM SIGPLAN Notices, 47(10):297–314, October 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khan:2015:UJW

[KFBK^{+}15] Faiz Khan, Vincent Foley-Bourgon, Sujay Kathrotia, Erick Lavoie, and Laurie Hendren. Using JavaScript and WebCL for numerical computations: a comparative study of native and web technologies. ACM SIG-PLAN Notices, 50(2):91– 102, February 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knoche:2018:UML

[KH18] H. Knoche and W. Hasselbring. Using microservices for legacy software modernization. IEEE Software, 35 (3):44–49, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kerschbaumer:2013:IFT

[KHL^{+}13] Christoph Kerschbaumer, Eric Hennigan, Per Larsen, Stefan Brunthaler, and Michael Franz. Information flow tracking meets justin-time compilation. ACM Transactions on Architecture and Code Optimization, 10(4):38:1–38:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kang:2017:PSR

[KHL^{+}17] Jeehoon Kang, Chung-Kil Hur, Ori Lahav, Viktor Vafeiadis, and Derek Dreyer. A promising semantics for relaxed-memory concurrency. ACM SIG-PLAN Notices, 52(1):175– 189, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kalibera:2011:FRT

[KHM^{+}11] Tomas Kalibera, Jeff Hagelberg, Petr Maj, Filip Pizlo, Ben Titzer, and Jan Vitek. A family of real-time Java benchmarks. Concurrency and Computation: Practice and Experience, 23(14):1679–1700, September 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabanov:2011:DSF

[KHR11] Jevgeni Kabanov, Michael Hunger, and Rein Raudjärv. On designing safe and flexible embedded DSLs with Java 5. Science of Computer Programming, 76(11): 970–991, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kienle:2010:ATT

[Kie10] Holger M. Kienle. It’s about time to take JavaScript (more) seriously. IEEE Software, 27(3):60–62, May/ June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kienle:2013:BRE

[Kie13] Holger M. Kienle. Book review: Effective JavaScript: 68 specific ways to harness the power of JavaScript by David Herman. ACM SIGSOFT Software Engineering Notes, 38(6):40–41, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:2017:TAA

[KKK^{+}17] Channoh Kim, Jaehyeok Kim, Sungmin Kim, Dooyoung Kim, Namho Kim, Gitae Na, Young H. Oh, Hyeon Gyu Cho, and Jae W. Lee. Typed architectures: Architectural support for lightweight scripting. ACM SIGARCH Computer Architecture News, 45(1):77–90, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Krieger:2011:AES

[KKW11] Matthias P. Krieger, Alexander Knapp, and Burkhart Wolff. Automatic and efficient simulation of operation contracts. ACM SIG-PLAN Notices, 46(2):53–62, February 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaiser:2014:WAM

[KKW14] Alexander Kaiser, Daniel Kroening, and Thomas Wahl. A widening approach to multithreaded program verification. ACM Transactions on Programming Languages and Systems, 36(4): 14:1–14:??, October 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ko:2010:EAW

[KM10] Andrew J. Ko and Brad A.

Myers. Extracting and answering why and why not questions about Java program output. ACM Transactions on Software Engineering and Methodology, 20(2):4:1–4:??, August 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Karakoidas:2015:TSE

[KMLS15] Vassilios Karakoidas, Dimitris Mitropoulos, Panagiotis Louridas, and Diomidis Spinellis. A type-safe embedding of SQL into Java using the extensible compiler framework J%. Computer Languages, Systems and Structures, 41(??):1– 20, April 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842415000020.

Kalibera:2014:FAS

[KMMV14] Tomas Kalibera, Petr Maj, Floreal Morandat, and Jan Vitek. A fast abstract syntax tree interpreter for R. ACM SIGPLAN Notices, 49 (7):89–102, July 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulkarni:2016:APA

[KMZN16] Sulekha Kulkarni, Ravi Mangal, Xin Zhang, and Mayur Naik. Accelerating program analyses by crossprogram training. ACM SIGPLAN Notices, 51(10): 359–377, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kolling:2010:GPE

[Köl10]  Michael Kölling.  The

Greenfoot programming environment. ACM Transactions on Computing Education, 10(4):14:1–14:??, November 2010. CODEN ???? ISSN 1946-6226.

Kroening:2015:CAV

[KP15] Daniel Kroening and Corina S. Păsăreanu, editors. Computer Aided Verification: 27th International Conference, CAV 2015, San Francisco, CA, USA, July 18–24, 2015, Proceedings, Part I, volume 9206 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-21689-9. URL

http://link.springer.
com/book/10.1007/978-3-
319-21690-4.

Kalibera:2011:SRT

[KPHV11] Tomas Kalibera, Filip Pizlo, Antony L. Hosking, and Jan Vitek. Scheduling realtime garbage collection on uniprocessors. ACM Transactions on Computer Systems, 29(3):8:1–8:??, August 2011. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Khyzha:2012:AP

[KPP12] Artem Khyzha, Pavel Parízek, and Corina S. Păsăreanu. Abstract pathfinder. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kintis:2018:HEM

[KPP^{+}18] Marinos Kintis, Mike Papadakis, Andreas Papadopoulos, Evangelos Valvis, Nicos Malevris, and Yves Le Traon. How effective are mutation testing tools? An empirical analysis of Java mutation testing tools with manual analysis and real faults. Empirical Software Engineering, 23(4):2426– 2463, August 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s10664-017-9582-
                                        5.

Kang:2012:FSJ

[KR12] Seonghoon Kang and Sukyoung Ryu. Formal specification of a JavaScript module system. ACM SIG-PLAN Notices, 47(10):621– 638, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kedlaya:2014:DDL

[KRCH14] Madhukar N. Kedlaya, Behnam Robatmili, Cglin Cascaval, and Ben Hardekopf. Deoptimization for dynamic language JITs on typed, stack-based virtual machines. ACM SIG-PLAN Notices, 49(7):103– 114, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kedlaya:2016:SST

[KRH16] Madhukar N. Kedlaya, Behnam Robatmili, and Ben Hardekopf. Server-side type profiling for optimizing client-side JavaScript engines. ACM SIGPLAN Notices, 51(2):140–153, February 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishnamurthi:2012:SAJ

[Kri12] Shriram Krishnamurthi. Semantics and analyses for JavaScript and the Web. Lecture Notes in Computer Science, 7460:4, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.
com/accesspage/chapter/
10.1007/978-3-642-33125-
1_3.

Kedlaya:2014:ITS

[KRR^{+}14] Madhukar N. Kedlaya, Jared Roesch, Behnam Robatmili, Mehrdad Reshadi, and Ben Hardekopf. Improved type specialization for dynamic scripting languages. ACM SIGPLAN Notices, 49(2):37–48, February 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Ko:2019:WSA

[KRR19] Yoonseok Ko, Xavier Rival, and Sukyoung Ryu. Weakly sensitive analysis for JavaScript objectmanipulating programs. Software—Practice and Experience, 49(5):840–884, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaufmann:2013:SCO

[KS13] Marco Kaufmann and Rainer G. Spallek. Superblock compilation and other optimization techniques for a Javabased DBT machine emulator. ACM SIGPLAN Notices, 48(7):33–40, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Krebs:2014:JJB

[KS14] Nico Krebs and Lothar Schmitz. Jaccie: a Javabased compiler-compiler for generating, visualizing and debugging compiler components. Science of Computer Programming, 79(??):101– 115, January 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642312000469.

Kroshko:2015:OPN

[KS15] Andrew Kroshko and Raymond J. Spiteri. odeTo-Java: a PSE for the numerical solution of IVPs. ACM Transactions on Mathematical Software, 41(3):17:1– 17:33, June 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kouneli:2012:MKD

[KSPK12] Aggeliki Kouneli, Georgia Solomou, Christos Pierrakeas, and Achilles Kameas. Modeling the knowledge domain of the Java programming language as an ontology. Lecture Notes in Computer Science, 7558:152– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-33642-
3_16/.

Korsholm:2014:RTJ

[KSR14] Stephan E. Korsholm, Hans Søndergaard, and Anders P. Ravn. A real-time Java tool chain for resource constrained platforms. Concurrency and Computation: Practice and Experience, 26(14):2407–2431, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kashyap:2014:TRS

[KSW^{+}14] Vineeth Kashyap, John Sarracino, John Wagner, Ben Wiedermann, and Ben Hardekopf. Type refinement for static analysis of JavaScript. ACM SIG-PLAN Notices, 49(2):17–26, February 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Keil:2014:EDA

[KT14] Matthias Keil and Peter Thiemann. Efficient dynamic access analysis using JavaScript proxies. ACM SIGPLAN Notices, 49(2): 49–60, February 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Keil:2015:BAH

[KT15] Matthias Keil and Peter Thiemann. Blame assignment for higher-order contracts with intersection and union. ACM SIGPLAN Notices, 50(9):375–386, September 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kersten:2014:RRA

[KvGS^{+}14] Rody W. J. Kersten, Bernard E. van Gastel, Olha Shkaravska, Manuel Montenegro, and Marko C. J. D. van Eekelen. ResAna: a resource analysis toolset for (real-time) JAVA. Concurrency and Computation: Practice and Experience, 26(14):2432–2455, September 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolesnikov:2014:CPB

[KvRHA14] Sergiy Kolesnikov, Alexander von Rhein, Claus Hunsen, and Sven Apel. A comparison of product-based, feature-based, and familybased type checking. ACM SIGPLAN Notices, 49(3): 115–124, March 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2010:EAE

[KW10] Minseong Kim and Andy Wellings. Efficient asynchronous event handling in the real-time specification for Java. ACM Transactions on Embedded Computing Systems, 10(1):5:1– 5:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2011:MAE

[KW11] Minseong Kim and Andy Wellings. Multiprocessors and asynchronous event handling in the real-time specification for Java. The Computer Journal, 54(8):1308–1324, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1308.
                                        full.pdf+html.

Lin:2012:UKT

[LBF12] Yi Lin, Stephen M. Blackburn, and Daniel Frampton. Unpicking the knot: teasing apart VM/application interdependencies. ACM SIGPLAN Notices, 47(7): 181–190, July 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Lauinger:2018:TSD

[LCW18] Tobias Lauinger, Abdelberi Chaabane, and Christo B. Wilson. Thou shalt not depend on me. Communications of the ACM, 61(6): 41–47, June 2018. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        cacm.acm.org/magazines/
                                        2018/6/228039/fulltext.

Li:2014:MHD

[LDL14] Pengcheng Li, Chen Ding, and Hao Luo. Modeling heap data growth using average liveness. ACM SIG-PLAN Notices, 49(11):71– 82, November 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lorenzen:2016:STD

[LE16] Florian Lorenzen and Sebastian Erdweg. Sound type-dependent syntactic language extension. ACM SIGPLAN Notices, 51(1): 204–216, January 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leijen:2017:TDC

[Lei17] Daan Leijen. Type directed compilation of row-typed algebraic effects. ACM SIG-PLAN Notices, 52(1):486– 499, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2010:FTJ

[Ler10] Reuven M. Lerner. At the forge: testing JavaScript. Linux Journal, 2010(191): 6:1–6:??, March 2010. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Lewis:2013:IAP

[Lew13] Mark C. Lewis. Introduction to the art of programming using Scala. Chapman and Hall/CRC textbooks in computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4398-9666-6 (paperback). ???? pp. LCCN QA76.73.S28 L49 2013.

Liu:2019:RIP

[LHR19] Bozhen Liu, Jeff Huang, and Lawrence Rauchwerger. Rethinking incremental and parallel pointer analysis. ACM Transactions on Programming Languages and Systems, 41 (1):6:1–6:??, March 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liu:2014:JNU

[Liu14] Alex Liu. JavaScript and the Netflix user interface. ACM Queue: Tomorrow’s Computing Today, 12(9):20, September 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Liva:2019:SDE

[LKP19] Giovanni Liva, Muhammad Taimoor Khan, and Martin Pinzger. Semanticsdriven extraction of timed automata from Java programs. Empirical Software Engineering, 24(5): 3114–3150, October 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/article/10.1007/s10664-
019-09699-5;   http://
link.springer.com/content/
pdf/10.1007/s10664-019-
09699-5.pdf.

Leino:2015:APS

[LL15] K. Rustan M. Leino and Paqui Lucio. An assertional proof of the stability and correctness of Natural Mergesort. ACM Transactions on Computational Logic, 17(1):6:1–6:22, December 2015. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Leung:2013:PEJ

[LLL13] Alan Leung, Ondrej Lhoták, and Ghulam Lashari. Parallel execution of Java loops on Graphics Processing Units. Science of Computer Programming, 78(5): 458–480, May 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642311001444.

Lin:2015:STU

[LM15] Yixiao Lin and Sayan Mitra. StarL: Towards a unified framework for programming, simulating and verifying distributed robotic systems. ACM SIGPLAN Notices, 50(5):9:1–9:??, May 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2016:ECP

[LMK16] Seong-Won Lee, Soo-Mook Moon, and Seong-Moo Kim. Extended conference papers: Flow-sensitive runtime estimation: an enhanced hot spot detection heuristics for embedded Java just-in-time compilers. Software—Practice and Experience, 46(6):841– 864, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loring:2017:SAJ

[LML17] Matthew C. Loring, Mark Marron, and Daan Leijen. Semantics of asynchronous JavaScript. ACM SIG-PLAN Notices, 52(11):51– 62, November 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Long:2012:COS

[LMS^{+}12] F. W. (Frederick W.) Long, Dhruv Mohindra, Robert Seacord, Dean Sutherland, and David Svoboda. The CERT Oracle secure coding standard for Java. The SEI series in software engineering. Addison-Wesley, Addison-Wesley, 2012. ISBN 0-321-80395-7 (paperback). xxxiv + 699 pp. LCCN QA76.73.J38 C44 2012.

URL http://ptgmedia.
pearsoncmg.com/images/
9780321803955/samplepages/
0321803957.pdf.

Long:2013:JCG

[LMS^{+}13] Fred Long, Dhruv Mohindra, Robert Seacord, Dean Sutherland, and David Svoboda. Java coding guidelines: 75 recommendations for reliable and secure programs. The SEI series in software engineering. Addison-Wesley, Reading, MA, USA, 2013. ISBN 0-13-343952-6, 0-13-343954-2 (e-book), 0-321-93315-X (paperback). LCCN QA76.73.J38 L664 2014.

URL http://0proquest.
safaribooksonline.com/
9780133439526.

Luo:2019:HDS

[LMZP19] Q. Luo, K. Moran, L. Zhang, and D. Poshyvanyk. How do static and dynamic test case prioritization techniques perform on modern software systems? An extensive study on GitHub projects. IEEE Transactions on Software Engineering, 45(11):1054–1080, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Leavens:2015:BSS

[LN15] Gary T. Leavens and David A. Naumann. Behavioral subtyping, specification inheritance, and modular reasoning. ACM Transactions on Programming Languages and Systems, 37(4):13:1–13:??, August 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lopes:2015:HSA

[LO15] Cristina V. Lopes and Joel Ossher. How scale affects structure in Java programs. ACM SIGPLAN Notices, 50(10):675–694, October 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lochbihler:2013:MJM

[Loc13] Andreas Lochbihler. Making the Java memory model safe. ACM Transactions on Programming Languages and Systems, 35(4):12:1– 12:??, December 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lochbihler:2018:MTS

[Loc18] Andreas Lochbihler. Mechanising a type-safe model of multithreaded Java with a verified compiler. Journal of Automated Reasoning, 61(1–4):243–332, June 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10817-018-9452-
x.

Long:2010:TDSa

[Lon10a] Brad Long. Towards the design of a set-based Java collections framework. ACM SIGSOFT Software Engineering Notes, 35(5):1–7, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Long:2010:TDSb

[Lon10b] Brad Long. Towards the design of a set-based Java collections framework. ACM SIGSOFT Software Engineering Notes, 35(6):1–7, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loureiro:2013:EDS

[LPA13] André Loureiro, João Paulo Porto, and Guido Araujo. Extending decoupled software pipeline to parallelize Java programs. Software —Practice and Experience, 43(5):525–541, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lerner:2014:TRT

[LPGK14] Benjamin S. Lerner, Joe Gibbs Politz, Arjun Guha, and Shriram Krishnamurthi. Te-JaS: retrofitting type systems for JavaScript. ACM SIGPLAN Notices, 49(2):1– 16, February 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Lux:2011:TSD

[LS11] Alexander Lux and Artem Starostin. A tool for static detection of timing channels in Java. Journal of Cryptographic Engineering, 1(4):303–313, December 2011. CODEN ???? ISSN 2190-8508 (print), 2190-8516 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s13389-011-0021-
                                        z.

Landman:2016:EAR

[LSBV16] Davy Landman, Alexander Serebrenik, Eric Bouwers, and Jurgen J. Vinju. Empirical analysis of the relationship between CC and SLOC in a large corpus of Java methods and C functions. Journal of Software: Evolution and Process, 28 (7):589–618, July 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic). See corrigendum [LSBV17].

Landman:2017:CEA

[LSBV17] Davy Landman, Alexander Serebrenik, Eric Bouwers, and Jurgen Vinju. Corrigendum: Empirical analysis of the relationship between CC and SLOC in a large corpus of Java methods and C functions published on 9 December 2015. Journal of Software: Evolution and Process, 29(10): ??, October 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic). See [LSBV16].

Larrucea:2018:M

[LSCPE18] X. Larrucea, I. Santamaria, R. Colomo-Palacios, and C. Ebert. Microservices. IEEE Software, 35(3):96– 100, May/June 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Luu:2014:MCC

[LSSD14] Loi Luu, Shweta Shinde, Prateek Saxena, and Brian Demsky. A model counter for constraints over unbounded strings. ACM SIG-PLAN Notices, 49(6):565– 576, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leopoldseder:2016:JJT

[LSWM16] David Leopoldseder, Lukas Stadler, Christian Wimmer, and Hanspeter Mössenböck. Java-to-JavaScript translation via structured control flow reconstruction of compiler IR. ACM SIG-PLAN Notices, 51(2):91– 103, February 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2011:JEC

[LT11] Siliang Li and Gang Tan. JET: exception checking in the Java Native Interface. ACM SIGPLAN Notices, 46 (10):345–358, October 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Li:2014:EAJ

[LT14] Siliang Li and Gang Tan. Exception analysis in the Java Native Interface. Science of Computer Programming, 89 (part C)(??):273– 297, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167642314000446.

Laskowski:2012:DJP

[LTD^{+}12] Eryk Laskowski, Marek Tudruj, Ivanoe De Falco, Umberto Scafuri, and Ernesto Tarantino. Distributed Java programs initial mapping based on extremal optimization. Lecture Notes in Computer Science, 7133: 75–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-28151-
8_8/.

Luckow:2017:HTP

[LTK17] Kasper Søe Luckow, Bent Thomsen, and Stephan Erbs Korsholm. HVM_{T P}: a time predictable and portable Java Virtual Machine for hard real-time embedded systems. Concurrency and Computation: Practice and Experience, 29(22): ??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2018:PGC

[LTMS18] Yue Li, Tian Tan, Anders Møller, and Yannis Smaragdakis. Precision-guided context sensitivity for pointer analysis. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):141:1–141:29, October 2018. URL

https://dl.acm.org/doi/
abs/10.1145/3276511.

Liu:2014:FFL

[LTZ14] Peng Liu, Omer Tripp, and Xiangyu Zhang. Flint: fixing linearizability violations. ACM SIGPLAN Notices, 49(10):543–560, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:2010:SDT

[LVG10] Benjamin S. Lerner, Herman Venter, and Dan Grossman. Supporting dynamic, third-party code customizations in JavaScript using aspects. ACM SIG-PLAN Notices, 45(10):361– 376, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2015:SGU

[LWB^{+}15] Yi Lin, Kunshan Wang, Stephen M. Blackburn, Antony L. Hosking, and Michael Norrish. Stop and go: understanding yieldpoint behavior. ACM SIG-PLAN Notices, 50(11):70– 80, November 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luckcuck:2017:SCJ

[LWC17] Matt Luckcuck, Andy Wellings, and Ana Cavalcanti. Safety-Critical Java: level 2 in practice. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2010:JSD

[LWH^{+}10] Byeongcheol Lee, Ben Wiedermann, Martin Hirzel, Robert Grimm, and Kathryn S. McKinley. Jinn: synthesizing dynamic bug detectors for foreign language interfaces. ACM SIG-PLAN Notices, 45(6):36– 49, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2019:PPY

[LX19] Jingbo Lu and Jingling Xue. Precision-preserving yet fast object-sensitive pointer analysis with partial context sensitivity. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 148:1–148:29, October 2019.

URL  https://dl.acm.
org/doi/abs/10.1145/3360574.

Li:2018:ATJ

[LXP18] Bing Li, Xueli Xiao, and Yi Pan. Automatic translation from Java to Spark. Concurrency and Computation: Practice and Experience, 30(20):e4459:1– e4459:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lindholm:2013:JVMa

[LYBB13a] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification. The Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2013. ISBN 0-13-326044-5 (paperback). ???? pp. LCCN ????

Lindholm:2013:JVMb

[LYBB13b] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification. Addison-Wesley, Addison-Wesley, Java SE 7 edition, 2013. ISBN 0-13-326049-6, 0-13-326044-5. xvii + 587 (est.) pp. LCCN QA76.73.J38 L56 1999.

                                        URL http://proquest.
                                        tech.safaribooksonline.
                                        de/9780133260496.

Lindholm:2014:JVM

[LYBB14] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification: Java SE 8 edition. Addison-Wesley, Addison-Wesley, 2014. ISBN 0-13-390590-X (paperback), 0-13-392274-X (e-book). xvi + 584 pp. LCCN QA76.73.J38 L56 2014.

Liu:2018:JIO

[LYM^{+}18] Xuanzhe Liu, Meihua Yu, Yun Ma, Gang Huang, Hong Mei, and Yunxin Liu. i-Jacob: an internetwareoriented approach to optimizing computation-intensive mobile Web browsing. ACM Transactions on Internet Technology (TOIT), 18(2): 14:1–14:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lyon:2012:JTW

[Lyo12] Doug Lyon. The Java tree withers. Computer, 45(1):83–85, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2012:PAA

[LZ12] P. Liu and C. Zhang. Pert:

The application-aware tailoring of Java object persistence. IEEE Transactions on Software Engineering, 38(4):909–922, July/August 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
5963692.

Li:2016:JJM

[LZYP16] Bing Li, Junbo Zhang, Ning Yu, and Yi Pan. J2M: a Java to MapReduce translator for cloud computing. The Journal of Supercomputing, 72(5):1928– 1945, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s11227-016-1695-
x.

McIntosh:2012:EJB

[MAH12] Shane McIntosh, Bram Adams, and Ahmed E. Hassan. The evolution of Java build systems. Empirical Software Engineering, 17(4– 5):578–608, August 2012. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/accesspage/article/
10.1007/s10664-011-9169-
5.

Maas:2016:THL

[MAHK16] Martin Maas, Krste Asanović, Tim Harris, and John Kubiatowicz. Taurus: a holistic language runtime system for coordinating distributed managed-language applications. Operating Systems Review, 50(2):457– 471, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maas:2019:HAT

[MAK19] M. Maas, K. Asanovic, and J. Kubiatowicz. A hardware accelerator for tracing garbage collection. IEEE Micro, 39(3):38–46, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McIntyre:2012:FJB

[MB12] Timothy J. McIntyre and Alexis I. Bishop. Fringe — a Java-based finite fringe analysis package. Computer Physics Communications, 183(9):2014–2018, September 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465512001506.

Martinez:2017:MBA

[MCC17] Salvador Martínez, Valerio Cosentino, and Jordi Cabot. Model-based analysis of Java EE web security misconfigurations. Computer Languages, Systems and Structures, 49(??):36– 61, September 2017. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842416301348.

McKinley:2016:PWU

[McK16] Kathryn S. McKinley. Programming the world of uncertain things (keynote). ACM SIGPLAN Notices, 51 (1):1–2, January 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McMillan:2011:SVM

[McM11] William W. McMillan. The soul of the virtual machine. IEEE Spectrum, 48(7):44– 59, July 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McNerney:2019:BBB

[McN19] P. J. McNerney. Beginning Bazel: Building and Testing for Java, Go, and More. Apress, Berkeley, CA, USA, 2019. ISBN 1-4842-5194-6. xi + 192 pp. LCCN QA75.5-

                                        76.95. URL http://www.
                                        vlebooks.com/vleweb/product/
                                        openreader?id=none\%26isbn=
                                        9781484251942.

Miyazawa:2019:SCS

[MCW19] Alvaro Miyazawa, Ana Cavalcanti, and Andy Wellings. SCJ-Circus: Specification and refinement of safetycritical Java programs. Science of Computer Programming, 181(??):140–176, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167642319300012.

McLane:2010:UIV

[MCY^{+}10] Jonathan C. McLane, W. Walter Czech, David A. Yuen, Mike R. Knox, Shuo Wang, Jim B. S. Greensky, and Erik O. D. Sevre. Ubiquitous interactive visualization of large-scale simulations in geosciences over a Java-based web-portal. Concurrency and Computation: Practice and Experience, 22(12):1750–1773, August 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marr:2015:TVP

[MD15] Stefan Marr and Stéphane Ducasse. Tracing vs. partial evaluation: comparing meta-compilation approaches for self-optimizing interpreters. ACM SIG-PLAN Notices, 50(10):821– 839, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mytkowicz:2010:EAJ

[MDHS10] Todd Mytkowicz, Amer Diwan, Matthias Hauswirth, and Peter F. Sweeney. Evaluating the accuracy of Java profilers. ACM SIG-PLAN Notices, 45(6):187– 197, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marr:2017:CLC

[MDM17] Stefan Marr, Benoit Daloze, and Hanspeter Mössenböck. Cross-language compiler benchmarking: are we fast yet? ACM SIGPLAN Notices, 52(2):120–131, February 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martinez:2017:ARR

[MDS^{+}17] Matias Martinez, Thomas Durieux, Romain Sommerard, Jifeng Xuan, and Martin Monperrus. Automatic repair of real bugs in Java: a large-scale experiment on the defects4j dataset. Empirical Software Engineering, 22(4):1936–1964, August 2017. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s10664-016-9470-
                                        4.

Meijer:2014:EJR

[Mei14] Erik Meijer. Evolution of Javascript revisited. ACM Queue: Tomorrow’s Computing Today, 12(2):50, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Martinsen:2014:HTL

[MGI14] Jan Kasper Martinsen, Hakan Grahn, and Anders Isberg. Heuristics for thread-level speculation in Web applications. IEEE Computer Architecture Letters, 13(2):77–80, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Martinsen:2017:CTL

[MGI17] Jan Kasper Martinsen, Håkan Grahn, and Anders Isberg. Combining thread-level speculation and just-in-time compilation in Google’s V8 JavaScript engine. Concurrency and Computation: Practice and Experience, 29(1):??, January 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehrabi:2019:PUP

[MGS19] Mostafa Mehrabi, Nasser Giacaman, and Oliver Sinnen. @PT: Unobtrusive parallel programming with Java annotations. Concurrency and Computation: Practice and Experience, 31 (1):e4831:1–e4831:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miller:2013:IPG

[MHBO13] Heather Miller, Philipp Haller, Eugene Burmako, and Martin Odersky. Instant pickles: generating object-oriented pickler combinators for fast and extensible serialization. ACM SIGPLAN Notices, 48(10): 183–202, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Matsakis:2015:TOJ

[MHL15] Nicholas D. Matsakis, David Herman, and Dmitry Lomov. Typed objects in JavaScript. ACM SIG-PLAN Notices, 50(2):125– 134, February 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGachey:2010:CJC

[MHM10] Phil McGachey, Antony L. Hosking, and J. Eliot B. Moss. Classifying Java class transformations for pervasive virtualized access. ACM SIGPLAN Notices, 45(2):75–84, February 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mastrangelo:2019:CAD

[MHN19] Luis Mastrangelo, Matthias Hauswirth, and Nathaniel Nystrom. Casting about in the dark: an empirical study of cast operations in Java programs. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):158:1–158:31, October 2019. URL

https://dl.acm.org/doi/
abs/10.1145/3360584.

Mayer:2012:ESI

[MHR^{+}12] Clemens Mayer, Stefan Hanenberg, Romain Robbes, Éric Tanter, and Andreas Stefik. An empirical study of the influence of static type systems on the usability of undocumented software. ACM SIGPLAN Notices, 47(10):683–702, October 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:2013:TSG

[Mil13] Mark Miller. A tested semantics for getters, setters, and eval in JavaScript. ACM SIGPLAN Notices, 48(2):1–16, February 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malhotra:2017:PPS

[MKG^{+}17] Geetika Malhotra, Rajshekar Kalayappan, Seep Goel, Pooja Aggarwal, Abhishek Sagar, and Smruti R. Sarangi. ParTejas: a parallel simulator for multicore processors. ACM Transactions on Modeling and Computer Simulation, 27(3): 19:1–19:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Misra:2012:JSC

[MKK^{+}12] Janardan Misra, Annervaz KM, Vikrant Kaulgud, Shubhashis Sengupta, and Gary Titus. Java sourcecode clustering: Unifying syntactic and semantic features. ACM SIGSOFT Software Engineering Notes, 37 (6):1–8, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2013:JSC

[MKK^{+}13] Janardan Misra, Annervaz KM, Vikrant Kaulgud, Shubhashis Sengupta, and Gary Titus. Java sourcecode clustering: Unifying syntactic and semantic features. ACM SIGSOFT Software Engineering Notes, 38 (1):41–43, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mazinanian:2017:UUL

[MKTD17] Davood Mazinanian, Ameya Ketkar, Nikolaos Tsantalis, and Danny Dig. Understanding the use of lambda expressions in Java. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA): 85:1–85:??, October 2017. CODEN ???? ISSN 2475-1421.

Marek:2014:SRC

[MKZ^{+}14] Lukás Marek, Stephen Kell, Yudi Zheng, Lubomír Bulej, Walter Binder, Petr Tuma, Danilo Ansaloni, Aibek Sarimbekov, and Andreas Sewe. ShadowVM: robust and comprehensive dynamic program analysis for the Java platform. ACM SIG-PLAN Notices, 49(3):105– 114, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martinez-Llario:2011:DJS

[MLGA11] J. Martinez-Llario and M. Gonzalez-Alcaide. Design of a Java spatial extension for relational databases. The Journal of systems and software, 84 (12):2314–2323, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121211001695.

Mesbah:2017:REJ

[MLM17] Abdelhak Mesbah, Jean-Louis Lanet, and Mohamed Mezghiche. Reverse engineering a Java Card memory management algorithm. Computers & Security, 66 (??):97–114, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167404817300093.

Mesbah:2019:REJ

[MLM19] Abdelhak Mesbah, Jean-Louis Lanet, and Mohamed Mezghiche. Reverse engineering Java Card and vulnerability exploitation: a shortcut to ROM. International Journal of Information Security, 18(1):85– 100, February 2019. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s10207-018-0401-
                                        9.

Madsen:2017:MRA

[MLT17] Magnus Madsen, Ondrej Lhoták, and Frank Tip. A model for reasoning about JavaScript promises. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA): 86:1–86:??, October 2017. CODEN ???? ISSN 2475-1421.

Mirshokraie:2012:JJA

[MM12] Shabnam Mirshokraie and Ali Mesbah. JSART: JavaScript assertion-based regression testing. Lecture Notes in Computer Science, 7387:238–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.

com/chapter/10.1007/978-
3-642-31753-8_18/.

McBurney:2016:ASC

[MM16] P. W. McBurney and C. McMillan. Automatic source code summarization of context for Java methods. IEEE Transactions on Software Engineering, 42(2):103–119, February 2016. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7181703.

Markstrum:2010:JDP

[MME^{+}10] Shane Markstrum, Daniel Marino, Matthew Esquivel, Todd Millstein, Chris Andreae, and James Noble. JavaCOP: Declarative pluggable types for Java. ACM Transactions on Programming Languages and Systems, 32(2):4:1– 4:37, January 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Martin:2014:TCR

[MME14] Marko Martin, Mira Mezini, and Sebastian Erdweg. Template constructors for reusable object initialization. ACM SIGPLAN Notices, 49(3):43–52, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mirzaei:2012:TAA

[MMP^{+}12] Nariman Mirzaei, Sam Malek, Corina S. Păsăreanu, Naeem Esfahani, and Riyadh Mahmood. Testing Android apps through symbolic execution. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mirshokraie:2015:GMT

[MMP15] S. Mirshokraie, A. Mesbah, and K. Pattabiraman. Guided mutation testing for JavaScript Web applications. IEEE Transactions on Software Engineering, 41 (5):429–444, May 2015. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=
                                        6960094.

Morgan:2018:SJW

[Mor18] Joe Morgan. Simplifying JavaScript: writing modern JavaScript with ES5, ES6, and beyond. The Pragmatic Bookshelf, Raleigh, North Carolina, 2018. ISBN 1-68050-288-3 (paperback), 1-68050-604-8. xiv + 260 pp. LCCN ????

Mastrangelo:2015:UYO

[MPM^{+}15] Luis Mastrangelo, Luca Ponzanelli, Andrea Mocci, Michele Lanza, Matthias Hauswirth, and Nathaniel Nystrom. Use at your own risk: the Java unsafe API in the wild. ACM SIG-PLAN Notices, 50(10):695– 710, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mercer:2012:CVI

[MPR12] Eric Mercer, Suzette Person, and Neha Rungta. Computing and visualizing the impact of change with Java PathFinder extensions. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Magazinius:2012:SWS

[MPS12] Jonas Magazinius, Phu H.

Phung, and David Sands. Safe wrappers and sane policies for self protecting JavaScript. Lecture Notes in Computer Science, 7127: 239–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-27937-
9_17/.

Mamouras:2017:SMS

[MRA^{+}17] Konstantinos Mamouras, Mukund Raghothaman, Rajeev Alur, Zachary G. Ives, and Sanjeev Khanna. StreamQRE: modular specification and efficient evaluation of quantitative queries over streaming data. ACM SIGPLAN Notices, 52(6): 693–708, June 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mace:2018:PTD

[MRF18] Jonathan Mace, Ryan Roelke, and Rodrigo Fonseca. Pivot tracing: Dynamic causal monitoring for distributed systems. ACM Transactions on Computer Systems, 35(4):11:1–11:??, December 2018. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3208104.

Meawad:2012:EBS

[MRMV12] Fadi Meawad, Gregor Richards, Floréal Morandat, and Jan Vitek. Eval begone!: semiautomated removal of eval from JavaScript programs. ACM SIGPLAN Notices, 47 (10):607–620, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McIlroy:2010:HJR

[MS10] Ross McIlroy and Joe Sventek. Hera-JVM: a runtime system for heterogeneous multi-core architectures. ACM SIGPLAN Notices, 45(10):205–222, October 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marinescu:2013:FSJ

[MS13] Maria-Cristina Marinescu and César Sánchez. Fusing statecharts and Java. ACM Transactions on Embedded Computing Systems, 12(1s):45:1–45:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Moller:2014:ADC

[MS14] Anders Møller and Mathias Schwarz. Automated detection of client-state manipulation vulnerabilities. ACM Transactions on Software Engineering and Methodology, 23(4): 29:1–29:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Marino:2010:DSE

[MSM^{+}10] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. DRFX: a simple and efficient memory model for concurrent programming languages. ACM SIG-PLAN Notices, 45(6):351– 362, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marino:2016:DXU

[MSM^{+}16] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. drf x: an understandable, high performance, and flexible memory model for concurrent languages. ACM Transactions on Programming Languages and Systems, 38(4):16:1–16:??, October 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mitchell:2010:FTL

[MSS10] Nick Mitchell, Edith Schonberg, and Gary Sevitsky. Four trends leading to Java runtime bloat. IEEE Software, 27(1):56–63, January/ February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marchetto:2019:CCR

[MSS19] A. Marchetto, G. Scanniello, and A. Susi. Combining code and requirements coverage with execution cost for test suite reduction. IEEE Transactions on Software Engineering, 45(4):363–390, April 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Mitropoulos:2016:HTY

[MSSK16] Dimitris Mitropoulos, Konstantinos Stroggylos, Diomidis Spinellis, and Angelos D. Keromytis. How to train your browser: Preventing XSS attacks using contextual script fingerprints. ACM Transactions on Privacy and Security (TOPS), 19(1):2:1– 2:??, August 2016. CO-DEN ???? ISSN 2471-2566 (print), 2471-2574 (electronic).

Malhotra:2013:DFT

[MT13] Ruchika Malhotra and Divya Tiwari. Development of a framework for test case prioritization using genetic algorithm. ACM SIG-SOFT Software Engineering Notes, 38(3):1–6, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Murawski:2014:GSI

[MT14] Andrzej S. Murawski and Nikos Tzevelekos. Game semantics for interface middleweight Java. ACM SIG-PLAN Notices, 49(1):517– 528, January 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). POPL ’14 conference proceedings.

Madsen:2015:SAE

[MTL15] Magnus Madsen, Frank Tip, and Ondrej Lhoták. Static analysis of eventdriven Node.js JavaScript applications. ACM SIG-PLAN Notices, 50(10):505– 519, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehrnezhad:2016:TIU

[MTSH16] Maryam Mehrnezhad, Ehsan Toreini, Siamak F. Shahandashti, and Feng Hao. TouchSignatures: Identification of user touch actions and PINs based on mobile sensor data via JavaScript. Journal of Information Security and Applications (JISA), 26 (??):23–38, February 2016. CODEN ???? ISSN 2214-2126. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2214212615000678.

Marz:2016:RPC

[MV16] Stephen Marz and Brad Vander Zanden. Reducing power consumption and latency in mobile devices using an event stream model. ACM Transactions on Embedded Computing Systems, 16(1):11:1–11:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mesbah:2012:CAB

[MvDL12] Ali Mesbah, Arie van Deursen, and Stefan Lenselink. Crawling Ajax-based Web applications through dynamic analysis of user interface state changes. ACM Transactions on the Web (TWEB), 6(1):3:1– 3:??, March 2012. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Motika:2015:LWS

[MvH15] Christian Motika and Reinhard von Hanxleden. Lightweight Synchronous Java (SJL): An approach for programming deterministic reactive systems with Java. Computing, 97(3):281–307, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mateos:2010:ANI

[MZC10a] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. An approach for non-intrusively adding malleable fork/join parallelism into ordinary JavaBean compliant applications. Computer Languages, Systems and Structures, 36 (3):288–315, October 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1477842410000035.

Mateos:2010:MJN

[MZC10b] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. m-JGRIM: a novel middleware for Gridifying Java applications into mobile Grid services. Software —Practice and Experience, 40(4):331–362, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nowicki:2018:MPI

[NBB18] Marek Nowicki, Davit Bzhalava, and Piotr Bała. Massively parallel implementation of sequence alignment with basic local alignment search tool using parallel computing in Java library. Journal of Computational Biology, 25(8):871– 881, August 2018. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

                                        tronic). URL https://www.
                                        liebertpub.com/doi/abs/
                                        10.1089/cmb.2018.0079;
                                         https://www.liebertpub.
                                        com/doi/pdf/10.1089/cmb.
                                        2018.0079.

Nasseri:2010:CMR

[NCS10] E. Nasseri, S. Counsell, and M. Shepperd. Class movement and re-location: an empirical study of Java inheritance evolution. The Journal of systems and software, 83(2):303–315, February 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nuzman:2013:JTC

[NED^{+}13] Dorit Nuzman, Revital Eres, Sergei Dyshel, Marcel Zalmanovici, and Jose Castanos. JIT technology with C/C++: Feedbackdirected dynamic recompilation for statically compiled languages. ACM Transactions on Architecture and Code Optimization, 10(4):59:1–59:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nguyen:2018:SCM

[NFN^{+}18] Khanh Nguyen, Lu Fang, Christian Navasca, Guoqing Xu, Brian Demsky, and Shan Lu. Skyway: Connecting managed heaps in distributed big data systems. ACM SIGPLAN Notices, 53(2):56–69, February 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newton:2015:ALF

[NFV15] Ryan R. Newton, Peter P.

Fogg, and Ali Varamesh. Adaptive lock-free maps: purely-functional to scalable. ACM SIGPLAN Notices, 50(9):218–229, September 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noll:2012:IDO

[NG12] Albert Noll and Thomas R. Gross. An infrastructure for dynamic optimization of parallel programs. ACM SIGPLAN Notices, 47(8): 325–326, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Noll:2013:OFD

[NG13] Albert Noll and Thomas Gross. Online feedbackdirected optimizations for parallel Java code. ACM SIGPLAN Notices, 48(10): 713–728, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Nunez:2016:PGC

[NGB16] Diogenes Nunez, Samuel Z. Guyer, and Emery D. Berger. Prioritized garbage collection: explicit GC support for software caches. ACM SIGPLAN Notices, 51 (10):695–710, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ngo:2012:BRE

[Ngo12] Terry Ngo. Book review:

Essential app engine: building high-performance Java apps with Google app engine by Adriaan de Jonge. ACM SIGSOFT Software Engineering Notes, 37(2): 37, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nilsen:2012:RTJ

[Nil12a] Kelvin Nilsen. Real-time Java in modernization of the Aegis Weapon System. ACM SIGADA Ada Letters, 32(3):63–70, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Nilsen:2012:TOU

[Nil12b]  Kelvin Nilsen.  Tuto-

rial overview: understanding dynamic memory management in safety critical Java. ACM SIGADA Ada Letters, 32(3):15–22, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Namjoshi:2010:NOP

[NK10] Manjiri A. Namjoshi and Prasad A. Kulkarni. Novel online profiling for virtual machines. ACM SIG-PLAN Notices, 45(7):133– 144, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Na:2016:JPC

[NKH16] Yeoul Na, Seon Wook Kim, and Youngsun Han. JavaScript parallelizing compiler for exploiting parallelism from data-parallel HTML5 applications. ACM Transactions on Architecture and Code Optimization, 12(4):64:1–64:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nolan:2014:XWT

[NL14] Deborah Ann Nolan and Duncan Temple Lang. XML and Web technologies for data sciences with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4614-7899-5. xxiv + 663 pp. LCCN QA76.76.H94 N65 2014. URL http:

//www.loc.gov/catdir/
enhancements/fy1503/2013954669-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy1503/2013954669-d.html;
 http://www.loc.gov/catdir/

enhancements/fy1503/2013954669-
t.html.

Nakaike:2010:LER

[NM10a] Takuya Nakaike and Maged M.

Michael. Lock elision for read-only critical sections in Java. ACM SIGPLAN Notices, 45(6):269–278, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nourie:2010:REJ

[NM10b] D. Nourie and M. Mc-Closkey. Regular expressions and the Java programming language. Web report, 2010. URL http://

java.sun.com/developer/
technicalArticles/releases/
1.4regex.

Noller:2019:CSS

[NNT^{+}19] Yannic Noller, Hoang Lam Nguyen, Minxing Tang, Timo Kehrer, and Lars Grunske. Complete shadow symbolic execution with Java PathFinder. ACM SIGSOFT Software Engineering Notes, 44(4):15–16, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://

dl.acm.org/doi/10.1145/
3364452.33644558.

Noller:2017:SSE

[NNTK17] Yannic Noller, Hoang Lam Nguyen, Minxing Tang, and Timo Kehrer. Shadow symbolic execution with Java PathFinder. ACM SIG-SOFT Software Engineering Notes, 42(4):1–5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nikolic:2012:DEA

[NS12] – Durica Nikolić and Fausto Spoto. Definite expression aliasing analysis for Java bytecode. Lecture Notes in Computer Science, 7521:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-32943-
2_6/.

Nikolic:2013:RAP

[NS13] Durica Nikolić and Fausto Spoto. Reachability analysis of program variables. ACM Transactions on Programming Languages and Systems, 35(4):14:1–14:??, December 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nicolay:2017:PAJ

[NSDD17] Jens Nicolay, Quentin Stiévenart, Wolfgang De Meuter, and Coen De Roover. Purity analysis for JavaScript through abstract interpretation. Journal of Software: Evolution and Process, 29(12): ??, December 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nguyen:2015:FCR

[NWB^{+}15] Khanh Nguyen, Kai Wang, Yingyi Bu, Lu Fang, Jianfei Hu, and Guoqing Xu. FACADE: a compiler and runtime for (almost) objectbounded big data applications. ACM SIGPLAN Notices, 50(4):675–690, April 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nguyen:2018:UCM

[NWB^{+}18] Khanh Nguyen, Kai Wang, Yingyi Bu, Lu Fang, and Guoqing Xu. Understanding and combating memory bloat in managed data-intensive systems. ACM Transactions on Software Engineering and Methodology, 26(4): 12:1–12:41, February 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Naik:2012:AT

[NYCS12] Mayur Naik, Hongseok Yang, Ghila Castelnuovo, and Mooly Sagiv. Abstractions from tests. ACM SIG-PLAN Notices, 47(1):373– 386, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Omar:2017:PSF

[OA17] Cyrus Omar and Jonathan Aldrich. Programmable semantic fragments: the design and implementation of typy. ACM SIGPLAN Notices, 52(3):81–92, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Obaidellah:2018:SUE

[OAC18] Unaizah Obaidellah, Mohammed Al Haek, and Peter C.-H. Cheng. A survey on the usage of eye-tracking in computer programming. ACM Computing Surveys, 51(1):5:1–5:??, April 2018. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Oaks:2014:JPD

[Oak14] Scott Oaks. Java Performance: The Definitive Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-5845-4 (paperback), 1-4493-6354-7 (e-book). xiv + 408 pp. LCCN QA76.73.J38.

                                        URL http://proquest.
                                        safaribooksonline.com/
                                        9781449363512.

Ocariza:2017:SCC

[OBPM17] F. S. Ocariza, K. Bajaj, K. Pattabiraman, and A. Mesbah. A study of causes and consequences of client-side JavaScript bugs. IEEE Transactions on Software Engineering, 43(2):128–144, February 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7501855.

Ortin:2014:RPI

[OCFLI14] Francisco Ortin, Patricia Conde, Daniel Fernandez-Lanvin, and Raul Izquierdo. The runtime performance of

invokedynamic: An evalu-

ation with a Java library. IEEE Software, 31(4):82– 90, July/August 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ou:2018:TUC

[OD18] Peizhao Ou and Brian Demsky. Towards understanding the costs of avoiding out-of-thin-air results. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):136:1–136:29, October 2018. URL

https://dl.acm.org/doi/
abs/10.1145/3276506.

Olivo:2015:SDA

[ODL15] Oswaldo Olivo, Isil Dillig, and Calvin Lin. Static detection of asymptotic performance bugs in collection traversals. ACM SIG-PLAN Notices, 50(6):369– 378, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogawa:2013:RJA

[OIA^{+}13] Yuki Ogawa, Masahiro Iida, Motoki Amagasaki, Morihiro Kuga, and Toshinori Sueyoshi. A reconfigurable Java accelerator with software compatibility for embedded systems. ACM SIGARCH Computer Architecture News, 41(5):71– 76, December 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Olszak:2012:RJP

[OJ12] Andrzej Olszak and Bo Nørregaard Jørgensen. Remodularizing Java programs for improved locality of feature implementations in source code. Science of Computer Programming, 77(3):131–151, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167642310001917.

Ogata:2010:SJN

[OMK^{+}10] Kazunori Ogata, Dai Mikurube, Kiyokuni Kawachiya, Scott Trent, and Tamiya Onodera. A study of Java’s non-Java memory. ACM SIGPLAN Notices, 45 (10):191–204, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odaira:2010:ERT

[OOK^{+}10] Rei Odaira, Kazunori Ogata, Kiyokuni Kawachiya, Tamiya Onodera, and Toshio Nakatani. Efficient runtime tracking of allocation sites in Java. ACM SIGPLAN Notices, 45 (7):109–120, July 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olson:2018:CLM

[OTR^{+}18] Matthew Benjamin Olson, Joseph T. Teague, Divyani Rao, Michael R. JANTZ, Kshitij A. Doshi, and Prasad A. Kulkarni. Cross-layer memory management to improve DRAM energy efficiency. ACM Transactions on Architecture and Code Optimization, 15(2):20:1–20:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ottoni:2018:HJP

[Ott18] Guilherme Ottoni. HHVM JIT: a profile-guided, regionbased compiler for PHP and hack. ACM SIG-PLAN Notices, 53(4):151– 165, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ohkawa:2013:RHO

[OUY^{+}13] Takeshi Ohkawa, Daichi Uetake, Takashi Yokota, Kanemitsu Ootsu, and Takanobu Baba. Reconfigurable and hardwired ORB engine on FPGA by Javato-HDL synthesizer for realtime application. ACM SIGARCH Computer Architecture News, 41(5):77– 82, December 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Olsson:2016:ERR

[OW16] Ronald A. Olsson and Todd Williamson. Experience reports: RJ: a Java package providing JR-like concurrent programming. Software —Practice and Experience, 46(5):685–708, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oh:2015:MWA

[OwKPM15] JinSeok Oh, Jin woo Kwon, Hyukwoo Park, and Soo-Mook Moon. Migration of Web applications with seamless execution. ACM SIGPLAN Notices, 50(7): 173–185, July 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paul:2014:RTP

[Pau14] Anand Paul. Real-time power management for embedded M2M using intelligent learning methods. ACM Transactions on Embedded Computing Systems, 13(5s):148:1–148:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pascarella:2019:CCC

[PBB19] Luca Pascarella, Magiel Bruntink, and Alberto Bacchelli. Classifying code comments in Java software systems. Empirical Software Engineering, 24(3):1499– 1537, June 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10664-019-09694-
w;   http://link.springer.
com/content/pdf/10.1007/
s10664-019-09694-w.pdf.

Ponzanelli:2019:AIC

[PBM^{+}19] L. Ponzanelli, G. Bavota, A. Mocci, R. Oliveto, M. D. Penta, S. Haiduc, B. Russo, and M. Lanza. Automatic identification and classification of software development video tutorial fragments. IEEE Transactions on Software Engineering, 45(5):464–488, May 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Parnin:2013:AUJ

[PBMH13] Chris Parnin, Christian Bird, and Emerson Murphy-Hill. Adoption and use of Java generics. Empirical Software Engineering, 18(6):1047–1089, December 2013. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/accesspage/
                                        article/10.1007/s10664-
                                        012-9236-6.

Pinto:2014:UEB

[PCL14] Gustavo Pinto, Fernando Castor, and Yu David Liu. Understanding energy behaviors of thread management constructs. ACM SIG-PLAN Notices, 49(10):345– 360, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philips:2017:DDD

[PDDD17] Laure Philips, Joeri De Koster, Wolfgang De Meuter, and Coen De Roover. Dependence-driven delimited CPS transformation for JavaScript. ACM SIG-PLAN Notices, 52(3):59– 69, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panizo:2012:EJP

[PdMG12] Laura Panizo and María del Mar Gallardo. An extension of Java PathFinder for hybrid systems. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Portillo-Dominguez:2016:ECP

[PDPM^{+}16] A. Omar Portillo-Dominguez, Philip Perry, Damien Magoni, Miao Wang, and John Murphy. Extended conference paper: TRINI: an adaptive load balancing strategy based on garbage collection for clustered Java systems. Software—Practice and Experience, 46(12): 1705–1733, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parker:2011:DPG

[PE11] Jon Parker and Joshua M.

Epstein. A distributed platform for global-scale agent-based models of disease transmission. ACM Transactions on Modeling and Computer Simulation, 22(1):2:1–2:??, December 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pradel:2012:FAP

[PG12] Michael Pradel and Thomas R. Gross. Fully automatic and precise detection of thread safety violations. ACM SIGPLAN Notices, 47(6): 521–530, June 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Pano:2018:FAL

[PGA18] Amantia Pano, Daniel Graziotin, and Pekka Abrahamsson. Factors and actors leading to the adoption of a JavaScript framework. Empirical Software Engineering, 23(6):3503–3534, December 2018. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s10664-018-9613-
                                        x.

Phan:2018:TIG

[Pha18] Quoc-Sang Phan. Test input generation using separation logic. ACM SIG-SOFT Software Engineering Notes, 43(4):55, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Park:2011:DCM

[PiLCH11] Heewan Park, Hyun il Lim, Seokwoo Choi, and Taisook Han. Detecting common modules in Java packages based on static object trace birthmark. The Computer Journal, 54(1):108– 124, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/108.
full.pdf+html.

Park:2017:PSS

[PIR17] Changhee Park, Hyeonseung Im, and Sukyoung Ryu. Precise and scalable static analysis of jQuery using a regular expression domain. ACM SIGPLAN Notices, 52(2):25–36, February 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pizlo:2017:JVM

[Piz17] Filip Pizlo. The JavaScript-Core virtual machine (invited talk). ACM SIG-PLAN Notices, 52(11):1, November 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pukall:2013:JFR

[PKC^{+}13] Mario Pukall, Christian Kästner, Walter Cazzola, Sebastian Götz, Alexander Grebhahn, Reimar Schröter, and Gunter Saake. JavAdaptor —flexible runtime updates of Java applications. Software—Practice and Experience, 43(2):153– 185, February 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Piao:2015:JJF

[PKO^{+}15] Xianglan Piao, Channoh Kim, Younghwan Oh, Huiying Li, Jincheon Kim, Hanjun Kim, and Jae W. Lee. JAWS: a JavaScript framework for adaptive CPU– GPU work sharing. ACM SIGPLAN Notices, 50(8): 251–252, August 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2019:ROC

[PKPM19] Hyukwoo Park, Sungkook Kim, Jung-Geun Park, and Soo-Mook Moon. Reusing the optimized code for JavaScript ahead-of-time compilation. ACM Transactions on Architecture and Code Optimization, 15(4): 54:1–54:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Parizek:2012:PAJ

[PL12] Pavel Parízek and OndYej Lhoták. Predicate abstraction of Java programs with collections. ACM SIG-PLAN Notices, 47(10):75– 94, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pan:2018:ASJ

[PLL^{+}18] Weifeng Pan, Bing Li, Jing Liu, Yutao Ma, and Bo Hu. Analyzing the structure of Java software systems by weighted K-core decomposition. Future Generation Computer Systems, 83(??): 431–444, June 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167739X17320940.

Park:2014:AAS

[PLR14] Changhee Park, Hongki Lee, and Sukyoung Ryu. All about the with statement in JavaScript: removing with statements in JavaScript applications. ACM SIG-PLAN Notices, 49(2):73–84, February 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Park:2018:SAJ

[PLR18] Changhee Park, Hongki Lee, and Sukyoung Ryu. Static analysis of JavaScript libraries in a scalable and precise way using loop sensitivity. Software—Practice and Experience, 48 (4):911–944, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2552.

Pawlak:2016:SLI

[PMP^{+}16] Renaud Pawlak, Martin Monperrus, Nicolas Petitprez, Carlos Noguera, and Lionel Seinturier. SPOON: a library for implementing analyses and transformations of Java source code. Software—Practice and Experience, 46(9):1155–1179, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Papadimitriou:2014:MLS

[PMTL14] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. MATLABlike scripting of Java scientific libraries in ScalaLab. Scientific Programming, 22 (3):187–199, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Phan:2012:SQI

[PMTP12] Quoc-Sang Phan, Pasquale Malacaria, Oksana Tkachuk, and Corina S. Păsăreanu.

Symbolic quantitative information flow. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Porter:2018:PJE

[Por18] Michael T. Porter. js-emass:

A flexible JavaScript implementation of the emass algorithm. Journal of Open Source Software, 3 (28):869:1, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00869.

Poslavsky:2019:REJ

[Pos19] Stanislav Poslavsky. Rings:

an efficient Java/Scala library for polynomial rings. Computer Physics Communications, 235(??):400–413, February 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465518303199.

Passerat-Palmbach:2015:TSS

[PPMH15] Jonathan Passerat-Palmbach, Claude Mazel, and David R. C. Hill. TaskLocalRandom: a statistically sound substitute to pseudorandom number generation in parallel Java tasks frameworks. Concurrency and Computation: Practice and Experience, 27(13):3383–3398, September 10, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pichon-Pharabod:2016:CSR

[PPS16] Jean Pichon-Pharabod and Peter Sewell. A concurrency semantics for relaxed atomics that permits optimisation and avoids thinair executions. ACM SIG-PLAN Notices, 51(1):622– 633, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pham-Quang:2012:JAD

[PQD12] Phuong Pham-Quang and Benoit Delinchant. Java automatic differentiation tool using virtual operator overloading. In Forth et al. [FHP^{+}12], pages 241–250. CODEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ????

                                        URL http://link.springer.
                                        com/content/pdf/10.1007/
                                        978-3-642-30023-3_22. Pro-

ceedings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Piedrahita-Quintero:2017:JGA

[PQTGS17] Pablo Piedrahita-Quintero, Carlos Trujillo, and Jorge Garcia-Sucerquia. JDiffraction: a GPGPU-accelerated JAVA library for numerical propagation of scalar wave fields. Computer Physics Communications, 214(??): 128–139, May 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465516303952.

Pironti:2010:PCJ

[PS10a] Alfredo Pironti and Riccardo Sisto. Provably correct Java implementations of Spi Calculus security protocols specifications. Computers & Security, 29(3): 302–314, May 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167404809000832.

Pitter:2010:RTJ

[PS10b] Christof Pitter and Martin Schoeberl. A real-time Java chip-multiprocessor. ACM Transactions on Embedded Computing Systems, 10(1): 9:1–9:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Palmer:2011:BJM

[PS11] Zachary Palmer and Scott F.

Smith. Backstage Java: making a difference in metaprogramming. ACM SIGPLAN Notices, 46(10): 939–958, October 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Park:2012:CB

[PS12] Chang-Seo Park and Koushik Sen. Concurrent breakpoints. ACM SIGPLAN Notices, 47(8):331–332, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Paquin:2018:AAS

[PSJ18] Maria Paquin, Elena Sherman, and Amit Jain. Assessing the adequacy of synthetic programs for learning SPF’s configurations. ACM SIGSOFT Software Engineering Notes, 43(4): 55, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pradel:2014:EAR

[PSNS14] Michael Pradel, Parker Schuh, George Necula, and Koushik Sen. Event-Break: analyzing the responsiveness of user interfaces through performanceguided test generation. ACM SIGPLAN Notices, 49 (10):33–47, October 2014.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:2015:KCF

[PSR15] Daejun Park, Andrei Stefanescu, and Grigore Rosu. KJS: a complete formal semantics of JavaScript. ACM SIGPLAN Notices, 50(6): 346–356, June 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pour:2011:MBD

[PSW11] Niusha Hakimi Pour, Paul Strooper, and Andy Wellings. A model-based development approach for the verification of real-time Java code. Concurrency and Computation: Practice and Experience, 23(13):1583–1606, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pinto:2015:LSS

[PTF^{+}15] Gustavo Pinto, Weslley Torres, Benito Fernandes, Fernando Castor, and Roberto S. M. Barros. A largescale study on the usage of Java’s concurrent programming constructs. The Journal of systems and software, 106(??):59–81, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121215000849.

Pape:2014:EJV

[PTHH14] Tobias Pape, Arian Treffer, Robert Hirschfeld, and Michael Haupt. Extending a Java Virtual Machine to Dynamic Object-oriented Languages, volume 82 of Technische Berichte des Hasso-Plattner-Instituts für Softwaresystemtechnik an der Universität Potsdam. Universitätsverlag Potsdam, Potsdam, Germany, 2014. ISBN 3-86956-266-

8. 163 pp. LCCN ???? URL http:/

                                        /d-nb.info/1046379119/
                                        04;   http://opus.kobv.
                                        de/ubp/volltexte/2013/
                                        6743/.

Papadimitriou:2011:SES

[PTML11]  Stergios  Papadimitriou,

Konstantinos Terzidis, Seferina Mavroudi, and Spiridon Likothanassis. ScalaLab: An effective Scala-based scientific programming environment for Java. Computing in Science and Engineering, 13(5):43–55, September/October 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Paletov:2018:ICA

[PTRV18] Rumen Paletov, Petar Tsankov, Veselin Raychev, and Martin Vechev. Inferring crypto API rules from code changes. ACM SIG-PLAN Notices, 53(4):450– 464, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Puffitsch:2013:SIP

[Puf13] Wolfgang Puffitsch. Special issue papers: Design and analysis of a hard realtime garbage collector for a Java chip multi-processor. Concurrency and Computation: Practice and Experience, 25(16):2269–2289, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Petrashko:2016:CGL

[PULO16] Dmitry Petrashko, Vlad Ureche, Ondrej Lhoták, and Martin Odersky. Call graphs for languages with parametric polymorphism. ACM SIGPLAN Notices, 51 (10):394–409, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powers:2017:BBG

[PVB17] Bobby Powers, John Vilk, and Emery D. Berger. Browsix: Bridging the gap between Unix and the browser. ACM SIGARCH Computer Architecture News, 45(1):253–266, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Pina:2014:RDJ

[PVH14] Luís Pina, Luís Veiga, and Michael Hicks. Rubah: DSU for Java on a stock JVM. ACM SIGPLAN Notices, 49(10):103–119, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plumbridge:2013:BPR

[PWA13] Gary Plumbridge, Jack Whitham, and Neil Audsley. Blueshell: a platform for rapid prototyping of multiprocessor NoCs and accelerators. ACM SIGARCH Computer Architecture News, 41(5):107– 117, December 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Pan:2017:GCF

[PWSG17] Y. Pan, J. White, Y. Sun, and J. Gray. Gray computing: A framework for computing with background JavaScript tasks. IEEE Transactions on Software Engineering, PP(99): 1, ???? 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
8105894.

Pan:2019:GCF

[PWSG19] Yao Pan, Jules White, Yu Sun, and Jeff Gray. Gray computing: A framework for computing with background JavaScript tasks. IEEE Transactions on Software Engineering, 45(??): 171–193, ???? 2019. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https:

//ieeexplore.ieee.org/
document/8105894/.

Pizlo:2010:SFT

[PZM^{+}10] Filip Pizlo, Lukasz Ziarek, Petr Maj, Antony L. Hosking, Ethan Blanton, and Jan Vitek. SCHISM: fragmentation-tolerant realtime garbage collection. ACM SIGPLAN Notices, 45(6):146–159, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qiu:2017:USR

[QLBS17] Dong Qiu, Bixin Li, Earl T.

Barr, and Zhendong Su. Understanding the syntactic rule usage in Java. The Journal of systems and software, 123(??):160–172, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121216302126.

Qian:2016:EFS

[QSaS^{+}16] Junjie Qian, Witawas Srisaan, Sharad Seth, Hong Jiang, Du Li, and Pan Yi. Exploiting FIFO scheduler to improve parallel garbage collection performance. ACM SIG-PLAN Notices, 51(7):109– 121, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rayns:2013:CJS

[R^{+}13] Chris Rayns et al. CICS and the JVM server developing and deploying Java applications. IBM redbooks. IBM Corporation, International Technical Support Organization, Poughkeepsie, NY, USA, 2013. ISBN 0-7384-3833-

2. ???? pp. LCCN ????

                                        URL http://proquest.
                                        tech.safaribooksonline.
                                        de/0738438332.

Rehman:2016:VMJ

[RAS16] Waqas Ur Rehman, Muhammad Sohaib Ayub, and Junaid Haroon Siddiqui. Verification of MPI Java programs using software model checking. ACM SIGPLAN Notices, 51(8):55:1–55:??, August 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rauschmayer:2014:SJD

[Rau14] Axel Rauschmayer. Speaking JavaScript: [an indepth guide for programmers]. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-6503-5. xvii + 437 pp. LCCN QA76.73.J39 R38 2014.

Rossi:2015:NPJ

[RB15] Gianfranco Rossi and Federico Bergenti. Nondeterministic programming in Java with JSetL. Fundamenta Informaticae, 140 (3–4):393–412, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Razafindralambo:2012:FFH

[RBL12] Tiana Razafindralambo, Guillaume Bouffard, and Jean-Louis Lanet. A friendly framework for hidding fault enabled virus for Java based Smartcard. Lecture Notes in Computer Science, 7371: 122–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-31540-
4_10/.

Raychev:2016:PMC

[RBV16] Veselin Raychev, Pavol Bielik, and Martin Vechev. Probabilistic model for code with decision trees. ACM SIGPLAN Notices, 51(10): 731–747, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rathee:2017:ROO

[RC17] Amit Rathee and Jitender Kumar Chhabra. Restructuring of objectoriented software through cohesion improvement using frequent usage patterns. ACM SIGSOFT Software Engineering Notes, 42(3): 1–8, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosa:2017:APV

[RCB17] Andrea Rosà, Lydia Y. Chen, and Walter Binder. Actor profiling in virtual execution environments. ACM SIGPLAN Notices, 52(3): 36–46, March 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robatmili:2014:MRL

[RCR^{+}14] Behnam Robatmili, Calin Cascaval, Mehrdad Reshadi, Madhukar N. Kedlaya, Seth Fowler, Vrajesh Bhavsar, Michael Weber, and Ben Hardekopf. MuscalietJS: rethinking layered dynamic web runtimes. ACM SIGPLAN Notices, 49 (7):77–88, July 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radoi:2015:ETS

[RD15] Cosmin Radoi and Danny Dig. Effective techniques for static race detection in Java parallel loops. ACM Transactions on Software Engineering and Methodology, 24(4):24:1–24:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ramirez-Deantes:2012:MTA

[RDCP12] D. Ramírez-Deantes, J. Correas, and G. Puebla. Modular termination analysis of Java bytecode and its application to phoneME core libraries. Lecture Notes in Computer Science, 6921: 218–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-27269-
1_13/.

Rhodes:2015:DDO

[RDF15] Dustin Rhodes, Tim Disney, and Cormac Flanagan. Dynamic detection of object capability violations through model checking. ACM SIGPLAN Notices, 50(2):103–112, February 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynders:2016:GSB

[RDP16] Bob Reynders, Dominique Devriese, and Frank Piessens. Generating safe boundary APIs between typed ED-SLs and their environments. ACM SIGPLAN Notices, 51 (3):31–34, March 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynolds:2013:MJB

[Rey13] Mark C. Reynolds. Modeling the Java Bytecode Verifier. Science of Computer Programming, 78(3):327– 342, March 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167642311000943.

Reza:2012:JS

[Rez12] Juan Rolando Reza. Java supervenience. Computer Languages, Systems and Structures, 38(1):73– 97, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842411000182.

Richard-Foy:2014:EHL

[RFBJ14] Julien Richard-Foy, Olivier Barais, and Jean-Marc Jézéquel. Efficient highlevel abstractions for Web programming. ACM SIG-PLAN Notices, 49(3):53– 60, March 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radoi:2014:TIC

[RFRS14] Cosmin Radoi, Stephen J.

Fink, Rodric Rabbah, and Manu Sridharan. Translating imperative code to MapReduce. ACM SIG-PLAN Notices, 49(10):909– 927, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roemer:2018:HCU

[RGB18] Jake Roemer, Kaan Genç, and Michael D. Bond. Highcoverage, unbounded sound predictive race detection. ACM SIGPLAN Notices, 53(4):374–389, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richards:2011:ACJ

[RGEV11] Gregor Richards, Andreas Gal, Brendan Eich, and Jan Vitek. Automated construction of JavaScript benchmarks. ACM SIG-PLAN Notices, 46(10):677– 694, October 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Ricci:2013:ETP

[RGM13] Nathan P. Ricci, Samuel Z. Guyer, and J. Eliot B. Moss. Elephant Tracks: portable production of complete and precise GC traces. ACM SIGPLAN Notices, 48(11): 109–118, November 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’13 conference proceedings.

Richards:2013:FAC

[RHN^{+}13] Gregor Richards, Christian Hammer, Francesco Zappa Nardelli, Suresh Jagannathan, and Jan Vitek. Flexible access control for JavaScript. ACM SIG-PLAN Notices, 48(10):305– 322, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Radoi:2015:WAR

[RHSD15] Cosmin Radoi, Stephan Herhut, Jaswanth Sreeram, and Danny Dig. Are web applications ready for parallelism? ACM SIGPLAN Notices, 50(8):289–290, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ravn:2013:EIS

[RHT13] Anders P. Ravn and M. Teresa Higuera-Toledano. Editorials: Introduction to the special issue on Java technologies for real-time and embedded systems: JTRES2011. Concurrency and Computation: Practice and Experience, 25(16): 2225–2226, November 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Richardson:2014:BEL

[Ric14] Matt Richardson. BeagleBone für Einsteiger: [Linux-basierte Elektronik-Projekte mit Python und JavaScript]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2014. ISBN 3-95561-409-3. xii + 134 pp. LCCN ????

Rimlinger:2012:TGS

[Rim12] Frank Rimlinger. Test generation via symbolic simulation. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roohitavaf:2019:AAF

[RK19] Mohammad Roohitavaf and Sandeep Kulkarni. Automatic addition of faulttolerance in presence of unchangeable environment actions. Future Internet, 11 (7):144, July 04, 2019. CO-DEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/7/144.

Raghothaman:2018:UGP

[RKHN18]  Mukund  Raghothaman,

Sulekha Kulkarni, Kihong Heo, and Mayur Naik. Userguided program reasoning using Bayesian inference. ACM SIGPLAN Notices, 53(4):722–735, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodchenko:2018:TIE

[RKN^{+}18] Andrey Rodchenko, Christos Kotselidis, Andy Nisbet, Antoniu Pop, and Mikel Luján. Type information elimination from objects on architectures with tagged pointers support. IEEE Transactions on Computers, 67(1):130– 143, January 2018. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/
document/7962268/.

Richards:2010:ADB

[RLBV10] Gregor Richards, Sylvain Lebresne, Brian Burg, and Jan Vitek. An analysis of the dynamic behavior of JavaScript programs. ACM SIGPLAN Notices, 45(6):1– 12, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodeghero:2015:ETS

[RLMM15] P. Rodeghero, C. Liu, P. W.

McBurney, and C. McMillan. An eye-tracking study of Java programmers and application to source code summarization. IEEE Transactions on Software Engineering, 41(11):1038– 1054, November 2015. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7118751.

Rompf:2012:LMS

[RO12] Tiark Rompf and Martin Odersky. Lightweight modular staging: a pragmatic approach to runtime code generation and compiled DSLs. Communications of the ACM, 55(6):121– 130, June 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ryu:2019:TAB

[RPP19] S. Ryu, J. Park, and J. Park. Toward analysis and bug finding in JavaScript Web applications in the wild. IEEE Software, 36(3):74–82, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rathje:2014:FMC

[RR14] William Rathje and Brad Richards. A framework for model checking UDP network programs with Java Pathfinder. ACM SIGADA Ada Letters, 34(3):81–86, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosa:2017:ARC

[RRB17] Andrea Rosà, Eduardo Rosales, and Walter Binder. Accurate reification of complete supertype information for dynamic analysis on the JVM. ACM SIGPLAN Notices, 52(12):104–116, December 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosa:2019:AOT

[RRB19] Andrea Rosà, Eduardo Rosales, and Walter Binder. Analysis and optimization of task granularity on the Java Virtual Machine. ACM Transactions on Programming Languages and Systems, 41(3):19:1– 19:??, July 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3338497.

Ravn:2012:SCJ

[RS12] Anders P. Ravn and Martin Schoeberl. Safety-critical Java with cyclic executives on chip-multiprocessors. Concurrency and Computation: Practice and Experience, 24(8):772–788, ???? 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rompf:2014:SPJ

[RSB^{+}14] Tiark Rompf, Arvind K.

Sujeeth, Kevin J. Brown, HyoukJoong Lee, Hassan Chafi, and Kunle Olukotun. Surgical precision JIT compilers. ACM SIG-PLAN Notices, 49(6):41– 52, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rastogi:2015:SEG

[RSF^{+}15] Aseem Rastogi, Nikhil Swamy, Cédric Fournet, Gavin Bierman, and Panagiotis Vekris. Safe & efficient gradual typing for TypeScript. ACM SIG-PLAN Notices, 50(1):167– 180, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reichenbach:2012:PPD

[RSI12]  Christoph  Reichenbach,

Yannis Smaragdakis, and Neil Immerman. PQL: a purely-declarative Java extension for parallel programming. Lecture Notes in Computer Science, 7313: 53–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-31057-
7_4/.

Reardon:2014:SSB

[RT14] Susan Reardon and Brendan Tangney. Smartphones, studio-based learning, and scaffolding: Helping novices learn to program. ACM Transactions on Computing Education, 14(4):23:1– 23:??, December 2014. CO-DEN ???? ISSN 1946-6226.

Ramos:2013:DSJ

[RTE^{+}13] Sabela Ramos, Guillermo L. Taboada, Roberto R. Expósito, Juan Touriño, and Ramón Doallo. Design of scalable Java communication middleware for multi-core systems. The Computer Journal, 56(2):214–228, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/214.
full.pdf+html.

Ramos:2015:NCS

[RTET15] Sabela Ramos, Guillermo L.

Taboada, Roberto R. Expósito, and Juan Touriño. Nonblocking collectives for scalable Java communications. Concurrency and Computation: Practice and Experience, 27(5):1169–1187, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rubin:2014:HCW

[Rub14] Norm Rubin. Heterogeneous computing: what does it mean for compiler research? ACM SIG-PLAN Notices, 49(8):315– 316, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowe:2014:STA

[RvB14] Reuben N. S. Rowe and S. J. van Bakel. Semantic types and approximation for Featherweight Java. Theoretical Computer Science, 517(??):34–74, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0304397513006415.

Raychev:2015:PPP

[RVK15] Veselin Raychev, Martin Vechev, and Andreas Krause. Predicting program properties from “Big Code”. ACM SIGPLAN Notices, 50 (1):111–124, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raychev:2019:PPP

[RVK19] Veselin Raychev, Martin Vechev, and Andreas Krause. Predicting program properties from ‘big code’. Communications of the ACM, 62(3):99–107, March 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        cacm.acm.org/magazines/
                                        2019/3/234932/fulltext.

Ricci:2011:SAO

[RVP11] Alessandro Ricci, Mirko Viroli, and Giulio Piancastelli. simpA: an agent-oriented approach for programming concurrent applications on top of Java. Science of Computer Programming, 76 (1):37–62, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ramos:2018:APS

[RVT18] M. Ramos, M. T. Valente, and R. Terra. AngularJS performance: A survey study. IEEE Software, 35(2):72–79, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rudafshani:2017:LDD

[RW17] Masoomeh Rudafshani and Paul A. S. Ward. LeakSpot: detection and diagnosis of memory leaks in JavaScript applications. Software— Practice and Experience, 47 (1):97–123, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramamohanarao:2017:SSM

[RXK^{+}17] Kotagiri Ramamohanarao, Hairuo Xie, Lars Kulik, Shanika Karunasekera, Egemen Tanin, Rui Zhang, and Eman Bin Khunayn. SMARTS: Scalable microscopic adaptive road traffic simulator. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):26:1–26:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ryu:2016:JFB

[Ryu16] Sukyoung Ryu. Journey to find bugs in JavaScript web applications in the wild. ACM SIGPLAN Notices, 51(9):2, September 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spadini:2019:MOT

[SABB19] Davide Spadini, Maurício Aniche, Magiel Bruntink, and Alberto Bacchelli. Mock objects for testing Java systems. Empirical Software Engineering, 24 (3):1461–1498, June 2019. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

http://link.springer.
com/article/10.1007/s10664-
018-9663-0;   http://
link.springer.com/content/
pdf/10.1007/s10664-018-
9663-0.pdf.

Serbanescu:2016:DPO

[SAdB^{+}16] V. Serbanescu, K. Azadbakht, F. de Boer, C. Nagarajagowda, and B. Nobakht. A design pattern for optimizations in data intensive applications using ABS and JAVA 8. Concurrency and Computation: Practice and Experience, 28(2):374–385, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samuelson:2012:LSO

[Sam12] Pamela Samuelson. Legally speaking: Oracle v. Google: are APIs copyrightable? Communications of the ACM, 55(11):25–27, November 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spoto:2019:SII

[SBE^{+}19] Fausto Spoto, Elisa Burato, Michael D. Ernst, Pietro Ferrara, Alberto Lovato, Damiano Macedonio, and Ciprian Spiridon. Static identification of injection attacks in Java. ACM Transactions on Programming Languages and Systems, 41(3):18:1– 18:??, July 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3332371.

Sartor:2010:ZRD

[SBF^{+}10] Jennifer B. Sartor, Stephen M.

Blackburn, Daniel Frampton, Martin Hirzel, and Kathryn S. McKinley. Zrays: divide arrays and conquer speed and flexibility. ACM SIGPLAN Notices, 45(6):471–482, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smaragdakis:2013:SBP

[SBK13] Yannis Smaragdakis, George Balatsouras, and George Kastrinis. Set-based preprocessing for points-to analysis. ACM SIGPLAN Notices, 48(10):253–270, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Shahriyar:2014:FCG

[SBM14] Rifat Shahriyar, Stephen M. Blackburn, and Kathryn S. McKinley. Fast conservative garbage collection. ACM SIGPLAN Notices, 49 (10):121–139, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scherr:2016:AFC

[SC16] Maximilian Scherr and Shigeru Chiba. Almost first-class language embedding: taming staged embedded DSLs. ACM SIG-PLAN Notices, 51(3):21– 30, March 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:2010:ERA

[Sch10a] Richard B. Schmidt. Experience report: Ada & Java integration in the FAA’s ERAM SWIM program. ACM SIGADA Ada Letters, 30(3):33–34, December 2010. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Schultz:2010:WAJ

[Sch10b] Christopher Schultz. Web applications with Java/JSP. Linux Journal, 2010(197): 4:1–4:??, September 2010. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Schmeisser:2013:MOE

[Sch13] Michael Schmeißer. Metriken und optimale Einsatzszenarien für Garbage Collectoren der Java HotSpot Virtual Machine. (German) [Metrics and best use scenarios for garbage collectors of the Java HotSpot Virtual Machine]. Masterarbeit, Hochschule für Technik, Wirtschaft und Kultur, Leipzig, Germany, 2013. iii + 103 pp.

Schildt:2014:JCRb

[Sch14] Herbert Schildt, editor.

Java: The Complete Reference. McGraw-Hill, New York, NY, USA, ninth edition, 2014. ISBN 0-07-180855-8 (paperback), 0-07-180925-2, 0-07-180856-6. xxxiv + 1274 pp. LCCN QA76.73.J38 S332 2014eb.

Sluanschi:2016:AAD

[SD16a] Emil I. Sluşanschi and Vlad Dumitrel. ADiJaC — automatic differentiation of Java classfiles. ACM Transactions on Mathematical Software, 43(2):9:1–9:33, September 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=2904901.

Sousa:2016:CHL

[SD16b] Marcelo Sousa and Isil Dillig. Cartesian Hoare logic for verifying k-safety properties. ACM SIGPLAN Notices, 51(6):57–69, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridharan:2012:CTP

[SDC^{+}12] Manu Sridharan, Julian Dolby, Satish Chandra, Max Schäfer, and Frank Tip. Correlation tracking for points-to analysis of JavaScript. Lecture Notes in Computer Science, 7313: 435–458, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-31057-
7_20/.

Schoeberl:2017:SCJ

[SDH^{+}17] Martin Schoeberl, Andreas Engelbredt Dalsgaard, René Rydhof Hansen, Stephan E. Korsholm, Anders P. Ravn, Juan Ricardo Rios Rivas, Tórur Biskopstø Strøm, Hans Søndergaard, Andy Wellings, and Shuai Zhao. Safety-critical Java for embedded systems. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shah:2012:AMJ

[SDM12] Syed Muhammad Ali Shah, Jens Dietrich, and Catherine McCartin. On the automated modularisation of Java programs using service locators. Lecture Notes in Computer Science, 7306: 132–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-30564-
1_9/.

Sartor:2012:EMT

[SE12] Jennfer B. Sartor and Lieven Eeckhout. Exploring multi-threaded Java application performance on multicore hardware. ACM SIG-PLAN Notices, 47(10):281– 296, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stolee:2014:SSS

[SED14] Kathryn T. Stolee, Sebastian Elbaum, and Daniel Dobos. Solving the search for source code. ACM Transactions on Software Engineering and Methodology, 23(3):26:1–26:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Staples:2019:SAB

[SEK^{+}19] J. Staples, C. Endicott, L. Krause, P. Pal, P. Samouelian, R. Schantz, and A. Wellman. A semi-autonomic bytecode repair framework. IEEE Software, 36(2):97– 102, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Simao:2019:GWS

[SEPV19] J. Simão, S. Esteves, André Pires, and L. Veiga. GC-Wise: a self-adaptive approach for memoryperformance efficiency in Java VMs. Future Generation Computer Systems, 100(??):674–688, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304898.

Serrano:2018:JAC

[Ser18] Manuel Serrano. JavaScript AOT compilation. ACM SIGPLAN Notices, 53(8): 50–63, October 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3393673.3276950.

Seth:2013:UJV

[Set13] Sachin Seth. Understanding Java Virtual Machine. Alpha Science International, Oxford, UK, 2013. ISBN 1-84265-815-8. 318 pp. LCCN QA76.73.J38 S437 2013.

Severance:2012:DJO

[Sev12a] Charles Severance. Discovering JavaScript object notation. Computer, 45(4): 6–8, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:JDL

[Sev12b] Charles Severance. JavaScript:

Designing a language in 10 days. Computer, 45(2):7– 8, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sewell:2012:TJ

[Sew12] Peter Sewell. Tales from the jungle. ACM SIG-PLAN Notices, 47(9):271– 272, September 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swamy:2014:GTE

[SFR^{+}14] Nikhil Swamy, Cedric Fournet, Aseem Rastogi, Karthikeyan Bhargavan, Juan Chen, Pierre-Yves Strub, and Gavin Bierman. Gradual typing embedded securely in JavaScript. ACM SIG-PLAN Notices, 49(1):425– 437, January 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). POPL ’14 conference proceedings.

Sherman:2015:DTB

[SGD15] Elena Sherman, Brady J. Garvin, and Matthew B. Dwyer. Deciding typebased partial-order constraints for path-sensitive analysis. ACM Transactions on Software Engineering and Methodology, 24 (3):15:1–15:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Subercaze:2017:UPT

[SGG^{+}17] Julien Subercaze, Christophe Gravier, Syed Gillani, Abderrahmen Kammoun, and Frédérique Laforest. Upsortable: programming topk queries over data streams. Proceedings of the VLDB Endowment, 10(12):1873– 1876, August 2017. CODEN ???? ISSN 2150-8097.

Simao:2012:CER

[SGV12] José Simão, Tiago Garrochinho, and Luís Veiga. A checkpointing-enabled and resource-aware Java Virtual Machine for efficient and robust e-Science applications in grid environments. Concurrency and Computation: Practice and Experience, 24(13):1421–1442, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stuchlik:2012:SVD

[SH12] Andreas Stuchlik and Stefan Hanenberg. Static vs. dynamic type systems: an empirical study about the relationship between type casts and development time. ACM SIGPLAN Notices, 47(2):97–106, February 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steimann:2016:CRA

[SHU16] Friedrich Steimann, Jörg Hagemann, and Bastian Ulke. Computing repair alternatives for malformed programs using constraint attribute grammars. ACM SIGPLAN Notices, 51(10): 711–730, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siebert:2010:CPR

[Sie10]  Fridtjof Siebert.  Con-

current, parallel, real-time garbage-collection. ACM SIGPLAN Notices, 45(8): 11–20, August 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siek:2017:CPT

[Sie17] Jeremy Siek. Challenges and progress toward efficient gradual typing (invited talk). ACM SIG-PLAN Notices, 52(11):2, November 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singer:2010:EGC

[SJBL10] Jeremy Singer, Richard E. Jones, Gavin Brown, and Mikel Luján. The economics of garbage collection. ACM SIGPLAN Notices, 45(8):103–112, August 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smans:2010:AVJ

[SJPS10] Jan Smans, Bart Jacobs, Frank Piessens, and Wolfram Schulte. Automatic verification of Java programs with dynamic frames. Formal Aspects of Computing, 22(3–4):423– 457, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s00165-010-0148-
1.

Shan:2012:OAC

[SK12] Zhe Shan and Akhil Kumar. Optimal adapter creation for process composition in synchronous vs. asynchronous communication. ACM Transactions on Management Information Systems (TMIS), 3(2): 8:1–8:??, July 2012. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Salkeld:2013:IDO

[SK13] Robin Salkeld and Gregor Kiczales. Interacting with dead objects. ACM SIG-PLAN Notices, 48(10):203– 216, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Singer:2011:GCA

[SKBL11] Jeremy Singer, George Kovoor, Gavin Brown, and Mikel Luján. Garbage collection auto-tuning for Java MapReduce on multi-cores. ACM SIGPLAN Notices, 46(11):109–118, November 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’11 conference proceedings.

Schoeberl:2011:HAL

[SKKR11] Martin Schoeberl, Stephan Korsholm, Tomas Kalibera, and Anders P. Ravn. A hardware abstraction layer in Java. ACM Transactions on Embedded Computing Systems, 10(4):42:1– 42:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sondergaard:2017:CTD

[SKR17] Hans Søndergaard, Stephan E. Korsholm, and Anders P. Ravn. Conformance test development with the Java modeling language. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stilkerich:2017:PGU

[SLE^{+}17] Isabella Stilkerich, Clemens Lang, Christoph Erhardt, Christian Bay, and Michael Stilkerich. The perfect getaway: Using escape analysis in embedded real-time systems. ACM Transactions on Embedded Computing Systems, 16(4):99:1– 99:30, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Stilkerich:2015:PGA

[SLES15] Isabella Stilkerich, Clemens Lang, Christoph Erhardt, and Michael Stilkerich. A practical getaway: Applications of escape analysis in embedded real-time systems. ACM SIGPLAN Notices, 50(5):4:1–4:??, May 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steele:2014:FSP

[SLF14] Guy L. Steele, Jr., Doug Lea, and Christine H. Flood. Fast splittable pseudorandom number generators. ACM SIGPLAN Notices, 49(10):453–472, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snellenburg:2012:GJB

[SLS^{+}12] Joris J. Snellenburg, Sergey Laptenok, Ralf Seger, Katharine M. Mullen, and Ivo H. M. van Stokkum. Glotaran: a Java-based graphical user interface for the R package TIMP. Journal of Statistical Software, 49(3): ??, June 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i03.

Shafiei:2012:MCL

[SM12] Nastaran Shafiei and Peter Mehlitz. Modeling class loaders in Java PathFinder version 7. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2012:EPS

[SMN^{+}12] Abhayendra Singh, Daniel Marino, Satish Narayanasamy, Todd Millstein, and Madan Musuvathi. Efficient processor support for DRFx, a memory model with exceptions. ACM SIGPLAN Notices, 47(4):53–66, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Santos:2018:JJV

[SMN^{+}18] José Fragoso Santos, Petar Maksimović, Daiva Naudžiūnienė, Thomas Wood, and Philippa Gardner. JaVerT: JavaScript verification toolchain. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):50:1– 50:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Spoto:2010:TAJ

[SMP10] Fausto Spoto, Fred Mesnard, and Étienne Payet. A termination analyzer for Java bytecode based on path-length. ACM Transactions on Programming Languages and Systems, 32 (3):8:1–8:70, March 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Storey:2019:SDP

[SMP19] Kyle Storey, Eric Mercer, and Pavel Parizek. A sound dynamic partial order reduction engine for Java Pathfinder. ACM SIG-SOFT Software Engineering Notes, 44(4):15, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://

dl.acm.org/doi/10.1145/
3364452.3364457.

Sewe:2012:NSI

[SMS^{+}12] Andreas Sewe, Mira Mezini, Aibek Sarimbekov, Danilo Ansaloni, Walter Binder, Nathan Ricci, and Samuel Z.

Guyer.  New Scala()
instanceof Java: a com-

parison of the memory behaviour of Java and Scala programs. ACM SIG-PLAN Notices, 47(11):97– 108, November 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’12 conference proceedings.

Sewe:2011:CCS

[SMSB11] Andreas Sewe, Mira Mezini, Aibek Sarimbekov, and Walter Binder. Da capo con Scala: design and analysis of a Scala benchmark suite for the Java Virtual Machine. ACM SIG-PLAN Notices, 46(10):657– 676, October 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Stein:2019:SAD

[SNCM19] Benno Stein, Benjamin Barslev Nielsen, Bor-Yuh Evan Chang, and Anders Møller. Static analysis with demanddriven value refinement. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 140:1–140:29, October 2019.

                                        URL  https://dl.acm.
                                        org/doi/abs/10.1145/3360566.

Stork:2014:APB

[SNS^{+}14] Sven Stork, Karl Naden, Joshua Sunshine, Manual Mohr, Alcides Fonseca, Paulo Marques, and Jonathan Aldrich. Æminium: a permission based concurrentby-default programming language approach. ACM SIGPLAN Notices, 49(6): 26, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schoeberl:2010:NRT

[SP10a] Martin Schoeberl and Wolfgang Puffitsch. Nonblocking real-time garbage collection. ACM Transactions on Embedded Computing Systems, 10(1):6:1– 6:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Spoto:2010:MSL

[SP10b] Fausto Spoto and Étienne Payet. Magic-sets for localised analysis of Java bytecode. Higher-Order and Symbolic Computation, 23 (1):29–86, March 2010. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1388-3690&volume=
23&issue=1&spage=29.

Serrano:2016:GH

[SP16] Manuel Serrano and Vincent Prunet. A glimpse of Hopjs. ACM SIGPLAN Notices, 51(9):180–192, September 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steimann:2010:TMI

[SPAK10] Friedrich Steimann, Thomas Pawlitzki, Sven Apel, and Christian Kästner. Types and modularity for implicit invocation with implicit announcement. ACM Transactions on Software Engineering and Methodology, 20(1):1:1–1:??, June 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Selakovic:2018:TGH

[SPKT18] Marija Selakovic, Michael Pradel, Rezwana Karim, and Frank Tip. Test generation for higher-order functions in dynamic languages. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):161:1–161:27, October 2018. URL

https://dl.acm.org/doi/
abs/10.1145/3276531.

Spoto:2016:JSA

[Spo16] Fausto Spoto. The Julia static analyzer for Java. In Static Analysis, pages 39– 57. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

2016. URL http://link.
springer.com/chapter/
10.1007/978-3-662-53413-
7_3.

Spring:2010:RAI

[SPP^{+}10] Jesper Honig Spring, Filip Pizlo, Jean Privat, Rachid Guerraoui, and Jan Vitek. Reflexes: Abstractions for integrating highly responsive tasks into Java applications. ACM Transactions on Embedded Computing Systems, 10(1):4:1– 4:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schoeberl:2010:WCE

[SPPH10] Martin Schoeberl, Wolfgang Puffitsch, Rasmus Ulslev Pedersen, and Benedikt Huber. Worst-case execution time analysis for a Java processor. Software—Practice and Experience, 40(6):507– 542, May 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Strom:2017:HLR

[SPS17] Tórur Biskopstø Strøm, Wolfgang Puffitsch, and Martin Schoeberl. Hardware locks for a real-time Java chip multiprocessor. Concurrency and Computation: Practice and Experience, 29(6):??, March 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stefanescu:2016:SBP

[SPY^{+}16] Andrei Stefanescu, Daejun Park, Shijiao Yuwen, Yilong Li, and Grigore Rosu. Semantics-based program verifiers for all languages. ACM SIGPLAN Notices, 51(10):74–91, October 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Samak:2014:MTS

[SR14a] Malavika Samak and Murali Krishna Ramanathan. Multithreaded test synthesis for deadlock detection. ACM SIGPLAN Notices, 49(10):473–489, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Samak:2014:TDD

[SR14b] Malavika Samak and Murali Krishna Ramanathan. Trace driven dynamic deadlock detection and reproduction. ACM SIGPLAN Notices, 49(8):29–42, August 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sun:2017:AJP

[SR17] Kwangwon Sun and Sukyoung Ryu. Analysis of JavaScript programs: Challenges and research trends. ACM Computing Surveys, 50(4):59:1–59:??, November 2017. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Sawant:2018:RDC

[SRB18] Anand Ashok Sawant, Romain Robbes, and Alberto Bacchelli. On the reaction to deprecation of clients of 4 + 1 popular Java APIs and the JDK. Empirical Software Engineering, 23(4):2158–2197, August 2018. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/
10.1007/s10664-017-9554-
9;   http://link.springer.
com/content/pdf/10.1007/
s10664-017-9554-9.pdf.

Samak:2015:SRT

[SRJ15] Malavika Samak, Murali Krishna Ramanathan, and Suresh Jagannathan. Synthesizing racy tests. ACM SIGPLAN Notices, 50(6):175–185, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scanniello:2017:FFC

[SRTR17] Giuseppe Scanniello, Michele Risi, Porfirio Tramontana, and Simone Romano. Fixing faults in C and Java source code: Abbreviated vs. full-word identifier names. ACM Transactions on Software Engineering and Methodology, 26 (2):6:1–6:??, October 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sutherland:2010:CTC

[SS10] Dean F. Sutherland and William L. Scherlis. Composable thread coloring.

ACM SIGPLAN Notices, 45 (5):233–244, May 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scheben:2012:VIF

[SS12] Christoph Scheben and Peter H. Schmitt. Verification of information flow properties of Java programs without approximations. Lecture Notes in Computer Science, 7421: 232–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/
                                        10.1007/978-3-642-31762-
                                        0_15/.

Stefik:2013:EIP

[SS13] Andreas Stefik and Susanna Siebert. An empirical investigation into programming language syntax. ACM Transactions on Computing Education, 13(4):19:1– 19:??, November 2013. CO-DEN ???? ISSN 1946-6226.

Sor:2014:MLD

[SS14] Vladimir Sor and Satish Narayana Srirama. Memory leak detection in Java: Taxonomy and classification of approaches. The Journal of systems and software, 96(??):139–151, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121214001423.

Surendran:2016:APP

[SS16] Rishi Surendran and Vivek Sarkar. Automatic parallelization of pure method calls via conditional future synthesis. ACM SIGPLAN Notices, 51(10):20–38, October 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sudarsan:2019:BDK

[SS19] V. Sudarsan and R. Sugumar. Building a distributed K-means model for Weka using remote method invocation (RMI) feature of Java. Concurrency and Computation: Practice and Experience, 31(14):e5313:1– e5313:??, July 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stark:2001:JJV

[SSB01] Robert F. Stärk, Joachim Schmid, and Egon Börger. Java and the Java Virtual Machine: definition, verification, validation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42088-6. x + 381 pp. LCCN QA76.73.J38 S785 2001. US$49.95. Includes CD-ROM with the entire text of the book and numerous examples and exercises.

Sarimbekov:2014:JCS

[SSB^{+}14a] Aibek Sarimbekov, Andreas Sewe, Walter Binder, Philippe Moret, and Mira Mezini. JP2: Call-site aware calling context profiling for the Java Virtual Machine. Science of Computer Programming, 79(??):146– 157, January 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642311002036.

Stark:2014:JJV

[SSB14b] Robert F. Stärk, Joachim Schmid, and Egon Börger. Java and the Java Virtual Machine: Definition, Verification, Validation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-63997-

6. x + 381 pp. LCCN QA76.76.C65. URL http:

//deposit.d-nb.de/cgi-
bin/dokserv?id=4745749\
%26prov=M\%26dok\_var=
1\%26dok\_ext=htm. Soft-

cover reprint of [SSB01].

Su:2014:CEM

[SSG^{+}14] Xueyuan Su, Garret Swart, Brian Goetz, Brian Oliver, and Paul Sandoz. Changing engines in midstream: a Java stream computational model for big data processing. Proceedings of the VLDB Endowment, 7(13): 1343–1354, August 2014. CODEN ???? ISSN 2150-8097.

Srikanth:2017:CVU

[SSH17] Akhilesh Srikanth, Burak Sahin, and William R. Harris. Complexity verification using guided theorem enumeration. ACM SIG-PLAN Notices, 52(1):639– 652, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Singh:2013:TGC

[SSK13] Pavitdeep Singh, Satwinder Singh, and Jatinder Kaur. Tool for generating code metrics for C# source code using abstract syntax tree technique. ACM SIG-SOFT Software Engineering Notes, 38(5):1–6, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saini:2018:CNC

[SSL18] Vaibhav Saini, Hitesh Sajnani, and Cristina Lopes. Cloned and non-cloned Java methods: a comparative study. Empirical Software Engineering, 23(4):2232– 2278, August 2018. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/
                                        10.1007/s10664-017-9572-
                                        7.

Sciampacone:2010:EMS

[SSMGD10] R. A. Sciampacone, V. Sundaresan, D. Maier, and T. Gray-Donald. Exploitation of multicore systems in a Java virtual machine. IBM Journal of Research and Development, 54(5): 1:1–1:11, ???? 2010. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Stone:2015:WMT

[ST15] James Stone and John Towse. A working memory test battery: Java-based collection of seven working memory tasks. Journal of Open Research Software, 3(1):e5–??, June 05, 2015. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/
                                        10.5334/jors.br/.

Stark:2010:BIA

[Sta10] Jonathan Stark. Building iPhone apps with HTML, CSS, and JavaScript: Making App Store apps without Objective-C or Cocoa. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 1-4493-8023-9, 0-596-80578-0. xv + 166 pp. LCCN ????

Sayed:2018:ITI

[STA18] Bassam Sayed, Issa Traoré, and Amany Abdelhalim. If-transpiler: Inlining of hybrid flow-sensitive security monitor for JavaScript. Computers & Security, 75 (??):92–117, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0167404818300397.

Santos:2013:DDS

[STCG13] Ivo Santos, Marcel Tilly, Badrish Chandramouli, and Jonathan Goldstein. DiAl: distributed streaming analytics anywhere, anytime. Proceedings of the VLDB Endowment, 6(12):1386– 1389, August 2013. CODEN ???? ISSN 2150-8097.

Stefanov:2010:JP

[Ste10] Stoyan Stefanov. JavaScript patterns. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2010. ISBN 1-4493-9911-8 (paperback), 0-596-80675-2 (paperback). xvi + 216 pp. LCCN QA76.73.J39 S734 2010.

Samak:2016:DSF

[STR16] Malavika Samak, Omer Tripp, and Murali Krishna Ramanathan. Directed synthesis of failing concurrent executions. ACM SIG-PLAN Notices, 51(10):430– 446, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sun:2013:BJW

[STS^{+}13] Mengtao Sun, Gang Tan, Joseph Siefers, Bin Zeng, and Greg Morrisett. Bringing Java’s wild native world under control. ACM Transactions on Information and System Security, 16(3):9:1– 9:??, November 2013. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Schafer:2012:CAN

[STST12] M. Schäfer, A. Thies, F. Steimann, and F. Tip. A comprehensive approach to naming and accessibility in refactoring Java programs. IEEE Transactions on Software Engineering, 38(6):1233–1257, November/December 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
6152131.

Su:2014:RVP

[STY^{+}14] Tzu-Hsiang Su, Hsiang-Jen Tsai, Keng-Hao Yang, Po-Chun Chang, Tien-Fu Chen, and Yi-Ting Zhao. Reconfigurable vertical profiling framework for the Android runtime system. ACM Transactions on Embedded Computing Systems, 13(2s):59:1–59:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Subramaniam:2011:PCJ

[Sub11] Venkat Subramaniam. Programming concurrency on the JVM: mastering synchronization, STM, and actors. The pragmatic programmers. Pragmatic Bookshelf, Dallas, TX, 2011. ISBN 1-934356-76-X. xvii + 270 pp. LCCN QA76.73.J38 S8467 2011.

Sun:2018:RAR

[Sun18]  Yu-Shan Sun.  Reason-

ing about reference behavior with RESOLVE. ACM SIGSOFT Software Engineering Notes, 43(3):18– 19, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Steindorfer:2015:CSM

[SV15a] Michael J. Steindorfer and Jurgen J. Vinju. Code specialization for memory efficient hash tries (short paper). ACM SIGPLAN Notices, 50(3):11–14, March 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steindorfer:2015:OHA

[SV15b] Michael J. Steindorfer and Jurgen J. Vinju. Optimizing hash-array mapped tries for fast and lean immutable JVM collections. ACM SIG-PLAN Notices, 50(10):783– 800, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steindorfer:2017:TSP

[SV17] Michael J. Steindorfer and Jurgen J. Vinju. Towards a software product line of trie-based collections. ACM SIGPLAN Notices, 52(3): 168–172, March 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steindorfer:2018:MOA

[SV18] Michael J. Steindorfer and Jurgen J. Vinju. Tomany or to-one? Allin-one! Efficient purely functional multi-maps with type-heterogeneous hashtries. ACM SIGPLAN Notices, 53(4):283–295, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silva:2017:ICL

[SVB^{+}17] Leonardo Humberto Silva, Marco Tulio Valente, Alexandre Bergel, Nicolas Anquetil, and Anne Etien. Identifying classes in legacy JavaScript code. Journal of Software: Evolution and Process, 29(8):??, August 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sverdlove:2014:JVL

[Sve14] Harry Sverdlove. The Java vulnerability landscape. Network Security, 2014(4): 9–14, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1353485814700418.

Siek:2012:FDT

[SW12] Konrad Siek and Paweł T.

Wojciechowski. A formal design of a tool for static analysis of upper bounds on object calls in Java. Lecture Notes in Computer Science, 7437: 192–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-32469-
7_13/.

Stancu:2015:SEH

[SWB^{+}15] Codrut Stancu, Christian Wimmer, Stefan Brunthaler, Per Larsen, and Michael Franz. Safe and efficient hybrid memory management for Java. ACM SIGPLAN Notices, 50(11): 81–92, November 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szweda:2012:ANB

[SWF12] Łukasz Szweda, Daniel Wilusz, and Jakub Flotyński. Application of NXT based robots for teaching Javabased concurrency. Lecture Notes in Computer Science, 7516:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/
10.1007/978-3-642-33466-
5_6/.

Sharma:2017:VCS

[SWMV17] Vaibhav Sharma, Michael W. Whalen, Stephen McCamant, and Willem Visser. Veritesting challenges in symbolic execution of Java. ACM SIGSOFT Software Engineering Notes, 42(4):1– 5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simon:2015:STH

[SWU^{+}15] Doug Simon, Christian Wimmer, Bernhard Urban, Gilles Duboscq, Lukas Stadler, and Thomas Würthinger. Snippets: Taking the high road to a low level. ACM Transactions on Architecture and Code Optimization, 12(2):20:1–20:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Savrun-Yeniceri:2014:EHI

[SYZZ^{+}14] Gülfem Savrun-Yeniçeri, Wei Zhang, Huahan Zhang, Eric Seckler, Chen Li, Stefan Brunthaler, Per Larsen, and Michael Franz. Efficient hosted interpreters on the JVM. ACM Transactions on Architecture and Code Optimization, 11(1): 9:1–9:24, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/2532642.

Servetto:2010:MMC

[SZ10] Marco Servetto and Elena Zucca. MetaFJig: a metacircular composition language for Java-like classes. ACM SIGPLAN Notices, 45 (10):464–483, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siegel:2011:AFV

[SZ11] Stephen F. Siegel and Timothy K. Zirkel. Automatic formal verification of MPIbased parallel programs. ACM SIGPLAN Notices, 46 (8):309–310, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Shen:2019:IRA

[SZZ^{+}19] Bo Shen, Wei Zhang, Haiyan Zhao, Guangtai Liang, Zhi Jin, and Qianxiang Wang. IntelliMerge: a refactoring-aware software merging technique. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 170:1–170:28, October 2019.

                                        URL  https://dl.acm.
                                        org/doi/abs/10.1145/3360596.

Tamayo:2012:UBD

[TABS12] Juan M. Tamayo, Alex Aiken, Nathan Bronson, and Mooly Sagiv. Understanding the behavior of database operations under program control. ACM SIG-PLAN Notices, 47(10):983– 996, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taft:2013:TPS

[Taf13] S. Tucker Taft. Tutorial: proving safety of parallel /multi-threaded programs. ACM SIGADA Ada Letters, 33(3):1–2, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tanyalcin:2018:LVL

[TAF^{+}18] Ibrahim Tanyalcin, Carla Al Assaf, Julien Ferte, François Ancien, Taushif Khan, Guillaume Smits, Marianne Rooman, and Wim Vranken. Lexicon visualization library and JavaScript for scientific data visualization. Computing in Science and Engineering, 20(1):50–65, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291800/.

Taibi:2013:ROS

[Tai13]  Fathi Taibi.  Reusabil-

ity of open-source program code: a conceptual model and empirical investigation. ACM SIGSOFT Software Engineering Notes, 38(4): 1–5, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tarau:2011:IST

[Tar11] Paul Tarau. Integrated symbol table, engine and heap memory management in multi-engine Prolog. ACM SIGPLAN Notices, 46(11):129–138, November 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’11 conference proceedings.

Tosch:2014:SPA

[TB14] Emma Tosch and Emery D. Berger. SurveyMan: programming and automatically debugging surveys. ACM SIGPLAN Notices, 49 (10):197–211, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomson:2015:LHB

[TD15] Paul Thomson and Alastair F. Donaldson. The lazy happens-before relation: better partial-order reduction for systematic concurrency testing. ACM SIGPLAN Notices, 50(8): 259–260, August 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tomescu:2017:CEN

[TD17] A. Tomescu and S. Devadas. Catena: Efficient non-equivocation via Bitcoin. In 2017 IEEE Symposium on Security and Privacy (SP), pages 393–409. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2017.

Teodorovici:2012:BRC

[Teo12] Vasile G. Teodorovici. Book review: Continuous testing with Ruby, Rails and JavaScript by Ben Rady and Rod Coffin. ACM SIG-SOFT Software Engineering Notes, 37(1):36, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRL

[Teo13] Vasile G. Teodorovici. Book review: Learning JavaScript: a hands-on guide to the fundamentals of modern JavaScript by Tim Wright. ACM SIGSOFT Software Engineering Notes, 38(3): 35–36, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teyton:2014:SLM

[TFPB14] Cédric Teyton, Jean-Rémy Falleri, Marc Palyart, and Xavier Blanc. A study of library migrations in Java. Journal of Software: Evolution and Process, 26 (11):1030–1052, November 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tommasel:2017:SJL

[TGZ17] Antonela Tommasel, Daniela Godoy, and Alejandro Zunino. SMArtOp: a Java library for distributing highdimensional sparse-matrix arithmetic operations. Science of Computer Programming, 150(??):26–30, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642317301260.

Tu:2014:PPP

[THC^{+}14] Chia-Heng Tu, Hui-Hsin Hsu, Jen-Hao Chen, Chun-Han Chen, and Shih-Hao Hung. Performance and power profiling for emulated Android systems. ACM Transactions on Design Automation of Electronic Systems, 19(2):10:1– 10:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tran-Jorgensen:2018:ATV

[TJLL18] Peter W. V. Tran-Jørgensen, Peter Gorm Larsen, and Gary T. Leavens. Automated translation of VDM to JML-annotated Java. International Journal on Software Tools for Technology Transfer (STTT), 20 (2):211–235, April 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

https://link.springer.
com/article/10.1007/s10009-
017-0448-3.

Tsai:2015:JPI

[TKL^{+}15] Chun-Jen Tsai, Han-Wen Kuo, Zigang Lin, Zi-Jing Guo, and Jun-Fu Wang. A Java processor IP design for embedded SoC. ACM Transactions on Embedded Computing Systems, 14(2): 35:1–35:??, March 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thiessen:2017:CTP

[TL17] Rei Thiessen and Ondrej Lhoták. Context transformations for pointer analysis. ACM SIGPLAN Notices, 52(6):263–277, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tate:2011:TWJ

[TLL11] Ross Tate, Alan Leung, and Sorin Lerner. Taming wildcards in Java’s type system. ACM SIG-PLAN Notices, 46(6):614– 627, June 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tetali:2013:MSA

[TLMM13] Sai Deep Tetali, Mohsen Lesani, Rupak Majumdar, and Todd Millstein. Mr-Crypt: static analysis for secure cloud computations. ACM SIGPLAN Notices, 48 (10):271–286, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Tan:2017:EPP

[TLX17] Tian Tan, Yue Li, and Jingling Xue. Efficient and precise points-to analysis: modeling the heap by merging equivalent automata. ACM SIGPLAN Notices, 52(6):278–291, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terra:2013:QCC

[TMVB13] Ricardo Terra, Luis Fernando Miranda, Marco Tulio Valente, and Roberto S. Bigonha. Qualitas.class corpus: a compiled version of the qualitas corpus. ACM SIGSOFT Software Engineering Notes, 38(5):1–4, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thakur:2019:PFP

[TN19] Manas Thakur and V. Krishna Nandivada. PYE: a framework for precise-yet-efficient just-in-time analyses for Java programs. ACM Transactions on Programming Languages and Systems, 41(3):16:1–16:??, July 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3337794.

Toledo:2012:AJA

[TNTN12] R. Toledo, A. Nunez, E. Tanter, and J. Noye. Aspectizing Java access control. IEEE Transactions on Software Engineering, 38(1):101–117, January/February 2012. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
5680915.

Topley:2011:JDG

[Top11] Kim Topley. JavaFX developer’s guide. Developer’s library. Addison-Wesley, Addison-Wesley, 2011. ISBN 0-321-60165-3. xxxix + 1110 pp. LCCN ???? Updated for JavaFX 1.3.

Toffola:2015:PPY

[TPG15] Luca Della Toffola, Michael Pradel, and Thomas R. Gross. Performance problems you can fix: a dynamic analysis of memoization opportunities. ACM SIG-PLAN Notices, 50(10):607– 622, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taboada:2013:JHP

[TRE^{+}13] Guillermo L. Taboada, Sabela Ramos, Roberto R.

Expósito, Juan Touriño, and Ramón Doallo. Java in the High Performance Computing arena: Research, practice and experience. Science of Computer Programming, 78(5):425–444, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642311001420.

Taboada:2011:DEJ

[TRTD11] Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Design of efficient Java messagepassing collectives on multicore clusters. The Journal of Supercomputing, 55(2):126–154, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
55&issue=2&spage=126.

Takikawa:2012:GTF

[TSD^{+}12] Asumu Takikawa, T. Stephen Strickland, Christos Dimoulas, Sam Tobin-Hochstadt, and Matthias Felleisen. Gradual typing for firstclass classes. ACM SIG-PLAN Notices, 47(10):793– 810, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Toledo:2011:ACJ

[TT11] Rodolfo Toledo and Eric Tanter. Access control in JavaScript. IEEE Software, 28(5):76–84, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Taboada:2011:DLC

[TTD^{+}11] Guillermo L. Taboada, Juan Touriño, Ramón Doallo, Aamir Shafi, Mark Baker, and Bryan Carpenter. Device level communication libraries for highperformance computing in Java. Concurrency and Computation: Practice and Experience, 23(18):2382– 2403, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taboada:2012:FMS

[TTD12] Guillermo L. Taboada, Juan Touriño, and Ramón Doallo. F-MPJ: scalable Java message-passing communications on parallel systems. The Journal of Supercomputing, 60(1):117–140, April 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0920-8542&volume=
                                        60&issue=1&spage=117.

Tatsubori:2010:EJT

[TTS^{+}10] Michiaki Tatsubori, Akihiko Tozawa, Toyotaro Suzumura, Scott Trent, and Tamiya Onodera. Evaluation of a just-in-time compiler retrofitted for PHP. ACM SIGPLAN Notices, 45 (7):121–132, July 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Torlak:2010:MCA

[TVD10] Emina Torlak, Mandana Vaziri, and Julian Dolby. MemSAT: checking axiomatic specifications of memory models. ACM SIG-PLAN Notices, 45(6):341– 350, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tardieu:2012:WSS

[TWL12] Olivier Tardieu, Haichuan Wang, and Haibo Lin. A work-stealing scheduler for X10’s task parallelism with suspension. ACM SIG-PLAN Notices, 47(8):267– 276, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Toegl:2012:SSJ

[TWNH12] Ronald Toegl, Thomas Winkler, Mohammad Nauman, and Theodore W. Hong. Specification and standardization of a Java Trusted Computing API. Software—Practice and Experience, 42(8):945–965, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Titzer:2010:ICR

[TWSC10] Ben L. Titzer, Thomas Würthinger, Doug Simon, and Marcelo Cintra. Improving compiler-runtime separation with XIR. ACM SIGPLAN Notices, 45(7): 39–50, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Teng:2010:TPA

[TWX^{+}10] Q. M. Teng, H. C. Wang, Z. Xiao, P. F. Sweeney, and E. Duesterwald. THOR: a performance analysis tool for Java applications running on multicore systems. IBM Journal of Research and Development, 54(5): 4:1–4:17, ???? 2010. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Urma:2015:JAL

[UFM15] Raoul-Gabriel Urma, Mario Fusco, and Alan Mycroft.

Java 8 in action: lambdas, streams, and functionalstyle programming. Manning Publications, Greenwich, CT, USA, 2015. ISBN 1-61729-199-4 (paperback). xxviii + 394 pp. LCCN QA76.73.J38 U76 2015.

URL http://proquest.
safaribooksonline.com/
?fpi=9781617291999;
http://proquest.tech.
safaribooksonline.de/
9781617291999.

Ugawa:2010:IRB

[UIY10] Tomoharu Ugawa, Hideya Iwasaki, and Taiichi Yuasa. Improved replication-based incremental garbage collection for embedded systems. ACM SIGPLAN Notices, 45(8):73–82, August 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ugawa:2014:ROP

[UJR14] Tomoharu Ugawa, Richard E. Jones, and Carl G. Ritson. Reference object processing in on-the-fly garbage collection. ACM SIGPLAN Notices, 49(11): 59–69, November 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Upadhyaya:2010:UDS

[UMP10] Gautam Upadhyaya, Samuel P. Midkiff, and Vijay S. Pai.

Using data structure knowledge for efficient lock generation and strong atomicity. ACM SIGPLAN Notices, 45(5):281–292, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uva:2018:AWJ

[UPR^{+}18] Marcelo Uva, Pablo Ponzio, Germán Regis, Nazareno Aguirre, and Marcelo F. Frias. Automated workarounds from Java program specifications based on SAT solving. International Journal on Software Tools for Technology Transfer (STTT), 20(6):665–688, November 2018. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

https://link.springer.
com/article/10.1007/s10009-
018-0503-8.

Upadhyaya:2015:EML

[UR15] Ganesha Upadhyaya and Hridesh Rajan. Effectively mapping linguistic abstractions for message-passing concurrency to threads on the Java Virtual Machine. ACM SIGPLAN Notices, 50 (10):840–859, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ugawa:2018:TSL

[URJ18] Tomoharu Ugawa, Carl G.

Ritson, and Richard E. Jones. Transactional Sapphire: Lessons in highperformance, on-the-fly garbage collection. ACM Transactions on Programming Languages and Systems, 40(4):15:1–15:??, December 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ureche:2013:MIS

[UTO13] Vlad Ureche, Cristian Talau, and Martin Odersky. Miniboxing: improving the speed to code size tradeoff in parametric polymorphism translations. ACM SIGPLAN Notices, 48(10): 73–92, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Vilk:2014:DBB

[VB14a] John Vilk and Emery D. Berger. Doppio: breaking the browser language barrier. ACM SIGPLAN Notices, 49(6):508–518, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vouillon:2014:BJJ

[VB14b] Jérôme Vouillon and Vincent Balat. From bytecode to JavaScript: the Js of ocaml compiler. Software—Practice and Experience, 44(8):951–972, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vilk:2018:BAD

[VB18] John Vilk and Emery D.

Berger. BLeak: automatically debugging memory leaks in web applications. ACM SIGPLAN Notices, 53 (4):15–29, April 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Villazon:2010:ARA

[VBAM10a] Alex Villazón, Walter Binder, Danilo Ansaloni, and Philippe Moret. Advanced runtime adaptation for Java. ACM SIGPLAN Notices, 45(2): 85–94, February 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Villazon:2010:HCA

[VBAM10b] Alex Villazón, Walter Binder, Danilo Ansaloni, and Philippe Moret. HotWave: creating adaptive tools with dynamic aspect-oriented programming in Java. ACM SIG-PLAN Notices, 45(2):95–98, February 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vidal:2016:ECJ

[VBDPM16] S. Vidal, A. Bergel, J. A. Díaz-Pace, and C. Marcos. Over-exposed classes in Java: an empirical study. Computer Languages, Systems and Structures, 46(??): 1–19, November 2016. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1477842415300531.

Villazon:2011:CAW

[VBMA11] Alex Villazón, Walter Binder, Philippe Moret, and Danilo Ansaloni. Comprehensive aspect weaving for Java. Science of Computer Programming, 76(11):1015– 1036, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vidal:2016:UAE

[VBMDP16] Santiago A. Vidal, Alexandre Bergel, Claudia Marcos, and J. Andrés Díaz-Pace. Understanding and addressing exhibitionism in Java empirical research about method accessibility. Empirical Software Engineering, 21(2): 483–516, April 2016. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

                                        http://link.springer.
                                        com/accesspage/article/

10.1007/s10664-015-9365-
9.

Vidal:2018:ARB

[VBZ^{+}18] Santiago Vidal, Iñaki Berra, Santiago Zulliani, Claudia Marcos, and J. Andrés Díaz Pace. Assessing the refactoring of brain methods. ACM Transactions on Software Engineering and Methodology, 27(1):2:1– 2:??, June 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

vanderMerwe:2012:VAA

[vdMvdMV12] Heila van der Merwe, Brink van der Merwe, and Willem Visser. Verifying Android applications using Java PathFinder. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Viotti:2017:HRH

[VDV17] Paolo Viotti, Dan Dobre, and Marko Vukolić. Hybris: Robust hybrid cloud storage. ACM Transactions on Storage, 13(3):27:1–27:??, October 2017. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

VanLoan:2010:ITC

[VF10] Charles F. Van Loan and K.-Y. Daisy Fan. Insight through computing: a MATLAB introduction to computational science and engineering. SIAM Press, Philadelphia, PA, USA, 2010. ISBN 0-89871-691-8. xviii + 434 pp. LCCN QA297 .V25 2010. URL http:

//www.loc.gov/catdir/
enhancements/fy1007/2009030277-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy1007/2009030277-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1007/2009030277-
t.html.

Vega-Gisbert:2016:DIJ

[VGRS16] Oscar Vega-Gisbert, Jose E. Roman, and Jeffrey M. Squyres. Design and implementation of Java bindings in Open MPI. Parallel Computing, 59(??):1–20, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819116300758.

Vikas:2014:MGA

[VGS14] Vikas, Nasser Giacaman, and Oliver Sinnen. Multiprocessing with GUIawareness using OpenMPlike directives in Java. Parallel Computing, 40(2):69– 89, February 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819113001439.

Vitek:2014:CTR

[Vit14] Jan Vitek. The case for the three R’s of systems research: repeatability, reproducibility and rigor. ACM SIGPLAN Notices, 49(7): 115–116, July 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vitek:2012:ISI

[VK12] Jan Vitek and Tomas Kalibera. Introduction to the Special Issue on Java Technologies for Real-Time and Embedded Systems. Concurrency and Computation: Practice and Experience, 24 (8):751–752, ???? 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

VanCutsem:2010:PDP

[VM10] Tom Van Cutsem and Mark S. Miller. Proxies: design principles for robust object-oriented intercession APIs. ACM SIGPLAN Notices, 45(12):59–72, December 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanCutsem:2015:RTC

[VM15] Tom Van Cutsem and Mark S. Miller. Robust trait composition for JavaScript. Science of Computer Programming, 98 (part 3)(??):422–438, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642312002079.

Verdu:2016:PSA

[VP16] Javier Verdu and Alex Pajuelo. Performance scalability analysis of JavaScript applications with Web Workers. IEEE Computer Architecture Letters, 15(2): 105–108, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

VanderHart:2010:PC

[VS10] Luke VanderHart and Stuart Sierra. Practical Clojure. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-7231-7, 1-4302-7230-9 (e-book). xvi + 210 pp. LCCN ????

V:2011:BBI

[VS11] Sharath Chandra V. and S. Selvakumar. BIXSAN: browser independent XSS sanitizer for prevention of XSS attacks. ACM SIG-SOFT Software Engineering Notes, 36(5):1–7, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Varier:2017:TNJ

[VSG17] K. Muraleedhara Varier, V. Sankar, and M. P. Gangadathan. TrackEtching

— a Java based code for etched track profile calculations in SSNTDs. Computer Physics Communications, 218(??):43–47, September 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465517301273.

VanNieuwpoort:2010:SHL

[VWJB10] Rob V. Van Nieuwpoort, Gosia Wrzesińska, Ceriel J. H. Jacobs, and Henri E. Bal. Satin: a high-level and efficient grid programming model. ACM Transactions on Programming Languages and Systems, 32 (3):9:1–9:39, March 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Vechev:2010:PPC

[VYY10] Martin Vechev, Eran Yahav, and Greta Yorsh. PHALANX: parallel checking of expressive heap assertions. ACM SIGPLAN Notices, 45(8):41–50, August 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wijayarathna:2019:WJC

[WA19] Chamila Wijayarathna and Nalin Asanka Gamagedara Arachchilage. Why Johnny can’t develop a secure application? A usability analysis of Java Secure Socket Extension API. Computers & Security, 80(??):54– 73, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167404818304887.

Wurthinger:2011:SAR

[WAB^{+}11] Thomas Würthinger, Danilo Ansaloni, Walter Binder, Christian Wimmer, and Hanspeter Mössenböck. Safe and atomic run-time code evolution for Java and its application to dynamic AOP. ACM SIGPLAN Notices, 46(10):825–844, October 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Walker:2012:SNJ

[Wal12] Henry M. Walker. SIGCSE by the numbers: JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(1):8, January 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wampler:2011:FPJ

[Wam11]  Dean Wampler.  Func-

tional programming for Java developers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-1265-9, 1-4493-1103-2. xi + 72 pp. LCCN QA76.62 .W36 2011.

URL http://proquest.
safaribooksonline.com/
9781449312657.

Wang:2011:EEU

[Wan11] Alf Inge Wang. Extensive evaluation of using a game project in a software architecture course. ACM Transactions on Computing Education, 11(1):5:1– 5:??, February 2011. CO-DEN ???? ISSN 1946-6226.

Wurthinger:2011:AED

[WBA^{+}11] Thomas Würthinger, Walter Binder, Danilo Ansaloni, Philippe Moret, and Hanspeter Mössenböck. Applications of enhanced dynamic code evolution for Java in GUI development and dynamic aspect-oriented programming. ACM SIGPLAN Notices, 46(2):123–126, February 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2018:HSA

[WBHN18] Kunshan Wang, Stephen M. Blackburn, Antony L. Hosking, and Michael Norrish. Hop, skip, & jump: Practical on-stack replacement for a cross-platform languageneutral VM. ACM SIG-PLAN Notices, 53(3):1–16, March 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welch:2010:ABS

[WBM^{+}10] Peter Welch, Neil Brown, James Moores, Kevin Chalmers, and Bernhard Sputh. Alting barriers: synchronisation with choice in Java using JCSP. Concurrency and Computation: Practice and Experience, 22(8): 1049–1062, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wellings:2016:ISC

[WCB16] A. J. Wellings, V. Cholpanov, and A. Burns. Implementing safety-critical Java missions in Ada. ACM SIGADA Ada Letters, 36 (1):51–62, June 2016. CO-DEN AALEE5. ISSN 0736-721X.

Wood:2014:LLD

[WCG14] Benjamin P. Wood, Luis Ceze, and Dan Grossman. Low-level detection of language-level data races with LARD. ACM SIGARCH Computer Architecture News, 42(1):671– 686, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wang:2018:PBJ

[WCG^{+}18] Kaiyuan Wang, Hayes Converse, Milos Gligoric, Sasa Misailovic, and Sarfraz Khurshid. A progress bar for the JPF search using program executions. ACM SIGSOFT Software Engineering Notes, 43(4):55, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2019:DEJ

[WCST19] Yuchen Wang, Kwok Sun Cheng, Myoungkyu Song, and Eli Tilevich. A declarative enhancement of JavaScript programs by leveraging the Java metadata infrastructure. Science of Computer Programming, 181(??):27–46, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167642319300723.

Wilcox:2018:VVH

[WFF18] James R. Wilcox, Cormac Flanagan, and Stephen N. Freund. VerifiedFT: a verified, high-performance precise dynamic race detector. ACM SIGPLAN Notices, 53(1):354–367, January 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wagner:2011:SJV

[WGF11] Gregor Wagner, Andreas Gal, and Michael Franz. “slimming” a Java virtual machine by way of cold code removal and optimistic partial program loading. Science of Computer Programming, 76(11):1037– 1053, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wagner:2011:CMM

[WGW^{+}11] Gregor Wagner, Andreas Gal, Christian Wimmer, Brendan Eich, and Michael Franz. Compartmental memory management in a modern web browser. ACM SIGPLAN Notices, 46(11): 119–128, November 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’11 conference proceedings.

Wu:2011:RTS

[WHIN11] Peng Wu, Hiroshige Hayashizaki, Hiroshi Inoue, and Toshio Nakatani. Reducing trace selection footprint for largescale Java applications without performance loss. ACM SIGPLAN Notices, 46 (10):789–804, October 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Wimmer:2013:MAV

[WHV^{+}13] Christian Wimmer, Michael Haupt, Michael L. Van De Vanter, Mick Jordan, Laurent Daynès, and Douglas Simon. Maxine: an approachable virtual machine for, and in, Java. ACM Transactions on Architecture and Code Optimization, 9(4):30:1–30:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wellings:2012:AEH

[WK12] Andy Wellings and Min-Seong Kim. Asynchronous event handling and Safety Critical Java. Concurrency and Computation: Practice and Experience, 24(8):813– 832, ???? 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2017:JRJ

[WKG17] Kaiyuan Wang, Sarfraz Khurshid, and Milos Gligoric. JPR: Replaying JPF traces using standard JVM. ACM SIGSOFT Software Engineering Notes, 42(4):1– 5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wade:2017:AVJ

[WKJ17] April W. Wade, Prasad A. Kulkarni, and Michael R. Jantz. AOT vs. JIT: impact of profile data on code quality. ACM SIG-PLAN Notices, 52(4):1–10, May 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2019:TRC

[WLL19] Lulu Wang, Jingyue Li, and Bixin Li. Tracking runtime concurrent dependences in Java threads using thread control profiling. The Journal of systems and software, 148(??):116–131, February 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0164121218302395.

Wimmer:2010:AFD

[WM10] Christian Wimmer and Hanspeter Mössenbösck. Automatic feedback-directed object fusing. ACM Transactions on Architecture and Code Optimization, 7(2): 7:1–7:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wendykier:2010:PCH

[WN10] Piotr Wendykier and James G.

Nagy. Parallel Colt: a highperformance Java library for scientific computing and image processing. ACM Transactions on Mathematical Software, 37(3):31:1– 31:22, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Witman:2010:TBR

[WR10] Paul D. Witman and Terry Ryan. Think big for reuse. Communications of the ACM, 53(1):142–147, January 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Westbrook:2010:MJM

[WRI^{+}10] Edwin Westbrook, Mathias Ricken, Jun Inoue, Yilong Yao, Tamer Abdelatif, and Walid Taha. Mint: Java multi-stage programming using weak separability. ACM SIGPLAN Notices, 45(6):400–411, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watt:2019:WW

[WRPP19] Conrad Watt, Andreas Rossberg, and Jean Pichon-Pharabod. Weakening WebAssembly. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):133:1–133:28, October 2019. URL

https://dl.acm.org/doi/
abs/10.1145/3360559.

Wimmer:2019:IOS

[WSH^{+}19] Christian Wimmer, Codrut Stancu, Peter Hofer, Vojin Jovanovic, Paul Wögerer, Peter B. Kessler, Oleg Pliss, and Thomas Würthinger. Initialize once, start fast: application initialization at build time. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):184:1–184:29, October 2019. URL

https://dl.acm.org/doi/
abs/10.1145/3360610.

Wehr:2010:JBP

[WT10] Stefan Wehr and Peter Thiemann. JavaGI in the battlefield: practical experience with generalized interfaces. ACM SIGPLAN Notices, 45(2):65–74, February 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wehr:2011:JIT

[WT11] Stefan Wehr and Peter Thiemann. JavaGI: The interaction of type classes with interfaces and inheritance. ACM Transactions on Programming Languages and Systems, 33(4):12:1– 12:83, July 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wang:2018:IDG

[WWG^{+}18] Wenwen Wang, Jiacheng Wu, Xiaoli Gong, Tao Li, and Pen-Chung Yew. Improving dynamicallygenerated code performance on dynamic binary translators. ACM SIGPLAN Notices, 53(3):17–30, March 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wurthinger:2017:PPE

[WWH^{+}17] Thomas Würthinger, Christian Wimmer, Christian Humer, Andreas Wöß, Lukas Stadler, Chris Seaton, Gilles Duboscq, Doug Simon, and Matthias Grimmer. Practical partial evaluation for high-performance dynamic language runtimes. ACM SIGPLAN Notices, 52(6):662–676, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wurthinger:2013:USD

[WWS13] Thomas Würthinger, Christian Wimmer, and Lukas Stadler. Unrestricted and safe dynamic code evolution for Java. Science of Computer Programming, 78(5): 481–498, May 1, 2013. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167642311001456.

Wei:2016:ESD

[WXR16] Shiyi Wei, Franceska Xhakaj, and Barbara G. Ryder. Empirical study of the dynamic behavior of JavaScript objects. Software—Practice and Experience, 46(7):867– 889, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2017:CJ

[WZdSOS17] Yanlin Wang, Haoyuan Zhang, Bruno C. d. S. Oliveira, and Marco Servetto. Classless Java. ACM SIG-PLAN Notices, 52(3):14– 24, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2019:OTA

[WZK^{+}19] Qingyang Wang, Shungeng Zhang, Yasuhiko Kanemasa, Calton Pu, Balaji Palanisamy, Lilian Harada, and Motoyuki Kawaba. Optimizing N-tier application scalability in the cloud: A study of soft resource allocation. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4 (2):10:1–10:??, June 2019. CODEN ???? ISSN 2376-3639. URL https:

//dl.acm.org/citation.
cfm?id=3326120.

Wu:2018:EBJ

[WZL^{+}18] Mingyu Wu, Ziming Zhao, Haoyu Li, Heting Li, Haibo Chen, Binyu Zang, and Haibing Guan. Espresso: Brewing Java for more nonvolatility with non-volatile memory. ACM SIGPLAN Notices, 53(2):70–83, February 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2019:EEG

[XGD^{+}19] Lijie Xu, Tian Guo, Wensheng Dou, Wei Wang, and Jun Wei. An experimental evaluation of garbage collectors on big data applications. Proceedings of the VLDB Endowment, 12(5): 570–583, January 2019. CO-DEN ???? ISSN 2150-8097.

Xi:2012:MDA

[XHH12] Kai Xi, Jiankun Hu, and Fengling Han. Mobile device access control: an improved correlation based face authentication scheme and its Java ME application. Concurrency and Computation: Practice and Experience, 24 (10):1066–1085, July 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2010:FLU

[XMA^{+}10] Guoqing Xu, Nick Mitchell, Matthew Arnold, Atanas Rountev, Edith Schonberg, and Gary Sevitsky. Finding low-utility data structures. ACM SIGPLAN Notices, 45(6):174–186, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2014:SRB

[XMA^{+}14] Guoqing Xu, Nick Mitchell, Matthew Arnold, Atanas Rountev, Edith Schonberg, and Gary Sevitsky. Scalable runtime bloat detection using abstract dynamic slicing. ACM Transactions on Software Engineering and Methodology, 23 (3):23:1–23:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xuan:2017:NAR

[XMD^{+}17] J. Xuan, M. Martinez, F. DeMarco, M. Clément, S. L. Marcote, T. Durieux, D. Le Berre, and M. Monperrus. Nopol: Automatic repair of conditional statement bugs in Java programs. IEEE Transactions on Software Engineering, 43(1):34–55, January 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7463060.

Xu:2010:DIU

[XR10] Guoqing Xu and Atanas Rountev. Detecting inefficientlyused containers to avoid bloat. ACM SIGPLAN Notices, 45(6):160–173, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2013:PML

[XR13] Guoqing Xu and Atanas Rountev. Precise memory leak detection for Java software using container profiling. ACM Transactions on Software Engineering and Methodology, 22 (3):17:1–17:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xue:2012:RJC

[Xue12] Jingling Xue. Rethinking Java call stack design for tiny embedded devices. ACM SIGPLAN Notices, 47 (5):1–10, May 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Xue:2019:ASC

[XXCL19] Y. Xue, Z. Xu, M. Chandramohan, and Y. Liu. Accurate and scalable crossarchitecture cross-OS binary code search with emulation. IEEE Transactions on Software Engineering, 45(11):1125–1149, November 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Xie:2013:AAE

[XXZ13] Xinwei Xie, Jingling Xue, and Jie Zhang. Acculock: accurate and efficient detection of data races. Software —Practice and Experience, 43(5):543–576, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2012:MPD

[YCYC12] Cheng-Hong Yang, Yu-Huei Cheng, Cheng-Huei Yang, and Li-Yeh Chuang. Mutagenic primer design for mismatch PCR-RFLP SNP genotyping using a genetic algorithm. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(3):837–845, May 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Yi:2015:CTC

[YDFF15] Jaeheon Yi, Tim Disney, Stephen N. Freund, and Cormac Flanagan. Cooperative types for controlling thread interference in Java. Science of Computer Programming, 112 (part 3)(??):227–260, November 15, 2015. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642315001483.

Yang:2013:CPP

[YHY13] Chao Yang, Zengyou He, and Weichuan Yu. A combinatorial perspective of the protein inference problem. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(6): 1542–1547, November 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Yoo:2014:WRR

[YK14] Danny Yoo and Shriram Krishnamurthi. Whalesong: running Racket in the browser. ACM SIG-PLAN Notices, 49(2):97– 108, February 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Yang:2019:MGL

[YKA^{+}19] Bing Yang, Kenneth B.

Kent, Eric Aubanel, Stephen MacKay, and Tobi Agila.

A multi-granularity locking scheme for Java PackedObjects based on a concurrent multiway tree. Concurrency and Computation: Practice and Experience, 31(11): e5024:1–e5024:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2017:EJV

[YKM17] Byung-Sun Yang, Jae-Yun Kim, and Soo-Mook Moon. Exceptionization: a Java VM optimization for non-Java languages. ACM Transactions on Architecture and Code Optimization, 14(1):5:1–5:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yessenov:2017:DAD

[YKSL17] Kuat Yessenov, Ivan Kuraj, and Armando Solar-Lezama. DemoMatch: API discovery from demonstrations. ACM SIGPLAN Notices, 52(6):64–78, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yim:2019:TFS

[YMHB19] Keun Soo Yim, Iliyan Malchev, Andrew Hsieh, and Dave Burke. Treble: Fast software updates by creating an equilibrium in an active software ecosystem of globally distributed stakeholders. ACM Transactions on Embedded Computing Systems, 18(5s):104:1–104:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358237.

Yang:2010:JIP

[YP10] L. Yang and M. R. Poppleton. Java implementation platform for the integrated state- and event-based specification in PROB. Concurrency and Computation: Practice and Experience, 22(8):1007–1022, June 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yerima:2012:AMB

[YPMM12] Suleiman Y. Yerima, Gerard P. Parr, Sally I. Mc-Clean, and Philip J. Morrow. Adaptive measurementbased policy-driven QoS management with fuzzyrule-based resource allocation. Future Internet, 4(3): 646–671, July 04, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/3/
646.

Yi:2015:SCC

[YQTR15] Jooyong Yi, Dawei Qi, Shin Hwei Tan, and Abhik Roychoudhury. Software change contracts. ACM Transactions on Software Engineering and Methodology, 24(3):18:1–18:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yiapanis:2013:OSR

[YRHBL13] Paraskevas Yiapanis, Demian Rosas-Ham, Gavin Brown, and Mikel Luján. Optimizing software runtime systems for speculative parallelization. ACM Transactions on Architecture and Code Optimization, 9(4): 39:1–39:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yahav:2010:VSP

[YS10] Eran Yahav and Mooly Sagiv. Verifying safety properties of concurrent heapmanipulating programs. ACM Transactions on Programming Languages and Systems, 32(5):18:1–18:50, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yan:2017:AAA

[YSCX17] Hua Yan, Yulei Sui, Shiping Chen, and Jingling Xue. AutoFix: an automated approach to memory leak fixing on value-flow slices for C programs. ACM SIGAPP Applied Computing Review, 16(4):38–50, January 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3040575.3040579.

Yue:2013:MSI

[YW13] Chuan Yue and Haining Wang. A measurement study of insecure JavaScript practices on the Web. ACM Transactions on the Web (TWEB), 7(2): 7:1–7:??, May 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yu:2018:NFN

[YWW^{+}18] Chunjiang Yu, Wentao Wu, Jing Wang, Yuxin Lin, Yang, Jiajia Chen, Fei Zhu, and Bairong Shen. NGS-FC: a next-generation sequencing data format converter. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(5):1683–1691, September 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Yan:2019:ACL

[YXS^{+}19] M. Yan, X. Xia, E. Shihab, D. Lo, J. Yin, and X. Yang. Automating change-level self-admitted technical debt determination. IEEE Transactions on Software Engineering, 45 (12):1211–1229, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zakas:2010:HPJ

[Zak10] Nicholas C. Zakas. High performance JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2010. ISBN 0-596-80279-x, 1-4493-8230-4. xviii + 209 pp. LCCN ????

Zakhour:2012:JTS

[Zak12] Sharon Zakhour. The Java tutorial: a short course on the basics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2012. ISBN 0-13-276169-6 (paperback). ???? pp. LCCN ????

Zakai:2018:FPW

[Zak18] Alon Zakai. Fast physics on the Web using C++, JavaScript, and Emscripten. Computing in Science and Engineering, 20(1):11–19, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        document/8254329/.

Zheng:2015:APP

[ZBB15] Yudi Zheng, Lubomír Bulej, and Walter Binder. Accurate profiling in the presence of dynamic compilation. ACM SIGPLAN Notices, 50(10):433–450, October 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2017:ACE

[ZBB17] Minjia Zhang, Swarnendu Biswas, and Michael D. Bond. Avoiding consistency exceptions under strong memory models. ACM SIG-PLAN Notices, 52(9):115– 127, September 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2019:DSJ

[ZBvdB^{+}19] Dan Zhang, Dragan Bošnački, Mark van den Brand, Cornelis Huizing, Bart Jacobs, Ruurd Kuiper, and Anton Wijs. Dependency safety for Java — Implementing and testing failboxes. Science of Computer Programming, 184(??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642319301145.

Zhang:2015:SYB

[ZCdSOvdS15] Haoyuan Zhang, Zewei Chu, Bruno C. d. S. Oliveira, and Tijs van der Storm. Scrap your boilerplate with object algebras. ACM SIG-PLAN Notices, 50(10):127– 146, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeuch:2019:AES

[ZDK^{+}19] Steffen Zeuch, Bonaventura Del Monte, Jeyhun Karimov, Clemens Lutz, Manuel Renz, Jonas Traub, Sebastian Breß, Tilmann Rabl, and Volker Markl. Analyzing efficient stream processing on modern hardware. Proceedings of the VLDB Endowment, 12(5):516–530, January 2019. CODEN ???? ISSN 2150-8097.

Zschaler:2014:SJF

[ZDS14] Steffen Zschaler, Birgit Demuth, and Lothar Schmitz. Salespoint: a Java framework for teaching objectoriented software development. Science of Computer Programming, 79(??):189– 203, January 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016764231200069X.

Zuo:2016:LOF

[ZFK^{+}16] Zhiqiang Zuo, Lu Fang, Siau-Cheng Khoo, Guoqing Xu, and Shan Lu. Lowoverhead and fully automated statistical debugging with abstraction refinement. ACM SIGPLAN Notices, 51(10):881–896, October 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2012:PTI

[Zha12] Tian Zhao. Polymorphic type inference for scripting languages with object extensions. ACM SIGPLAN Notices, 47(2):37–50, February 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2015:LOS

[ZHCB15] Minjia Zhang, Jipeng Huang, Man Cao, and Michael D. Bond. Low-overhead software transactional memory with progress guarantees and strong semantics. ACM SIGPLAN Notices, 50(8):97–108, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2012:RAJ

[ZHL^{+}12] Ying Zhang, Gang Huang, Xuanzhe Liu, Wei Zhang, Hong Mei, and Shunxiang Yang. Refactoring Android Java code for ondemand computation offloading. ACM SIGPLAN Notices, 47(10):233–248, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zacharopoulos:2017:EMM

[ZIvdS17] Theologos Zacharopoulos, Pablo Inostroza, and Tijs van der Storm. Extensible modeling with managed data in Java. ACM SIG-PLAN Notices, 52(3):25– 35, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zheng:2016:CMD

[ZKB^{+}16] Yudi Zheng, Stephen Kell, Lubomir Bulej, Haiyang Sun, and Walter Binder. Comprehensive multiplatform dynamic program analysis for Java and Android. IEEE Software, 33(4):55–63, July/ August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        so/2016/04/mso2016040055-
                                        abs.html.

Zhao:2013:INT

[ZLB^{+}13] Jisheng Zhao, Roberto Lublinerman, Zoran Budimlić, Swarat Chaudhuri, and Vivek Sarkar. Isolation for nested task parallelism. ACM SIGPLAN Notices, 48 (10):571–588, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Zhang:2014:AIO

[ZLBF14] Wei Zhang, Per Larsen, Stefan Brunthaler, and Michael Franz. Accelerating iterators in optimizing AST interpreters. ACM SIG-PLAN Notices, 49(10):727– 743, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeyda:2014:CMS

[ZLCW14] Frank Zeyda, Lalkhumsanga Lalkhumsanga, Ana Cavalcanti, and Andy Wellings. Circus models for safety-critical Java programs. The Computer Journal, 57(7):1046–1091, July 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/7/1046.
full.pdf+html.

Zabolotnyi:2015:JCG

[ZLHD15] Rostyslav Zabolotnyi, Philipp Leitner, Waldemar Hummer, and Schahram Dustdar. JCloudScale: Closing the gap between IaaS and PaaS. ACM Transactions on Internet Technology (TOIT), 15(3):10:1– 10:??, September 2015. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zheng:2018:ADS

[ZLNP18] Guolong Zheng, Quang Loc Le, ThanhVu Nguyen, and Quoc-Sang Phan. Automatic data structure repair using separation logic. ACM SIGSOFT Software Engineering Notes, 43(4): 66, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2014:ARP

[ZMG^{+}14] Xin Zhang, Ravi Mangal, Radu Grigore, Mayur Naik, and Hongseok Yang. On abstraction refinement for program analyses in Datalog. ACM SIGPLAN Notices, 49(6):239–248, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhou:2016:IRO

[ZMM^{+}16] Minghui Zhou, Audris Mockus, Xiujuan Ma, Lu Zhang, and Hong Mei. Inflow and retention in OSS communities with commercial involvement: a case study of three hybrid projects. ACM Transactions on Software Engineering and Methodology, 25(2):13:1–13:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2014:HTB

[ZMNY14] Xin Zhang, Ravi Mangal, Mayur Naik, and Hongseok Yang. Hybrid top-down and bottom-up interprocedural analysis. ACM SIG-PLAN Notices, 49(6):249– 258, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zakkak:2014:JJM

[ZP14] Foivos S. Zakkak and Polyvios Pratikakis. JDMM: a Java memory model for non-cache-coherent memory architectures. ACM SIG-PLAN Notices, 49(11):83– 92, November 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zibin:2010:OIG

[ZPL^{+}10] Yoav Zibin, Alex Potanin, Paley Li, Mahmood Ali, and Michael D. Ernst. Ownership and immutability in generic Java. ACM SIG-PLAN Notices, 45(10):598– 617, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zerzelidis:2010:FFS

[ZW10] Alexandros Zerzelidis and Andy Wellings. A framework for flexible scheduling in the RTSJ. ACM Transactions on Embedded Computing Systems, 10(1):3:1– 3:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2013:EAZ

[ZW13] Daming Zhu and Lusheng Wang. An exact algorithm for the zero exemplar breakpoint distance problem. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(6):1469–1477, November 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Zhu:2015:APL

[ZWSS15] Xiaoyan Zhu, E. James Whitehead, Jr., Caitlin Sadowski, and Qinbao Song. An analysis of programming language statement frequency in C, C++, and Java source code. Software—Practice and Experience, 45(11):1479–1495, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2014:CSP

[ZWZ^{+}14] Zhijia Zhao, Bo Wu, Mingzhou Zhou, Yufei Ding, Jianhua Sun, Xipeng Shen, and Youfeng Wu. Call sequence prediction through probabilistic calling automata. ACM SIGPLAN Notices, 49(10):745–762, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2016:NVC

[ZXL16] Kebo Zhang, Hailing Xiong, and Chao Li. A new version of code Java for 3D simulation of the CCA model. Computer Physics Communications, 204(??):214– 215, July 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465516300741.

Zhou:2019:AJM

[ZYY^{+}19] Yu Zhou, Xin Yan, Wenhua Yang, Taolue Chen, and Zhiqiu Huang. Augmenting Java method comments generation with context information based on neural networks. The Journal of systems and software, 156(??):328–340, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121219301529.

Zhang:2012:SRB

[ZYZ^{+}12] Yuan Zhang, Min Yang, Bo Zhou, Zhemin Yang, Weihua Zhang, and Binyu Zang. Swift: a registerbased JIT compiler for embedded JVMs. ACM SIG-PLAN Notices, 47(7):63– 74, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Zhang:2013:IMF

[ZZK13] Lingming Zhang, Lu Zhang, and Sarfraz Khurshid. Injecting mechanical faults to localize developer faults for evolving software. ACM SIGPLAN Notices, 48(10): 765–784, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.